A Complete Bibliography of Publications in IEEE Transactions on Software Engineering (2000–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 February 2018 Version 1.00

Title word cross-reference

3 [HCC^{+}08]. ^{TA} [DHS09]. N [MP00]. π [NPPW09].

-Calculus [NPPW09]. -dimensional [MP00].

10th [vDS05]. 15504 [EB00].

2002 [Gri03]. 2004 [Rot05]. 2009-2010 [Ano08-28].

60-Year [Ano06-30, Ano06-29, Ano06-28].

9 [GR01, SJJ06]. ’99 [WW00].

Abstract [GGV04, NPD01, SLLL07b].

Abstractions [SYFP08, DMCS05, SSE01]. acceptance [RHD02]. Access [MBGM09, BBG06]. accesses [YBT01]. accuracy [SHP01]. Achieving [SP05]. ACM [DY08a]. across [BMW02, YCCX09]. Actions [BCH^{+}07]. activation [LRS00]. active [ZC06]. activities [HCM02a, HCM02b]. activity [EW04]. Adaptable [ZME06]. Adaptation [CPS08, CC08, BSHL01, LL03, MA04]. Adaptive [ABB^{+}00, AP07, FN01]. Addendum [BEK^{+}04]. Adding [CFP^{+}03]. Address [Ano05-29, Kra09b, Kra08c, RL00]. Adequacy [RCW08]. adjustment [PS04]. Advanced [BM00]. Advancing [HDS06]. Advertisement [Ano04f, Ano04h, Ano04j, Ano04e, Ano04b, Ano04g, Ano04i, Ano04k, Ano04c, Ano04l, Ano04d, Ano05l, Ano05g, Ano05m, Ano05i, Ano05k, Ano05e, Ano05b, Ano05h, Ano05n, Ano05j, Ano05f, Ano05c, Ano05d, Ano06c, Ano07a, Ano07b, Ano08r, Ano09b]. AE [Ano04-30, BCH04, Kni03, Kni04a, Kni04c, Kni04b]. AEIC [Kni03]. Affective [Ano09i]. affine [MP00]. after [PRS04]. against [AH00, DVJP08]. agent [CLZ02, DIH00, XYDD03, XS03]. agent-oriented [XS03]. aggregating [GHR01]. Agile [OEGQ07]. ahead [Yan02]. aid [SRC05]. al [OAF^{+}04]. algorithm [CW02a, GHR01, KT00, VFdM^{+}03, WCJ02]. Algorithms [LHH07, BDM01, BHHK03, Pik06, Rus99]. Allocation [NTR09]. Alternative [VEBD07]. alternatives [TGK^{+}04]. among [FBS05]. Analogy [KKJ08, ATG^{+}06, KKJ08]. Analogy-Based [KKJ08, ATG^{+}06]. Analogy-X [KKJ08]. Analyses [KLZR06]. Analysis [AR07, ABLN06, Bak07, BH04, CGV09, CZvD^{+}09, CPR07, DXLN07, ER08, HLMN08, HB02, KT09, KMS04, LW07b, PD07, RCR06, RMR04, RAW08, SSG^{+}07, SR06, SH00, SLLL07a, SLLL07b, SBV01, TW07, WHY06, WS07, YCCX09, ZL06, vMW06, AZ03, BNL05, BGK^{+}02, BFGP03, BS00, BC03, BP04, BFSV04, BSV05, CDM03, CRL01, CG01, CGPA^{+}05, DS03, DN02, Egy03, EB00, EBN02, FO00, GZ05, GL05, GPHG^{+}03, GH02, GL00b, HLP01, HW02a, HT02, IR01, JKS04, JY06, JMO04, KLK00, KBYC04, KCC04, LM04, MYC05, Neu02, OFdL^{+}04, PSJ00, RDKN03, Spi03, SK03, TYB06, Ton01, Ton03, Vic01, VZJ03, WSY^{+}04, Wan05, WS06a, XHD02, YA02, ZWN^{+}06, dBFT00, Rot05]. analytic [Cha00]. Analyzing [BA08, MSO01, XS05, MPS02, SA03]. anchoring [PS04]. Anniversaries [Ano06-30, Ano06-29, Ano06-28]. annotated [DLDvL05]. Announcements [TW07]. Annual [Ano04m, Ano06a, Ano07c, Ano08a]. anomalies [LM04]. anonymity [VR03]. anonymous [Cam01]. Answering [SMD08]. API [SRK07, XS07]. API-Based [SRK07]. API-Evolution [XS07]. applets [MR00]. Applicability [SJJ06, BMW02, GR01, ZX02]. Application [Ano08p, Ano09r, DVJP08, GHM08, KMSL08, LPRT07, RCR06, BT03, CDF01, ERKF05, HLT^{+}04, KBYC04, LRSF04, MKL^{+}04, SC02, SC03, TGK^{+}04, YSCO04, ZZC01]. application-level [LRSF04]. application-specific [KBYC04]. Applications [DYZ07, HOM08, SSG^{+}07, ZZZ07, AW04, CLZ02, CCW03, CEM03, CW00, DAP01b, FMRW05, HCC04a, HCC04b, JR06, LGF05, NPGM00, RSSS02, WIK^{+}04]. applied [PSJ00]. Applying [HALM08, SSG^{+}07, JDM01, PS04]. Apportioning [IR01]. Approach [APS07, BJ09, BG09, CH06, CDS08, CM09, DM06, EA08, GA08, HWM06, LFY^{+}06, SWE09, VDDR07, BKO05, CW00, CFTV05, DMM01, Egy03, EMT05, GL05, HS01, HCC03, JMO04, Kho02, LD03, Lev00, LRS00, MB00a, MYC05, NFLJ06, RR00, WV00]. Approximate [HWM06]. Arbitrary [KW07]. Architectural [GPHG^{+}03, PIM09, MMRM05, NKL05, VDGD05, XYDD03]. Architectural-level [GPHG^{+}03]. Architecture [DP09, GHM08, MB07, MLLF09, PSJ00, Big04, BW02, Bra00, DN02, MT00, MIB04, Rav02, TSJ00b, TSJ00a, WMC01, YA02]. Architecture-based [PSJ00]. architecture-independent [Bra00, TSJ00b, TSJ00a]. architecture-level [YA02]. Architectures [SAG^{+}06, FG02]. arcs [GH02]. Arguments [LW07b, Lit00]. aritlces [Ano07a, Ano07b]. ARMOR [WIK^{+}04]. ARMOR-based [WIK^{+}04]. array [GL00b]. arrays [YCP06]. art [Men02]. artifact [PS04]. Artifacts [Rei06]. Artificial [HJ08]. ask [Jor04]. Asking [SMD08]. aspect [WBY05, XN06]. aspect-oriented [WBY05, XN06]. aspectizable [TC05]. Aspects [BCH^{+}06]. Aspectual [ALS08]. assertion [JS01]. Assessing [ABLN06, ACDD04, BMW02, BDL04, ED05, PULPT02, EGK^{+}01]. Assessment [NMJ09, BD02, Cha03, GS05, Jor04, Jor05, Tia02, TC05]. Assessments [DR06b, JG09]. asset [MYC05]. assignment [SHP01]. Assist [CH09]. assists [DIH00]. Associate [Ano07r, Kni05b, Kra06c, Kra06b, Kra08b, Kra09a, Ano05-28, Kni05a, Kra08a]. Association [Mil07, SSCM06]. Associations [Mil07, ZLX09, SP03]. Assurance [PYM^{+}07, YPK^{+}07, dLAF06, CYK^{+}05, PL03, WBY05]. Asynchronous [PDB00, Mit05, RP00]. Atomicity [YCCX09, Cam01, FFQ05, WS06a]. atomicity-generating [Cam01]. attribute [Yan02]. Attributes [MGF07, MDDG07, ZZ07, VDGD05]. Aug [HRD08]. augmented [GH02]. authentication [SW02]. Author [Ano01b, Ano02a, Ano02b, Ano03a]. Authors [Ano04-42, Ano04-45, Ano04-46, Ano04-56, Ano04-55, Ano04-59, Ano04-57, Ano04-60, Ano04-63, Ano04-61, Ano04-64, Ano04-62, Ano04-58, Ano05-61, Ano05-62, Ano05-56, Ano05-65, Ano05-55, Ano05-54, Ano05-60, Ano05-58, Ano05-63, Ano05-57, Ano05-59, Ano05-64, Ano06-56, Ano06-53, Ano06-52, Ano06-50, Ano06-55, Ano06-57, Ano06-49, Ano06-54, Ano06-48, Ano06-46, Ano06-47, Ano06-51, Ano07-53, Ano07-52, Ano07-54, Ano07-58, Ano07-60, Ano07-57, Ano07-51, Ano07-61, Ano07-59, Ano07-56, Ano07-50, Ano07-55, Ano08-33, Ano08-32, Ano08-30, Ano08-29, Ano08-31, Ano08-34, Ano09-28, Ano09-30, Ano09-27, Ano09y, Ano09-29, Ano09z]. Automata [DHQ^{+}08, SYFP08]. Automata-Based [SYFP08]. Automated [CW00, KT09, WBY05, HS02, MPS01]. Automatic [BD03, CG01, DR09, FFH04, HW02a, JY06, MP00, Mil02, NFLJ06, VDGD05, WH05, GL00b, MS00b, MM06]. Automatically [AH00]. automating [MA04]. automation [MG02]. automorphism [WSY^{+}04]. Available [KMSL08]. avoidance [WCJ02]. Avoiding [DeL01a, DeL01b]. Aware [HOM08, HVC09, LPRT07, BCMS03, CEM03, CC03, JR06, MMRM05, ZBN^{+}04]. axioms [Pik06].

Back [Ano04p, Ano04n, Ano04q, Ano04t, Ano04r, Ano04u, Ano04s, Ano04o, Ano05v, Ano05q, Ano05z, Ano05u, Ano05r, Ano05w, Ano05p, Ano05o, Ano05s, Ano05x, Ano05t, Ano05y, Ano06h, Ano06d, Ano06n, Ano06k, Ano06j, Ano06m, Ano06o, Ano06g, Ano06l, Ano06f, Ano06e, Ano06i, Ano07h, Ano07g, Ano07i, Ano07m, Ano07o, Ano07l, Ano07f, Ano07p, Ano07n, Ano07e, Ano07j, Ano07k, Ano08g, Ano08d, Ano08c, Ano08e, Ano08f, Ano08h, Ano09h, Ano09g, Ano09d, Ano09f, Ano09e, Ano09c, BHS05]. Background [CNP08]. Backward [SD06a]. base [NKL05]. Based [Ano08i, Ano09j, Ano09k, BHM06, BDFZ08, CPS08, CM09, DR06a, HCC^{+}08, KP09, KKJ08, Kou06, KRM06, LW07b, LFY^{+}06, MBGM09, MPL07, PMT^{+}08, PGM^{+}07, RCW08, SSG^{+}07, SRK07, SR06, SYFP08, SCDP07, UBKW05, VDDR07, ZLW07, ATG^{+}06, BDIS04, Big04, BDB00, BLW03, BDL04, BLDYB05, CHCC03, CGPA^{+}05, CDF01, DH00, Dol00, EFYB04, FKGS04, FRC^{+}02, GL05, Gor00, GS05, HS01, HS02, HL03, IYY^{+}05, Jor05, KKI02, Kao03, LEH01, LC04, LS03b, LGF05, MY04, MG02, MYC05, MRS02, MSO01, NAD03, PP02, PSJ00, PP03, RM02, SMBK03, Sel05, SD06b, TRW03, TRL04, TA04, VZJ03, WSY^{+}04, WS06b, WIK^{+}04, XS03, XC03, Yan02]. Basis [Moo09, SG00]. Bayesian [DXLN07, HFMN09, MM08b, PD07, WGC02]. BBN [LW07b]. BDD [WSY^{+}04, Wan05].

BDD-based [WSY^{+}04]. BDD-like [Wan05]. Behavior [PV02, UBC09, DLDvL05, DAP01b, Yan02]. Behavioral [CPS08, CJ03, HK02, UKM03]. behaviorally [SJLY00]. Benchmarking [LBMP08, JWB02]. benefit [BH03]. Benefits [DAB08]. Best [MCHL06]. Better [Tia02]. Between [CFM08, ACC^{+}02, CG01, DC06, HCC03, JY06, JDM01, MNS01, ZLX09]. bidding [JC04, KPLJ03]. binaries [LKV05]. Bisimulation [SD06a]. biting [SBMJ01]. black [BHS03, XN05]. blueprint [DL05]. bootstrap [LS03a]. bounded [CYK^{+}05]. Bounding [LLH^{+}01]. bounds [PSMK03]. box [XN05]. branch [HW02a]. Branching [dAFS09]. bridging [MNS01]. brokerages [MKL^{+}04]. browsing [DIH00]. Budget [NH09]. bug [WH05]. Buggy [KWZ08]. bugs [LLMZ06]. Build [dJ05]. Build-level [dJ05]. building [Lev00]. Bunch [MM06]. Business [ZZZ07, DMCS05, VZJ03, WCD^{+}05]. Bytecode [SMBZ07, JW01].

c [EBN02, CCG^{+}04, ESWH04, SHFJ08]. cache [LLH^{+}01]. cache-related [LLH^{+}01]. calculation [BNL05]. Calculus [NPPW09, San03, dBFT00]. calibrating [BBS02a]. Call [Ano04-28, Ano04-29, Ano05-27, Ano07q, Ano08i, Ano09i, Ano09j, Ano09k, DAP01a, MM08a, OJ02]. Call-for-Papers [Ano07q]. Call-Stack [MM08a]. calls [BD03, LKV05]. candidate [HDS06]. Capability [JG07, EB00, MP03]. capture [BEFL00, EL01]. capture-recapture [BEFL00, EL01]. Career [Ano08q]. CARISMA [CEM03]. Carving [ECDJ09]. Case [DM08, DR06b, LHH07, MPL07, MLLF09, NMJ09, VJB09, ABGM02, BS00, CPT03, DTB05, EMR02, Kao03, MPS01, MRS02, NFLJ06, SP05, TGK^{+}04, WV00, BBS02a, CDFG06]. Cases [ECDJ09, JG07, MO00, MX05, RUCH01]. catalog [DGR04]. CatchUp [XS07]. Categorization [YSCO04]. category [CPT03]. category-partition [CPT03]. Cause [EZS^{+}08]. Causes [RCR06]. CCFinder [KKI02]. Celebrates [Ano06-30, Ano06-28, Ano06-29]. Center [Ano08q]. centered [Lev00, MB00b]. centralized [AS04, Lee04]. Centric [CHMB08, LL03]. Certifiably [HALM08]. Chain [SD06a]. Chains [BCH^{+}07, BHHK03]. Change [DLL09, FWPG07, RCR06, RAW08, SMD08, ZLX09, CHCC03, EGK^{+}01, GKMS00, KT05, TCS05, VMA^{+}04, YMNCC04]. Change-Proneness [ZLX09]. Changes [KWZ08, EGK^{+}02, PP05, YMNCC04, ZZWD05]. Channel [Mit08]. characteristics [Kel06]. characterization [YCP06]. Characterizing [Mit08]. CHARMY [PIM09]. Charts [SC06, AEY03, BHS05]. check [HH06]. Checker [BCY01, HB07, Hol97]. Checking [BCH^{+}07, DHS09, HKB09, HB08, NPPW09, SD06a, XN05, AZ03, AH00, BHHK03, BKO05, CAB^{+}01, ESWH04, GCD03, San03]. checklist [LEH01, TRW03]. checklist-based [TRW03]. Chief [Ano08-28, Kni06, Kra06a]. choice [CPT03]. CK [SK03]. Claims [LW07b, DS00]. Class [Cas09, CFTV05, Min06, RMR04, SWE09, ZLX09, BLW03, BDL04, DL05, EBGR01, EBG^{+}02, Eva03, KBYC04, MS00b]. Class-Oriented [Cas09]. Classes [HRD07, HRD08, MPF08, OEGQ07, VSC09, CKB04, ESWH04]. Classification [LBMP08, MT00, Yan02]. Classifications [VJB09]. Classifying [HKL^{+}07, KWZ08]. Clean [KWZ08]. client [HCC04a, HCC04b, MG00, RP00]. client-server [RP00]. client/server [HCC04a, HCC04b, MG00]. clock [HW02b]. Clone [BKA^{+}07, BvDvET05, KKI02]. Clones [Bak07, BJ09]. closed [PSE04]. closed-source [PSE04]. Clustering [KT00, LPRT07, MB07, AT05, Tia02]. CMM [AC07, MP03]. Co [DLL09]. Co-Change [DLL09]. COCOMO [BBS02a]. Code [BCH^{+}06, FWPG07, KP09, ML09, MGF07, MDDG07, SCDP07, ZZ07, ACC^{+}02, BD03, BEK^{+}04, BvDvET05, DRW03, EGK^{+}01, EKS03, GZ05, HCM02a, HCM02b, HH06, KKI02, LEH01, LLMZ06, MIB04, PP05, RCM04, Vok04, WH05, YMNCC04]. codesign [BZCJ02]. Cognitive [PS04, HHC04, MY04]. cognitive-based [MY04]. Cohesion [MPF08, CKB04]. Collect [KMCA06]. collection [JMO04]. collocation [TCKO02]. Combining [BZCJ02, NMJ09, FBV05, LPSS00]. Comments [BCY01, BT03, CJ03, Eva03, FMPR06, JKS04, JY06, Kao03, MDDG07, MET03, RGRF01, SC03, YSW^{+}01, ZX02, ZZ07, BNR06, OAF^{+}04]. Commerce [ZZZ07, AW04]. commercial [Cha00]. Common [HGP09, YSCO04]. commonality [MYC05]. Communicating [MD00]. Communication [Mit08, BDM01, GGV04, HM03, HL03, LH03]. CoMoM [Cas09]. Companies [ASM09]. Company [KMT07]. Comparative [MM08b, MO00, MSS05]. Comparing [ABLN06, BSC03, KT05, SK01, TM05, PUT^{+}01, XN05]. Comparison [BKA^{+}07, LEH01, MT00, MOJ05, RHD02, TRW03]. comparisons [Nta01]. Complex [AR07, DC06, FO00, LZBQ04, YCP06]. Complexity [AGDS07, CRL01, BVT03, DKST05, GML05, GR01, SK03, ZX02]. compliant [SP05]. comply [MP03]. Component [Hat09, Kou06, LW07a, PMT^{+}08, ZME06, Dol00, LGF05, MA04, RDKN03]. Component-Based [Kou06, PMT^{+}08, Dol00, LGF05]. Components [CPS08, KMSL08, LCS^{+}08, RL00, BSHL01, CCG^{+}04, DS03, IYY^{+}05, LHS03, PV02, VZJ03, dJ05]. composite [MKL^{+}04]. Composition [AP07, YCCX09, ZBN^{+}04]. Compositional [BHH^{+}09, BDHK06, XHD02, ZC09]. Compositions [DYZ07]. Comprehension [CZvD^{+}09, DHOH03, HCM02a, HCM02b, SGMB03]. comprehensive [BEFL00]. Compression [SMBZ07]. computation [EH01, RSSS02]. Computational [ABB^{+}00, DS03]. Computer [Ano04-39, Ano04-41, Ano04-44, Ano04-42, Ano04-40, Ano04-43, Ano06-28, Ano06-30, Ano06-29, Ano06-45, Ano06-44, Ano07a, Ano07b, Ano07-46, Ano07-47, Ano07-48, Ano07-45, Ano07-49, Ano08p, Ano08q, Ano08r, Ano09b, Ano09r, HKT09, VR03]. Computer-mediated [VR03]. computers [CCL^{+}00]. Computing [SMBZ07, WW00, dLAF06, CC03, LL03, Sat03, ZC06, Ano09i]. Concept [SSG^{+}07, Ton01, Ton03]. Conceptual [MPF08, AAG^{+}04, CdPL04]. concern [BvDvET05]. Concerns [EZS^{+}08]. concurrent [IR01, KCT02, LC06, LHS03, MG02, WSY^{+}04, dBFT00]. condition [JH03]. condition/decision [JH03]. Conditions [CPY08, Mit05]. Conference [Ano03c, BKM07, CTM09, DT04, FKW06, Gri03, GN06, GR06, vDS05, BDB00, CA03, HS03, KGR01, DY08a]. Confidence [LW07b, LS03a]. Configurable [CDS08, SSE01]. configuration [YCP06, vdHCHW02]. configurations [PBG04]. Confirming [PBG04]. conformance [CJMR07, EFYB04]. Confounding [ZLX09, EBGR01, Eva03]. Congress [WW00]. conjunctive [CW02a]. consensus [GS01]. considering [CKB04]. Consistency [KLZR06, CG01, JY06]. constrained [MMRM05]. constraint [Bol00]. constraint-object-oriented [Bol00]. Constraints [CDS08, HW02a, MKL^{+}04]. Constructing [CDS08, CDFG06, Moo09, MY04]. construction [MC02, SP03, WS06b, ZZC01]. constructs [SH00]. Container [HRD07, HRD08]. Content [PD07, BEFL00, LL03, PRS04]. Contents [Ano04z, Ano04-27, WCL04]. Context [BCMS03, HWM06, CLZ02, CEM03, CC03, JR06]. Context-aware [BCMS03, CEM03, CC03, JR06]. context-dependent [CLZ02]. contextual [SP03]. continuity [JWB02]. Continuous [PYM^{+}07, YPK^{+}07, BHHK03, MNH01]. continuous-time [BHHK03]. Contract [ESWH04, LBJ06]. Contract-checking [ESWH04]. Control [HRJ08, MBFSV07, MBGM09, ZC09, AS04, ABGM02, HP00, HCM02a, HCM02b, JS02, WBY05, WS06b, dBFT00]. control-flow [dBFT00]. controlled [BBD01, PUT^{+}01, PULPT02, SHH^{+}05]. controller [HLP01, MS00a]. Controversies [BHS03]. conversations [FBS05]. converters [WO00]. coordination [CLZ02, MC02]. copy [LLMZ06]. copy-paste [LLMZ06]. CORBA [CFP^{+}03]. core [MYC05]. Correct [SWE09, DS03]. Correction [HCC04a, SSCM06]. Corrections [DeL01b, HCM02a]. Correctness [BSV05]. Cost [JS07, KKJ08, KMT07, YHJ08, ATG^{+}06, BH03, FBV05, Jor05, KHCP00, SEM01, XY03]. cost-effective [KHCP00]. cost-effectiveness [FBV05]. Costs [DAB08]. Counterexample [HKB09]. coupling [ABF04, YSCO04]. cover [Ano04p, Ano04n, Ano04q, Ano04t, Ano04r, Ano04u, Ano04s, Ano04o, Ano04-35, Ano04-32, Ano04-37, Ano04-31, Ano04-36, Ano04-38, Ano04-33, Ano04-34, Ano04-48, Ano04-53, Ano04-47, Ano04-52, Ano04-50, Ano04-51, Ano04-54, Ano04-49, Ano05v, Ano05q, Ano05z, Ano05u, Ano05r, Ano05w, Ano05p, Ano05o, Ano05s, Ano05x, Ano05t, Ano05y, Ano05-36, Ano05-40, Ano05-33, Ano05-30, Ano05-35, Ano05-37, Ano05-31, Ano05-32, Ano05-34, Ano05-41, Ano05-39, Ano05-38, Ano05-52, Ano05-45, Ano05-47, Ano05-43, Ano05-44, Ano05-51, Ano05-48, Ano05-42, Ano05-49, Ano05-46, Ano05-53, Ano05-50, Ano06h, Ano06d, Ano06n, Ano06k, Ano06j, Ano06m, Ano06o, Ano06g, Ano06l, Ano06f, Ano06e, Ano06i, Ano06z, Ano06p, Ano06y, Ano06x, Ano06s, Ano06v, Ano06w, Ano06t, Ano06r, Ano06u, Ano06-27, Ano06q, Ano06-42, Ano06-32, Ano06-41]. cover [Ano06-35, Ano06-39, Ano06-34, Ano06-37, Ano06-43, Ano06-40, Ano06-38, Ano06-36, Ano06-33, Ano07h, Ano07g, Ano07i, Ano07m, Ano07o, Ano07l, Ano07f, Ano07p, Ano07n, Ano07e, Ano07j, Ano07k, Ano07x, Ano07-29, Ano07v, Ano07y, Ano07-27, Ano07-28, Ano07t, Ano07-30, Ano07u, Ano07z, Ano07s, Ano07w, Ano07-38, Ano07-43, Ano07-36, Ano07-42, Ano07-34, Ano07-44, Ano07-40, Ano07-39, Ano07-41, Ano07-35, Ano07-33, Ano07-37, Ano08g, Ano08d, Ano08c, Ano08e, Ano08f, Ano08h, Ano08m, Ano08n, Ano08l, Ano08k, Ano08j, Ano08o, Ano08z, Ano08w, Ano08v, Ano08y, Ano08x, Ano08-27, Ano09h, Ano09g, Ano09d, Ano09f, Ano09e, Ano09c, Ano09m, Ano09o, Ano09n, Ano09p, Ano09q, Ano09l, Ano09t, Ano09v, Ano09u, Ano09w, Ano09x, Ano09s]. Coverage [ABLN06, CH09, MM08a, DS00, JH03, MB03]. Covering [YCP06]. CP [LLMZ06]. CP-Miner [LLMZ06]. craft [HLP01]. Criteria [ABLN06, RCW08]. criterion [FSKM03, Kel06]. Critical [CHMB08, dLAF06, LM04, PMM00]. Cross [KMT07]. Crosscutting [EZS^{+}08, BvDvET05]. Cryptographic [DS00]. CSDP [Ano08r]. CSL [DHS09]. CTK [SSE01]. CTTE [MPS02]. currencies [Cam01]. Current [PRT00b, PRT00a]. customer [SBV01]. customizable [DS03]. Cycle [AC07, NH09]. cycles [SA02].

Data [BJ09, DR06a, DVJP08, DM06, HVC09, HGP09, KP09, KLZR06, MGF07, MDDG07, ZZ07, ABGM02, BW05, Bra00, EGK^{+}01, ERKF05, FBV05, GL00b, JMO04, KHL01, LG03, MMS01, MB00b, MSO01, NPGM00, SC01, SG00, SEM01, Tia02, TRL04, WSY^{+}04, Wan05, YBT01]. Data-Aware [HVC09]. data-flow [GL00b]. data-parallel [NPGM00]. data-structures [Wan05]. database [HCC04a, HCC04b]. Datatypes [APS07]. DBM [CGV09]. Deadline [BP04]. deadlock [WCJ02]. Deadlocks [Mit08, Lee04, VFdM^{+}03]. Debugging [LFY^{+}06, CFC01, XY03]. Decay [HCC^{+}08, EGK^{+}01]. decision [Cha00, JH03]. decision-analytic [Cha00]. decisions [EW06]. decomposition [Ton03, WBY05]. decorate [BW02]. deep [SRC05]. def [SP03]. def-use [SP03]. Defect [HCC^{+}08, KZEL09, LBMP08, MGF07, MDDG07, Vok04, ZZ07, BH03, BEFL00, PRS04, SSCM06]. Defects [EZS^{+}08, HCC^{+}08, ML09, SK03]. defined [TFD03]. Defining [JDM01]. definition [Bou00, BMB02]. Degradation [YPK^{+}07]. Delay [JTMZ01, LLH^{+}01]. delegated [AS04]. Deliberate [HJ08]. delivery [MRV02]. Delta [dLM08b]. DeMIMA [GA08]. Dense [VSC09, WHY06]. Dense-Time [VSC09, WHY06]. Density [CGV09]. Dependability [BHH^{+}09, BFGP03, MO00, WBY05, WIK^{+}04]. dependable [MRV02]. Dependence [CPY08, BH04]. Dependencies [CMRH09, DVJP08, dLAF06]. dependency [Egy03]. dependent [BR03, CLZ02, CKB04, Vic01]. Deployed [HKL^{+}07, ED05]. deployment [BSHL01, ZC06]. derive [RR00]. description [BKO05, MT00, Rav02]. descriptions [MPS04]. Design [ABB^{+}00, ART03, BHM06, BDFZ08, BCRW00, CCW03, CFM08, DYZ07, GA08, HJ08, HB07, KP09, LBJ06, LGF05, LCN08, MNH01, MBFSV07, MPS04, SD06b, TCSH06, YBT01, ZZC01, ZL06, BVT03, BD02, BGM04, BR03, BW02, BW01, CCF^{+}00, Cha03, DMCS05, EW06, Gor00, HH06, HLT^{+}04, Kel06, MS00a, MG00, MG02, MKL^{+}04, MPS02, MNS01, NPGM00, PUT^{+}01, PULPT02, SK03, TYB06, VZJ03, Vok04, WS06b, XS05, XS03]. Design-level [LGF05]. design-time [EW06]. Designing [LRS00, PIM09]. designs [BBD01, CG01, JY06]. detect [CW02a, GZ05]. Detecting [BJ09, KW07, LKV05]. Detection [BKA^{+}07, JG07, TCSH06, VDDR07, AV01, BD03, BH03, BvDvET05, FFH04, HW02a, KKI02, Lee04, LPSS00, MX05, ZWN^{+}06]. determinants [GKMG02]. Determining [FBV05, MJJ^{+}07]. Developed [ASM09, OEGQ07]. developer [RHD02]. Developers [KMCA06, SRS^{+}09, RCM04]. developing [MYC05, MPS02, OBMP03]. Development [BG09, BLHS^{+}09, CFM08, JS07, LCS^{+}08, MLLF09, NH09, OEGQ07, WW00, BDIS04, BT03, Bol00, BW01, BLDYB05, CMSB05, CDF01, DTB05, DRW03, GKMG02, HP00, HM03, Jor05, JR06, Kao03, LZBQ04, MOJ05, MPS04, MRS02, PSR05, PPW^{+}02, RGFR98, RGRF01, RL00, SJLY00, Sel05, SC02, TCKO02, XY03, YSW^{+}01, SC03]. Diagnosing [LI00]. diagnosis [BD03, XC03]. Diagrams [APS07, BLL06, Mit08, EW04, HCM02a, HCM02b, HHC04, Mil02]. diamonds [BHS03]. Diff [XS07]. Differencing [FWPG07]. different [AL03]. Differential [ECDJ09]. Digital [Ano07a, Ano07b]. dimensional [MP00]. Directed [DR06a]. Disaggregating [BBS02a]. Disclosure [CCR07]. Discovered [ML09]. Discovering [CPY08, HRD07, HRD08, SAG^{+}06, ECGN01]. Discrete [KT09]. Discrete-Event [KT09]. Disruptive [VEBD07]. Disseminate [CCR07]. distance [JDM01]. Distilling [FWPG07]. distinguishing [ZyC03]. Distributed [BLL06, DLM08a, Kou06, PJJ^{+}07, PYM^{+}07, RCW08, YPK^{+}07, BSHL01, BDB00, BG00, CW02a, GGV04, HP00, HM03, HS02, HW02b, Kho02, Lee04, MMRM05, MP00, MT04b, Rav02, SZ02, WCJ02]. Distribution [Zha08]. Distributions [AR07, Hat09]. diverse [DRW03, LPSS00, PSMK03]. Diversity [vdMR08, Lit00]. Do [EZS^{+}08]. Documentation [HRD07, HRD08, LZBQ04, ACC^{+}02, ABHL06, PULPT02]. documents [LEH01]. Does [EGK^{+}01]. Domain [Bot05, CM09, DH00, EW05, HT02, MYC05]. Domain-Specific [CM09]. Domains [CGV09]. dominance [SGMB03]. Driven [EA08, LPRT07, MLLF09, BMB02, BP04, Egy03, LZBQ04, NFLJ06, XN06]. dual [BF00]. dual-tier [BF00]. Dup [Bak07]. during [JG07, KMCA06, LM04, SMD08]. Dynamic [ABF04, BSHL01, CC08, CZvD^{+}09, HFMN09, VEBD07, CW02b, DAP01b, MACE05, MC02, Rav02, SM03a, Vic01]. Dynamically [ECGN01].

e-commerce [AW04, ZZZ07]. e-Transactions [FG02]. Eager [VDDR07]. early [SRC05]. Economic [BG09]. ED [HCC^{+}08]. Editor [Ano05-28, ABGM02, Gri03, HMHJ05, Kni05a, Kni06, Kra06a, Kra08a, Rot05, Woh01, vDS05]. Editor-in-Chief [Kra06a, Kni06]. Editorial [And01, And02, Ano04-30, Ano05-28, Ano05-29, Ano07r, BCH04, HKL05, Kni02, Kni03, Kni04a, Kni04c, Kni04b, Kni04d, Kni05a, Kni05b, Kni06, Kra06a, Kra06b, Kra07, Kra08a, Kra08b, Kra09a, Kra09b, UBKW05, Kra06c]. Editors [Ano07r, Ano08-28, Kni05b, Kra06c, Kra06b, Kra08b, Kra09a, BKM07, CA03, CTM09, CCW00a, CCW00b, DPP00, DT04, ER08, FGLW09, FKW06, GN06, GR06, HB02, HS03, HKT09, KGR01, MN08, MO03, TSJ00b, TSJ00a]. Educational [CNP08]. Effect [LCN08, ZLX09, AS04, EBGR01, Eva03, XY03].

Effective [DMJN08, MBGM09, KHCP00, LL03, RCM04]. Effectiveness [CNP08, vdMR08, BH03, EMT05, FBV05, HCM02a, HCM02b, KD00, MX05, SJLY00]. Effects [WIK^{+}04, YPK^{+}07, DHOH03, LPSS00, PU01, SBV01]. Efficiency [CCR07, SP05]. Efficient [BR03, BNL05, Cas09, GL00b, KW07, KCLL03, SWE09, dLM08b, CW02a, GHR01, IR01, KLK00, RR00, WCJ02, YCP06]. Effort [AC07, CFM08, JG09, MM08b, MCHL06, NH09, BW01, FN01, HS01, Jor04, JMO04, PSR05, SHP01, SSCM06]. EFSM [PBG04]. EgoSpaces [JR06]. EIC [And02]. Electronic [DM08, MKL^{+}04]. Elements [HJ08]. elicitation [CdPL04]. Eliciting [JMSS07]. Eliminating [MO00, SA02]. Embedded [CM09, MBFSV07, SMBZ07, HLT^{+}04, JWB02, KBYC04, KCC04, SA03, WS06b]. EMERALDS [ZS01]. Empirical [CFM08, CNP08, DC06, DR06b, DAB08, GFS05, HRRW01, KP09, LM04, MPL07, OEGQ07, PD07, RAW08, SG00, SK03, TW07, ZL06, AL03, BT03, CS00, DHOH03, Dyb05, EMR02, EBN02, HM03, JC04, KPP^{+}02, MO00, MSO01, PSE04, SV02, SHP01, SC02, SC03, TC05, VR03, VZJ03, Vok04]. empirically [SMBK03]. empirically-based [SMBK03]. Employee [DM08]. employing [JS02]. Emulation [DM06]. Enabling [Sel05]. End [ABB^{+}00, DLDvL05, FG02]. End-to-End [ABB^{+}00, FG02]. end-user [DLDvL05]. Ends [Mil07]. Engineering [ASM09, Ano03c, Ano07q, Ano08i, ACS09, AGK09, BLL06, CLZ02, DC06, DY08a, DT04, FGLW09, GN06, HLMN08, HRJ08, HSD07, HJ08, HKL^{+}07, MN08, Moo09, NMJ09, Rav02, SC09, VJB09, WHY06, ZME06, vDS05, BBG06, Bou00, Bro01, HS03, HW02b, KPP^{+}02, KGR01, LD03, LS03b, MO03, NKL05, PS04, SRC05, SG00, SV02, SHH^{+}05, SDA05, vLL00, Gri03, Ano04-29, Ano04x, Ano04w, Ano04y, Ano04-27, Ano04v, Ano04z]. Enhanced [FAOW^{+}09, Neu02]. Enhancing [KMSL08, TA04]. Ensuring [VEBD07]. entities [GZ05]. Entropy [DXLN07]. environment [KT00, RBR06, SBV01, WIK^{+}04, ZZC01]. Environments [SMBZ07, BCRW00, Mil02, WCL04]. Epidemiological [AG06]. equational [LC04]. equations [MP00]. Equipped [APS07]. ERP [SM03b]. Errata [HRD08]. error [JMO04]. Errors [Bot05, MO00, XE03]. establishment [SM03a]. estimate [LS03a]. Estimating [PSMK03, BEFL00, LH03, PRS04]. Estimation [FN01, HCC^{+}08, JS07, JG09, KKJ08, KMT07, MCHL06, ZWL08, ATG^{+}06, BZCJ02, CFTV05, Dol00, HS01, HLK03, Jor04, JMO04, RL01, SHP01, SEM01]. Ethical [SV02]. evaluate [ABGM02]. Evaluating [AS04, AGDS07, EL01, RCW08, TRL04, TGK^{+}04, BBD01]. Evaluation [BKA^{+}07, BHH^{+}09, Cas09, DAB08, FKW06, GHM08, HOM08, HKT09, Kou06, LS03a, VSC09, ABHL06, BR03, BEFL00, CCW03, DRW03, FSKM03, MSO01, NPD01, SMBK03]. Event [CHCC03, KT09, CDF01]. Event-based [CHCC03, CDF01]. events [MKL^{+}04]. Evidence [Jor05, EGK^{+}01]. Evidence-based [Jor05]. Evolution [BKM07, DLL09, GR06, HKL05, XS07, AP03, ECGN01, MMS01, SMBK03]. evolutionary [CHCC03, XS05]. Evolving [YPK^{+}07, Kel06, MX05]. Examining [ZLX09]. Example [LW07b]. Examples [HWM06]. Exception [BAvZ00, HA00, BS00, BM00, FFH04, KD00, MO00, MT04b, PRT00b, PRT00a, SH00]. exceptions [CRL01]. executables [SDA05]. Execution [PGM^{+}07, dLM08b, BW05, MR00, PMM00]. Executions [HKL^{+}07]. Exercising [MPL07]. exhaustive [CYK^{+}05]. Existing [BCH^{+}06, MB00a]. Experience [WV00, SMBK03]. Experiment [Bak07, BW02, BBD01, PUT^{+}01, PU01]. experimental [ABHL06, BLDYB05, DKST05, LEH01, TRW03]. experimenting [PD00]. Experiments [HSD07, HJ08, BDL04, PULPT02, SHH^{+}05]. expert [FBV05, LS03b]. Expertise [AGDS07]. Explaining [RHD02]. exploitation [HW02a]. Exploiting [FFQ05, CDFG06]. Exploration [SEHM06, dLM08b, GCD03]. Exploratory [KMCA06, Kao03, MRS02, RCM04]. explosion [AV01, SBMJ01]. extended [Mil02]. extending [PS04]. Extensible [SLLL07a, SLLL07b]. Extension [HB07]. Extensions [NPPW09, MP03]. extracted [TRL04]. Extracting [PMT^{+}08]. Extraction [DR09, FWPG07, HS02, JW01].

facilitating [JR06]. factorial [BR03].

Factors [RGFR98, YSW^{+}01, Dyb05, MD03, MET02, MET03, RGRF01]. Failure [HGP09, LZH08, RCR06, MD03, MET02, MET03, TRL04, ZH02]. failure-inducing [ZH02]. Failures [CMRH09, FO00]. Families [CM09]. family [EMR02]. FAML [BLHS^{+}09]. Farewell [Ano05-28, Kni04a, Kni04c, Kni03]. Fast [Lee04]. Fault [AR07, HGP09, JG07, MPF08, OEGQ07, PD07, BFGP03, BMW02, DGR04, GKMS00, GFS05, KWG04, LPSS00, MX05, Pik06, Rus99, RBR06, YCP06, ZWN^{+}06]. fault-detection [MX05]. Fault-Proneness [OEGQ07, BMW02]. fault-tolerant [Pik06, Rus99]. Faults [DR06b, DM06, Zha08, ZL06, BD03, FO00, MMG03, OWB05]. Feature [AG06, ALS08, PGM^{+}07, ATG^{+}06]. features [BEK^{+}04, EKS03]. FEDERATE [BBG06]. federated [BBG06]. Field [DM06]. file [AZ03, MNH01]. Filtering [MPL07]. Filters [ZWL08]. find [XE03].

Finding [Bak07, WM08, LLMZ06, WH05]. Findings [LBMP08]. Fine [FWPG07, ART03, SA03]. Fine-Grained [FWPG07, ART03]. finite [ZyC03]. finite-state [ZyC03]. Firm [TW07]. First [WW00, EMT05, FKW06]. five [RHD02]. Fixture [VDDR07]. Flexible [AP07, AGK09, DeL01a, DeL01b, MOJ05, TM05]. flow [BZCJ02, GL00b, dBFT00]. Flows [MPL07]. Fluid [GH02]. flush [GH02]. flush-out [GH02]. FM [WW00]. forces [GMK05]. Form [KZEL09, BBS02b]. Formal [APS07, BT03, HLP01, HP00, HALM08, SC02, SC03, WW00, AH00, ACLdSS04, BGK^{+}02, BH04, BDB00, BG00, CDM02, CJMR07, FRC^{+}02, HL03, LH03, LKW00, Pik06, Rus99, SW02, XYDD03]. Formalism [BDHK06]. formality [BLDYB05]. Formalization [BHSLB03, THY03]. Formalizing [Bot05, CW02b, EW05]. formally [Gor00, LRSF04]. formally-based [Gor00]. foundation [Bro01]. Foundations [Ano03c, DY08a, Gri03, JS01, PP03]. Four [ASM09]. fractional [BR03]. Fragment [RMR04]. Framework [ACS09, GHM08, HLMN08, LBMP08, PIM09, AAG^{+}04, BBG06, BPT00, BG00, CPT03, DCC^{+}02, FRC^{+}02, HLT^{+}04, JKS04, Kao03, KBYC04, LRSF04, MT00, MRS02, MB00b, OFdL^{+}04, Sat03, SLB00, SZ02, XS03]. framework-based [Kao03, MRS02]. Framework-Specific [ACS09]. frameworks [MB00a]. FreeBSD [DTB05]. frequency [Vok04]. Front [Ano04-35, Ano04-32, Ano04-37, Ano04-31, Ano04-36, Ano04-38, Ano04-33, Ano04-34, Ano05-36, Ano05-40, Ano05-33, Ano05-30, Ano05-35, Ano05-37, Ano05-31, Ano05-32, Ano05-34, Ano05-41, Ano05-39, Ano05-38, Ano06z, Ano06p, Ano06y, Ano06x, Ano06s, Ano06v, Ano06w, Ano06t, Ano06r, Ano06u, Ano06-27, Ano06q, Ano07x, Ano07-29, Ano07v, Ano07y, Ano07-27, Ano07-28, Ano07t, Ano07-30, Ano07u, Ano07z, Ano07s, Ano07w, Ano08m, Ano08n, Ano08l, Ano08k, Ano08j, Ano08o, Ano09m, Ano09o, Ano09n, Ano09p, Ano09q, Ano09l, Ano04-48, Ano04-53, Ano04-47, Ano04-52, Ano04-50, Ano04-51, Ano04-54, Ano04-49, Ano05-52, Ano05-45, Ano05-47, Ano05-43, Ano05-44, Ano05-51, Ano05-48, Ano05-42, Ano05-49, Ano05-46, Ano05-53, Ano05-50, Ano06-42, Ano06-32, Ano06-41, Ano06-35, Ano06-39, Ano06-34, Ano06-37, Ano06-43, Ano06-40, Ano06-38]. front [Ano06-36, Ano06-33, Ano07-38, Ano07-43, Ano07-36, Ano07-42, Ano07-34, Ano07-44, Ano07-40, Ano07-39, Ano07-41, Ano07-35, Ano07-33, Ano07-37, Ano08z, Ano08w, Ano08v, Ano08y, Ano08x, Ano08-27, Ano09t, Ano09v, Ano09u, Ano09w, Ano09x, Ano09s]. FSM [EFYB04]. FSM-based [EFYB04]. function [Bou00, SM03a]. Functional [AP03, KZEL09, TYB06]. Functionalities [JMSS07]. Functions [CGV09, BSC03, SG00]. Fusion [WCD^{+}05]. future [FK05].

gap [MNS01]. General

[AZ03, Hat09, Min06, VDDR07].

generalized [BBS02b, Lee04, TFD03].

generated [BD03]. Generating

[DLDvL05, MMS01, Cam01]. Generation [HKB09, KRM06, MBGM09, PMM00, CPT03, DPP00, DS03, HDS06, KLK00, MS00b, MPS01, NFLJ06, TL02, ZZC01]. Generators [SCDP07, BCRW00, Big04, CDFG06]. Generic [BLHS^{+}09, GS01, ZZC01]. GenVoca [BCRW00]. glasses [DL05]. Global [Spi03, AV01, BH04]. globally [HM03]. Goal [CHMB08, DR06a, BMB02, vLL00]. Goal-Centric [CHMB08]. Goal-Directed [DR06a]. goal-driven [BMB02]. goal-oriented [vLL00]. Governance [CFM08]. Grained [FWPG07, ART03].

Grammar [KDM09]. Grammars [HB08, Yan02]. graph [BLW03, WSY^{+}04]. graph-based [BLW03]. graphical [BW02, WGC02]. graphics [BGM04]. Graphs [CPY08, DS03, ZD04]. Great [Ano09b]. Greedy [CDS08]. grids [BSRR02]. Group [BDM01, VR03]. groups [SM03a]. growth [Tur02]. guardian [MT04b]. Guest [HKL05, UBKW05, BKM07, CA03, CTM09, CCW00a, CCW00b, DPP00, DT04, ER08, FGLW09, FKW06, Gri03, GN06, GR06, HB02, HS03, HMHJ05, HKT09, KGR01, MN08, MO03, Rot05, TSJ00b, TSJ00a, Woh01, vDS05]. GUI [MM08a, MPS01, MX05]. Guide [DM08, ZZWD05]. Guided [KDM09, BHS05]. Guidelines [JMSS07, BBD01, Jor05, KPP^{+}02].

Hadas [BSHL01]. hammock [ZD04]. Handling [MB00b, vLL00, BS00, BM00, BAvZ00, FFH04, HA00, KD00, MO00, MT04b, PRT00b, PRT00a, SH00]. Hard [BDHK06]. hardware [BZCJ02, BLD02, LKW00]. hardware/software [BZCJ02, BLD02]. Heap [DR09]. help [EH01]. Heuristic [San03]. heuristics [PS04]. Hidden [dLAF06]. Hiding [SHFJ08]. Hierarchical [MB07, MPS01, BD02]. High [ZL06, dLAF06, BFGP03, KT05, SM03b, WBY05]. high-change [KT05]. high-level [BFGP03]. Higher [BJ09]. Higher-Level [BJ09]. highest [KT05]. Highly [CDS08, OEGQ07]. Highly-Configurable [CDS08]. Hipikat [CMSB05]. histories [ZZWD05]. history [GKMS00, XS05, YMNCC04]. homogeneous [BSRR02]. Horizontal [BHM06]. hostile [DS02]. human [Lev00, MB00b]. human-centered [Lev00]. hybrid [Wan05]. hypermedia [DAP01b]. Hypothesis [LFY^{+}06].

I/O [CW00, RSSS02]. Idealized [LW07b].

Identification [AG06, GA08, TC09]. Identifying [MP03, RCR06, SM03b, BvDvET05]. IEC [EB00]. IEEE [Rot05, Ano04-39, Ano04-41, Ano04-44, Ano04-42, Ano04-40, Ano04-43, Ano06-28, Ano06-30, Ano06-29, Ano06-45, Ano06-44, Ano07a, Ano07b, Ano07-46, Ano07-47, Ano07-48, Ano07-45, Ano07-49, Ano08p, Ano08q, Ano08r, Ano09b, Ano09r, Ano04-29, Ano04x, Ano04w, Ano04y, Ano04-27, Ano04v, Ano04z, Ano09i]. IFIP [BDB00]. II [BBS02a]. III [Bra00]. image [BW05]. image-processing [BW05]. Impact [CNP08, CMRH09, DM08, JG09, KP09, NH09, RCR06, RAW08, TW07, YHJ08, ABHL06, ABGM02, GMK05, JMO04, RGFR98, RGRF01, Ton03, YSW^{+}01]. imperfect [XY03]. implement [BPT00]. Implementation [dLAF06, ART03, AH00, DCC^{+}02, FDR04, LH03, MNH01, MMG03, MNS01, RGRF01, TYB06]. implementations [RR00]. implementing [LG03, RGFR98, YSW^{+}01]. implications [KWG04, SK03]. Importance [RAW08]. Improve [LBJ06, SHP01, WH05]. improved [VMA^{+}04]. improvement [ATG^{+}06, Dyb05]. Improving [CKB04, JG07, SA02, ZZZ07, BDL04, MB00a]. imputation [MSO01]. INCA [SA02]. incidence [GKMS00]. inclusion [VDGD05]. incomplete [SEM01]. inconsistent [Pik06]. Increase [LW07b]. Incremental [KCT02, MS00a, Rei06, EFYB04]. independent [Bra00, CCL^{+}00, RL01, TSJ00b, TSJ00a]. Index [Ano04m, Ano06a, Ano07c, Ano08a, Ano01b, Ano01c, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03d]. inducing [ZH02]. Industrial [NMJ09, TGK^{+}04, Vok04, VMA^{+}04]. Inevitability [WHY06]. Inference [AEY03, DS03, KKJ08]. inferences [RM02]. influence [DHOH03]. influences [LG03].

Information [AT05, Ano04-45, Ano04-46, Ano04-39, Ano04-41, Ano04-42, Ano04-40, Ano04-56, Ano04-55, Ano04-59, Ano04-57, Ano04-60, Ano04-63, Ano04-61, Ano04-64, Ano04-62, Ano04-58, Ano05-61, Ano05-62, Ano05-56, Ano05-65, Ano05-55, Ano05-54, Ano05-60, Ano05-58, Ano05-63, Ano05-57, Ano05-59, Ano05-64, Ano06-56, Ano06-53, Ano06-52, Ano06-50, Ano06-55, Ano06-57, Ano06-49, Ano06-54, Ano06-48, Ano06-46, Ano06-47, Ano06-51, Ano07-53, Ano07-52, Ano07-54, Ano07-58, Ano07-60, Ano07-57, Ano07-51, Ano07-61, Ano07-59, Ano07-56, Ano07-50, Ano07-55, Ano08-33, Ano08-32, Ano08-30, Ano08-29, Ano08-31, Ano08-34, Ano09-28, Ano09-30, Ano09-27, Ano09y, Ano09-29, Ano09z, CH09, DLL09, KMCA06, MPL07, PGM^{+}07, SRK07, SHFJ08, BW02, JMO04, Nix00]. Information-Theoretic [SRK07, AT05]. Infrastructure [AGK09, PYM^{+}07, CDF01]. inheritance [CJ03, GR01, HK02, ZX02]. input [ZH02]. inspection [ART03, BH03, DRW03, FBV05, MY04, PRS04, PL03, PPW^{+}02, VR03, Xu03]. Inspections [CNP08]. inspectors [EL01]. instance [CKB04]. instantiation [JKS04, OFdL^{+}04]. institutional [GMK05]. instruction [BLD02]. instrumentation [CFC01]. Integrating [CJMR07, CPR07, CW02b]. Integration [APS07, BLW03, DMM01, KCT02, MS00b]. Integrative [BG09]. integrity [DS02]. Intent [Lev00, HH06]. Interaction [CDS08, DMCS05, UBKW05, SD06b]. Interactions [PMT^{+}08, KWG04]. Interactive [RBR06, BGM04, MPS02]. Interface [DMM01, HB08, JS01, SMBK03]. interfaces [ACL02, MPS04, TC05]. internally [LEH01]. International [Ano03c, BKM07, CTM09, DT04, ER08, FKW06, GN06, GR06, HB02, Rot05, Woh01, CA03, HS03, KGR01]. Internet [BCMS03, MO03, OBMP03]. InterPlay [BHM06]. interprocedural [GL00b]. interrupt [BP04]. interrupt-driven [BP04]. Interval [BNR06, FMPR06, MN01, PPW^{+}02]. intervals [LS03a]. Int’l [BDB00]. Introduction [And02, Ano04-30, Ano05-28, Ano07r, BKM07, CTM09, CCW00b, CGH02, DY08a, DT04, ER08, FGLW09, FKW06, GN06, GR06, HB02, HMHJ05, HKT09, Kni04c, Kni04b, Kni05a, Kni05b, Kra06c, Kra06b, Kra08a, Kra08b, Kra09a, MN08, Rot05, vDS05, BCH04, CA03, CCW00a, DPP00, Gri03, HS03, Kni03, KL02, KGR01, MO03, TSJ00b, TSJ00a, Woh01]. invariant [YBT01]. invariants [ECGN01]. investigate [RCM04]. Investigating [BH03, SLB00]. Investigation [KZEL09, BLW03, BLDYB05, CS00, DHOH03, Dyb05, GML05, VR03]. IP [ZC09]. irregularities [MB00b]. ISO [EB00]. ISO/IEC [EB00]. isolating [ZH02]. Issue [Ano04-29, Ano07q, Ano09j, Ano09k, GR06, HKL05, HMHJ05, HKT09, Ano06-31, Ano07-31, Ano07-32, Ano08u, Ano08s, Ano08t]. Issues [RAW08, SV02, TSJ00b, TSJ00a, WV00, WW00]. Itanium [SDA05]. Iterative [BCMV03, OEGQ07, AL03].

Java [HRD08, BS00, BLL06, CRL01,

FDR04, FMRW05, HRD07, KMS04, LHS03, MR00, RCR06, RMR04, SMBZ07, SR06]. Java/JVM [BS00]. JEDI [CDF01]. Jiau [OAF^{+}04]. Join [Ano06-45, Ano06-44, Ano07-46, Ano07-47, Ano07-48, Ano07-45, Ano07-49]. Joining [Ano09b]. Journal [Ano05-29, Kra07, Kra08c, Kra09b]. Journals [Ano08-28]. JVM [BS00].

kernel [YSCO04]. Key [SM03a, Dyb05]. Knowledge [CCR07, MG02, VJB09, VZJ03, NKL05, SRC05]. Knowledge-based [MG02, VZJ03].

Labels [BCH^{+}07]. Language [Ano07q, AGK09, BHSLB03, Bot05, FGLW09, THY03, CCL^{+}00, CDFG06, EW05, Rav02, ZZC01]. Languages [ACS09, CM09, MT00, Spi03, TSJ00b, TSJ00a]. Large [ABB^{+}00, CMPS07, SKR08, vdMR08, KKI02, KT05, LLMZ06, OWB05, PPW^{+}02, RSSS02, Sel05, SM03a]. Large-Scale [SKR08, KT05, LLMZ06, PPW^{+}02, Sel05]. lattice [Ton03]. Law [Hat09]. Laws [CMPS07]. Layered [FAOW^{+}09, WMC01]. Lead [DC06]. Learn [MGF07, MDDG07, ZZ07]. Learned [JG09]. Learning [HFMN09, DIH00, Kao03, MRS02, PRS04, SLB00]. legacy [BCMV03]. Lessons [JG09]. Lessons-Learned [JG09]. Level [AC07, BJ09, BFGP03, GPHG^{+}03, LRSF04, LGF05, TFD03, YA02, dJ05]. levels [MP03]. Leveraging [ERKF05]. libraries [DIH00]. Library [Ano07a, Ano07b]. Licenses [MJJ^{+}07]. Lightweight [JW01, LD03]. like [Wan05]. likelihood [MSO01]. likelihood-based [MSO01]. likely [ECGN01]. limitations [ATG^{+}06]. limits [JS02]. line [MYC05]. Linear [dAFS09, Wan05]. link [HDS06]. Linking [MLLF09, SHFJ08]. links [ACC^{+}02]. Linux [YSCO04]. List [Ano04-44, Ano04-43, Ano04y, Ano06b, Ano07d, Ano08b, Ano09a, Ano00, Ano01a, Ano03b, Ano04a, Ano05a]. live [BHS05]. local [San03]. localization [RBR06]. Locating [BEK^{+}04, EKS03]. Location [PGM^{+}07, OWB05]. log [AZ03]. Logic [BNR06, FMPR06, GCD03, MN01, SW02]. Logical [HW02b, ACL02, MPS04, XS05, XYDD03]. logs [TRL04]. Longitudinal [DM08]. look [Yan02]. look-ahead [Yan02]. Low [VEBD07, ZL06].

M [HCC^{+}08]. machine [CCL^{+}00, PRS04]. machines [BHS05, NPD01, ZyC03]. Maintain [CHMB08]. maintainability [AS04, BBD01, HL03, YSCO04]. Maintenance [BKM07, CTM09, DAB08, GR06, HKL05, KMCA06, Rei06, ACDD04, ABHL06, BVT03, CA03, FN01, PUT^{+}01, PULPT02, TM05]. majorization [BSC03]. malicious [LKV05]. manage [WCD^{+}05]. Management [BG09, DC06, LCS^{+}08, Nix00, BCMS03, BBG06, EGK^{+}01, HA00, SMBK03, WMC01, vdHCHW02]. manager [RL00]. managing [CHCC03, EW06]. mapping [MP00]. Markov [BHHK03, BCH^{+}07, RM02, SD06a, ZLW07]. Markovian [BPT00, CGV09, HT02]. MAS [BLHS^{+}09]. masking [FFH04]. Matching [HWM06, BKO05, KC04]. Mathematical [Cha03, Bro01]. mathematics [Bou00]. matters [Jor04]. Maturing [VJB09]. maturity [MP03]. Maximum [DXLN07]. Maximum-Entropy [DXLN07]. Measure [ZLW07, EB00]. Measurement [GKMG02, ABF04, HS01, KHL01, KM04, KT05, MB00a]. measures [BMB02, JDM01, KM04, LS03b]. Measuring [HL03, KT01, SRK07, SKR08, PU01]. mechanism [MS00b, MY04]. Mechanisms [CCR07, vMW06, BM00]. Media [DLM08a, ZC09, MNH01]. mediated [VR03]. Membership [Ano08p, Ano09r]. memory [CMSB05, MP00, ZS01]. Merging [DMJN08, GZ05, Men02]. Message [Kra06a, SC06, AEY03, Gor00, MRV02]. messages [RP00]. Meta [SLLL07a, CDFG06]. meta-CASE [CDFG06]. Meta-Model [SLLL07a]. Metamodel [BLHS^{+}09, SLLL07b, ZME06]. Metaphor [AG06]. Method [TC09, BS00, Dol00, ENDK02, HS02, JS01, JMO04, LKV05, MS00b, MG00, MG02, TFD03]. methodological [AAG^{+}04]. methodologies [RHD02]. methodology [BF00, BZCJ02, BSRR02, MS00a, NPGM00, RGFR98, RGRF01, YA02, YSW^{+}01].

Methods [HALM08, PD07, PGM^{+}07, WW00, BT03, BDB00, Bro01, DN02, EFYB04, FDR04, GL00a, HDS06, JTMZ01, LS03a, MSO01, PD00, SC02, SC03]. Metrics [KDM09, OEGQ07, SRK07, SKR08, VDDR07, Woh01, ZL06, ZLX09, dAFS09, AL03, BVT03, CKB04, EBGR01, Eva03, FN01, GML05, GMK05, GR01, GFS05, LS03a, SJJ06, SK03, ZX02]. Metrics-Based [VDDR07]. Metrics-Guided [KDM09]. microkernel [ZS01]. microprocessor [LKW00]. Middleware [CC08, BCMS03, CEM03, CC03, MMRM05, Rav02, VDGD05, WCL04, ZBN^{+}04, ZC06]. Miner [LLMZ06]. Mining [Ano04-29, BJ09, CPY08, HMHJ05, MGF07, MDDG07, SYFP08, ZZ07, ZZWD05, SSCM06, WH05, YMNCC04]. mismatch [DeL01a, DeL01b]. Mismatching [CPS08]. missing [MSO01]. Mixed [DLM08a]. MMRE [FSKM03]. MNav [ZLW07]. MobiGATE [ZC06]. Mobile [CC08, SMBZ07, ZME06, CLZ02, CCW03, CEM03, CC03, JR06, LL03, MRV02, NPD01, Sat03, ZC06]. mobility [XYDD03]. MobiPADS [CC03]. Mobius [DCC^{+}02]. Model [BHHK03, BCH^{+}07, BDIS04, BCY01, CPS08, CGPA^{+}05, CM09, DHS09, EA08, HCC^{+}08, HKB09, HFMN09, HB07, HB08, MLLF09, NPPW09, SWE09, SD06a, SLLL07a, SCDP07, WS07, ZWL08, ZLW07, BD02, BKO05, CDM02, CDM03, CAB^{+}01, CW02b, FSKM03, GGV04, GCD03, HRRW01, Hol97, HS02, JWB02, LH03, Lee04, LKW00, MP03, Mil02, MT04b, NAD03, PD00, PSR05, RM02, RP00, San03, TYB06, TGK^{+}04, TA04, Tur02, WS06b, WCJ02, XYDD03, XS03]. Model-Based [CPS08, CM09, SCDP07, ZLW07, BDIS04, CGPA^{+}05, NAD03, TA04, WS06b, XS03]. Model-checking [BHHK03]. Model-Driven [EA08, MLLF09].

Modeling

[ACS09, BHSLB03, BDHK06, Bot05, BW01, DXLN07, DAP01b, FAOW^{+}09, KHL01, KPLJ03, Kou06, LPSS00, THY03, UBKW05, AAG^{+}04, AW04, BPT00, CW02b, EW05, HCC03, KT01, MG00, Mil02, SG00, XN06]. Models [CGV09, Cas09, CPR07, KT09, LW07a, LBMP08, MBFSV07, MM08b, Min06, PJJ^{+}07, RSSS02, RCW08, UBC09, BW02, BEFL00, BMW02, CGH02, CdPL04, DLDvL05, DH00, EL01, GHR01, HLK03, JW01, MOJ05, MPS02, MNS01, MSS05, RJ01, RHD02, UKM03, VDGD05, WGC02]. MODEST [BDHK06]. modified [JH03]. Modular [CCG^{+}04, HB08, KLZR06, RM02, SHFJ08, PMM00]. Modularization [SRK07, SKR08, MM06]. module [JS01, Ton01]. Modules [ALS08, KZEL09, KT05]. Monitoring [YPK^{+}07, DGR04]. monitors [PP02]. Moral [SGMB03]. motivated [SJLY00]. Move [TC09]. moving [HCC03]. mu [San03]. mu-calculus [San03]. multi [FDR04, YBT01]. multi-invariant [YBT01]. multi-methods [FDR04]. Multiclass [Cas09]. multicomputers [BF00]. Multicore [HB07]. multidevice [MPS04]. multifactor [BR03]. multilayer [RP00]. Multilayered [GA08]. Multilegged [LW07b]. Multilevel [AGK09]. multilinguistic [KKI02]. multimedia [HCC03]. Multiphase [Gou02]. multiple [KM04, MPS04]. multiplicity [PD00]. multiprocessor [BZCJ02, JWB02, YBT01]. multiprocessors [KT00, SSE01]. multithreaded [SP05, WS06a]. Mutation [AE09, ABLN06, DMM01, DR06b].

Navigability [ZLW07]. NDT [EA08]. Need [DLM08a]. needs [ACDD04]. Neglected [CPY08]. negotiation [LL03]. Negotiations [DLM08a]. net [BFGP03, HCC03, TFD03]. net-level [TFD03]. Nets [JTMZ01, Kou06, BBS02b, BSV05, CGH02, GH02, HT02, TFD03, XHD02, XN06]. Network [HFMN09, MM08b, BGK^{+}02, Neu02, RP00]. Networks [FAOW^{+}09]. neural [Neu02]. next [DPP00]. nominal [BH03]. Non [CGV09, BPT00, HT02]. Non-Markovian [CGV09, BPT00, HT02]. nonatomic [FFH04]. noncircular [Yan02]. nondeterministic [HCC04a, HCC04b, ZyC03]. Nonfunctional [CdPL04, GS05]. nonhomogeneous [HLK03]. Nonparametric [WS07, LS03a]. nonpreemptive [MMG03]. Notations [Moo09, NAD03]. note [Pik06]. Novel [LBMP08]. Number [MJJ^{+}07, OWB05].

O [CW00, RSSS02]. obfuscated [LKV05]. obfuscation [CT02]. Object [BCH^{+}06, CMPS07, DR09, DMJN08, KCC04, MD00, MPF08, OEGQ07, SKR08, SR06, ZL06, ZLX09, dLM08b, AL03, ABF04, AS04, BVT03, BD02, Bol00, BDB00, BBD01, BW01, BMW02, BG00, CS00, CN00, Cha03, CJ03, CW02b, CFTV05, DH00, DRW03, EBGR01, EBG^{+}02, Eva03, FN01, GR01, GFS05, HK02, IR01, JW01, Mil02, SJJ06, SLB00, SSE01, SP03, SK03, TCS05, XS05, ZX02]. Object-Oriented [BCH^{+}06, CMPS07, DR09, DMJN08, MPF08, OEGQ07, SKR08, ZL06, ZLX09, dLM08b, AL03, ABF04, AS04, BVT03, BD02, BBD01, BW01, BMW02, BG00, CS00, CN00, Cha03, CW02b, CFTV05, DH00, DRW03, EBGR01, EBG^{+}02, Eva03, FN01, GR01, GFS05, IR01, SLB00, SP03, SK03, TCS05, XS05, ZX02]. objects [ACL02, CFP^{+}03, HCC03, dBFT00]. observable [ZyC03]. obstacles [vLL00]. Off [LCS^{+}08]. Off-the-Shelf [LCS^{+}08]. one [SM03a]. one-way [SM03a]. online [CW02a, KCLL03]. only [SLLL07b]. onto [MP00]. OO [TYB06]. Open [CFM08, BDB00, DTB05, GFS05, KT05, PSE04]. open-source [KT05, PSE04]. Operating [DY08b, KD00, LRSF04]. operational [BMB02, RM02]. operations [LM04]. Operators [AE09]. opinion [FBV05, LS03b]. Opportunities [TC09, ED05]. OPS5 [KC04]. OPSS [CDF01]. Optimal [ATG^{+}06, YHJ08, ZyC03, ZWL08, EBG^{+}02]. Optimization [Ano08i, BG09, CH06]. Optimized [NTR09]. Optimizing [CAB^{+}01, LC04]. Optimum [JS02]. Order [WS07, BLW03, Mit05]. Order-Statistic [WS07]. Organizations [YCCX09]. Oriented [BCH^{+}06, Cas09, CMPS07, DR09, DMJN08, DP09, HRJ08, MPF08, OEGQ07, SKR08, ZL06, ZLX09, dLM08b, AL03, AAG^{+}04, ABF04, AS04, BVT03, BD02, Bol00, BBD01, BW01, BMW02, BG00, CS00, CN00, Cha03, CW02b, CFTV05, DH00, DRW03, EBGR01, EBG^{+}02, Eva03, FN01, GR01, GFS05, IR01, SJJ06, SLB00, SP03, SK03, TCS05, WBY05, XS05, XS03, XN06, ZX02, vLL00]. origin [GZ05]. Other [DC06]. overruns [MOJ05].

Package [HRJ08]. packaging

[DeL01a, DeL01b]. Pair [AGDS07]. Pairs [LCN08]. pairwise [TL02]. paleontology [AP03]. Papers [Ano04-28, Ano04-29, Ano05-27, Ano07q, Ano08i, Ano09i, Ano09j, Ano09k, BDB00, DAP01a, OJ02]. paradigm [BAvZ00]. Parallaxis [Bra00]. Parallaxis-III [Bra00]. Parallel [ABB^{+}00, BW05, Bra00, CCL^{+}00, CW00, FRC^{+}02, GL00a, MP00, NPGM00, PDB00, RR00, RSSS02, SZ02, TSJ00b, TSJ00a]. parallelization [GL00b]. parallelizing [KT00]. parameters [BH04]. Parametric [BFGP03, Wan05]. Parse [KDM09]. Part [BHSLB03, THY03]. Partial [UBC09, Mit05]. Participation [DM08]. partition [BSC03, CPT03, Nta01]. partitionable [BDM01]. passing [Gor00]. paste [LLMZ06]. Pattern [GA08, TCSH06, FKGS04, PULPT02]. Patterns [DYZ07, DHQ^{+}08, CCF^{+}00, KCC04, PUT^{+}01, SHP01, Vok04]. Payoffs [DC06]. PEPA [GHR01]. Perceived [LG03]. Perceptions [NMJ09]. Performance [ABB^{+}00, Cas09, CPR07, Kou06, NPD01, YPK^{+}07, ZWL08, AW04, BZCJ02, BDIS04, BVT03, BR03, BSV05, CW00, CCW00a, CCW00b, CGH02, CGPA^{+}05, LGF05, MG00, Nix00, PSJ00, SM03b, VDGD05, WV00, WIK^{+}04]. performed [EH01]. personal [PU01]. perspective [BLD02, LEH01]. Petri [BBS02b, BFGP03, BSV05, CGH02, GH02, HT02, HCC03, JTMZ01, Kou06, TFD03, XHD02, XN06]. Petri-net [HCC03]. phase [SRC05]. Physics [Moo09]. Planning [CH06, NTR09, MPS01]. plans [RM02]. Platform [CPR07]. platforms [DS02]. playground [MR00]. Pluggable [KLZR06]. Plumblines [CHMB08]. POEMS [ABB^{+}00]. point [CFTV05]. pointer [WSY^{+}04]. points [BC03, CRL01, MACE05]. points-to [BC03, CRL01, MACE05]. Poisson [HLK03]. policies [BPT00, TM05]. policy [BBG06, MC02, vdHCHW02]. Polymetric [LD03]. Polymorphism [RMR04]. Population [vdMR08]. populations [vO05]. portability [SP05]. portable [CW00]. Positive [HOM08]. POSIX [KD00, LRSF04, SP05]. POSIX-compliant [SP05]. possible [BW02]. Potentially [ZLX09]. Power [CMPS07, Hat09, MS00a]. Power-Law [Hat09]. Power-Laws [CMPS07]. Practical [DM06, BS00]. Practice [LCS^{+}08]. Practices [MCHL06, RDKN03]. Precision [MDDG07, SA02]. precondition [CN00]. predicate [AV01, BH04]. predicates [CW02a]. Predict [OEGQ07]. Predicting [BVT03, GKMS00, HFMN09, OWB05, SC01, TCS05, YMNCC04, ZL06, PSR05]. Prediction [LBMP08, MPF08, MM08b, BDIS04, FN01, GFS05, HRRW01, LGF05, MSS05, SK01, SSCM06, Tia02]. Predictor [MDDG07]. Predictors [MGF07, ZZ07]. preemption [BPT00, LLH^{+}01]. preemptive [BFSV04, BSV05]. Preliminary [KPP^{+}02]. preprocessed [Spi03]. preprocessor [EBN02]. Prescriptions [NMJ09]. Presence [CDS08, DMJN08, CRL01]. Preserving [MBFSV07]. Pressure [NH09]. PRI [HT02]. Price [MJJ^{+}07, TW07]. primitives [MT04b]. principal [RDKN03]. Principle [DXLN07]. priorities [TFD03]. Prioritization [DR06b, LHH07, EMR02, JH03]. Prioritizing [RUCH01]. Priority [GL00a]. Privacy [BA08, SC09]. Privately [WM08]. Probabilistic [HKB09, NPPW09, PGM^{+}07, PSR05]. probability [TCS05]. Problem [HRJ08, PD00, SBMJ01]. Problems [MDDG07, LI00]. procedure [BD03]. Process [DM08, DP09, NMJ09, PYM^{+}07, YPK^{+}07, BMB02, CDM02, CDM03, Dyb05, EB00, HLK03, JS02, JKS04, LRS00, MP03, OBMP03, OFdL^{+}04, PU01, VR03, WBY05]. process-control [WBY05]. Process-Oriented [DP09]. Processes [AP07, DC06, OEGQ07, ZZZ07, AL03, DMCS05, MB00b, NPD01]. processing [BW05, Bra00, JWB02, TSJ00b, TSJ00a]. processors [BLD02]. Product [BBS02b, MYC05, vO05]. production [KC04]. Productivity [DC06, Kao03, KM04, MRS02]. products [KT05, PSE04]. profile [ACLdSS04]. Profiling [ED05]. Program [CZvD^{+}09, LCN08, MACE05, SLLL07a, SLLL07b, CFC01, DHOH03, ECGN01, GL00b, PULPT02, SJLY00, SGMB03, Ton03, WCD^{+}05]. Programmer [AGDS07]. Programming [AGDS07, BHM06, SMD08, BF00, BCRW00, CCL^{+}00, EMT05, SZ02, XC03, vdHCHW02].

Programs [DR09, RCR06, dLM08b, vdMR08, BW05, CN00, GKMG02, GMK05, Gor00, IR01, KCT02, LC06, PDB00, SP05, SH00, WS06a, XC03, ZD04]. Project [HFMN09, ACDD04, ATG^{+}06, Ber00, CMSB05, DTB05, FBV05, JC04, MOJ05, RL00]. Projects [AC07, CFM08, BMW02, SM03b]. Proneness [OEGQ07, PD07, ZLX09, BMW02]. proof [KHCP00, Lit00]. proofing [CT02]. Proper [MJJ^{+}07]. Properties [DR09, DHS09, HVC09, KW07, UBC09, BKO05, BW01, SA03]. Property [GR01, ZX02, SJJ06]. proportional [Nta01]. Proposed [LBMP08]. Protecting [HOM08]. Protection [DVJP08, CT02]. protocol [Cam01, HL03, LH03]. protocols [DMCS05, PV02]. prototype [SMBK03]. prototyping [JWB02, KL02]. Provable [DVJP08]. Providing [KKJ08]. Proximity [LZH08]. PSP [PU01, KP09]. publish [CCW03]. publish/subscribe [CCW03]. purity [FFQ05].

QEST [FKW06]. QoS [CC08, LPRT07, ZBN^{+}04]. QoS-Aware [LPRT07, ZBN^{+}04]. qualitative [BW02]. Qualities [CHMB08]. Quality [AC07, CFM08, DC06, Kao03, KP09, MRS02, PYM^{+}07, SRK07, SKR08, YPK^{+}07, BD02, BBD01, Cha03, PL03]. Quantitative [AR07, FO00, FKW06, HKT09, VSC09, LRS00]. quasi [LEH01]. quasi-experimental [LEH01]. query [GCD03]. Questions [SMD08]. queue [DS02]. Queueing [FAOW^{+}09, Kou06, RP00]. queues [GL00a]. queuing [ACDD04]. Quiescence [VEBD07].

race [Mit05]. Radar [DLL09]. radial [SG00]. railway [HP00]. random [BSC03, Nta01]. Ranking [IYY^{+}05, PGM^{+}07, LS03b]. Rapid [TCKO02, JR06, KL02]. rapidly [MX05]. rational [MP03]. Reachability [LC06, IR01]. Reactive [VSC09, CJMR07]. reading [LEH01, SLB00, TRW03]. Real [BNR06, FMPR06, ACLdSS04, BKO05, BFSV04, BSV05, ENDK02, HLT^{+}04, JTMZ01, KLK00, KCLL03, LLH^{+}01, LC04, LRSF04, LZBQ04, MN01, MMG03, PD00, PP02, Rav02, SA03, XHD02, ZS01]. Real-Time [BNR06, FMPR06, ACLdSS04, BKO05, BFSV04, BSV05, ENDK02, HLT^{+}04, JTMZ01, KLK00, KCLL03, LLH^{+}01, LC04, LRSF04, LZBQ04, MN01, MMG03, PD00, PP02, Rav02, SA03, XHD02, ZS01]. Realism [Jor04]. Realistic [DAB08]. Really [ML09]. Reasoning [DR06a]. Reasons [Ano09b, JMO04]. recapture [BEFL00, EL01]. recency [SW02]. Recommend [HWM06]. Recomputing [CH09]. Reconfigurable [BLD02]. Reconstruction [DP09]. Recovering [ACC^{+}02]. Recovery [KDM09, MB07, MMG03]. recurrence [MP00]. rediscovered [LI00]. Reducing [PPW^{+}02]. Reduction [JG07, JTMZ01, MM08a, JH03, WSY^{+}04]. redundancies [XE03]. redundant [BFGP03]. reengineering [BCMV03]. Refactoring [BCH^{+}06, KDM09, TC05, TC09, MT04a]. Refactorings [DMJN08]. Reference [DR09, OBMP03, RJ01, Tur02]. References [SR06]. Refinding [SRS^{+}09]. Refinement [SWE09, BSR04, CN00]. reflective [CEM03, CC03, FDR04]. reflexion [MNS01]. Regression [CH09, LHH07, HRRW01, RUCH01, XN05]. Regulatory [BA08]. Relate [KMCA06]. Related [DVJP08, LLH^{+}01, LLMZ06]. relation [CPT03]. relational [BNL05]. relations [HCC03, IYY^{+}05, SGMB03]. Relationship [CFM08, KZEL09, BHSLB03, THY03]. Relationships [DC06]. Release [NTR09, YHJ08]. Relevant [HWM06, KMCA06]. Reliability [DXLN07, MSS05, RL01, WS07, FG02, GL05, HLK03, KT01, KBYC04, PSMK03, RM02, Tia02, TRL04, YA02]. Reliable [YPK^{+}07]. Reminding [SRS^{+}09]. remote [MR00]. Repair [DR06a]. replacement [TM05]. Replaying [ECDJ09]. Replicated [AR07, LEH01]. replication [DTB05, LRS00]. Reply [BNR06, YSW^{+}01]. Repositories [Ano04-29, HMHJ05, WH05]. repository [VZJ03]. Representation [HLMN08, DH00, JKS04, OFdL^{+}04]. Reproducibility [ASM09]. required [MP03]. Requirements [BA08, CNP08, DC06, DLM08a, EA08, HLMN08, NMJ09, PP02, CG01, CdPL04, EB00, GS05, HDS06, HW02b, JY06, MYC05, Nix00, RJ01, RSSS02, SRC05, Xu03, vLL00]. Requirements-based [PP02]. research [Ber00, FK05, KPP^{+}02, SJLY00]. resolution [Lee04]. Resolving [Mit05, VFdM^{+}03]. Resource [NTR09, BCMS03, MMRM05]. resource-constrained [MMRM05]. Respect [AGDS07]. Respectful [WO00]. respondent [JMO04]. Response [MDDG07, OAF^{+}04, RGRF01, SC03]. Restart [vMW06]. restructuring [Ton01]. result [AZ03]. Retaining [JG07]. Retargeting [BW05]. Retrieval [PGM^{+}07]. retrieving [VZJ03]. reuse [DPP00, FK05, MD03, MA04, MET02, MET03, PS04, RDKN03, Sel05, SBV01, TGK^{+}04, vO05]. reuse-based [Sel05]. Reverse [BLL06, vDS05, HW02b, LD03, SDA05]. Reversible [CFC01]. Review [HSD07, JS07, KMT07, TA04]. Reviewer [Ano08b]. Reviewers [Ano03b, Ano04a, Ano05a, Ano06b, Ano07d, Ano09a, Ano00, Ano01a]. Reviewing [HHC04]. Reviews [KP09, ML09, RL01, SJLY00]. revisited [Tur02]. Revisiting [SMBZ07]. revocation [SW02]. rhetoric [PP05]. Risk [BG09, DC06, LCS^{+}08, CGPA^{+}05, GPHG^{+}03, Neu02, RL00, YA02]. risks [KPLJ03]. RMI [SR06]. RMI-Based [SR06]. robust [YBT01]. Robustness [FMRW05]. Role [HJ08, MBGM09, AW04, JMO04, PL03]. Role-Based [MBGM09]. roles [CFP^{+}03]. Router [HRJ08]. RPG [SBV01]. rule [Cha00, LC04]. rule-based [LC04]. Rules [BA08, CCF^{+}00]. Running [SAG^{+}06]. Runtime [WS06a, DGR04]. RUP [MP03]. Rushby [Pik06].

Safe [SHFJ08, VEBD07, VFdM^{+}03]. Safety [LW07b, LM04, WSY^{+}04, Wan05]. safety-analysis [WSY^{+}04]. safety-critical [LM04]. Same [ASM09]. SARE [MP00]. SATIN [ZME06]. satisfaction [SBV01]. Saturation [Min06]. Scalable [MBGM09, SWE09]. Scale [BHM06, SKR08, KKI02, KT05, LLMZ06, PPW^{+}02, Sel05]. Scale-Up [BHM06]. Scaling [BSR04]. Scenario [BHM06, GS05, BKO05, Egy03]. Scenario-Based [BHM06, GS05]. scenario-driven [Egy03]. scenario-matching [BKO05]. Scenarios [PGM^{+}07, UBC09, DLDvL05, HW02b, Mit05, UKM03]. schedulability [KCLL03, XHD02]. Schedule [NH09]. scheduler [MMG03]. scheduling [LRSF04]. scheme [HLK03, VZJ03]. Schur [BSC03]. Scientific [Moo09, CW00]. SCL [HH06]. Scoring [TCSH06]. Screening [YPK^{+}07]. Search [Ano08i, LHH07, San03]. Search-Based [Ano08i]. Searching [BC03]. second [BDB00]. Section [Ano03c, BKM07, CTM09, DY08a, FGLW09, FKW06, GN06, MN08, UBKW05, CGH02]. Secure [BDFZ08, HALM08, MR00, MN08, XN06]. Security [BA08, HLMN08, LKW00]. see [BSRR02]. Seek [Ano08-28, KMCA06]. Selected [BDB00]. Selecting [MCHL06]. selection [HRRW01]. selective [San03]. Selectively [JG07]. Self [DY08b]. Self-Stabilizing [DY08b]. Semantics [BDFZ08, Bot05, MBFSV07, Mil07, BFGP03, CN00, EW05, NAD03, SW02]. Semantics-Based [BDFZ08]. Semantics-Preserving [MBFSV07]. sensitivity [CDM03]. Sequence [BLL06, Mit08, SC06, AEY03, BHS05, PP03]. sequence-based [PP03]. sequences [PMM00]. sequential [BW05, MOJ05, RL01]. series [BDL04]. Server [KMSL08, FMRW05, HCC04a, HCC04b, MNH01, MG00, RP00, TRL04]. Servers [LPRT07]. Service [Ano09j, Ano09k, AP07, HVC09, YCCX09, CCW03, GS01, LL03, NKL05]. Service-Based [Ano09j, Ano09k]. Services [Ano09j, Ano09k, BDFZ08, AP03, FBS05, OBMP03, ZBN^{+}04, ZC06]. Session [DVJP08, KRM06, SSG^{+}07, ERKF05]. Session-Based [KRM06]. Sessions [JG09]. set [BLD02]. sets [MB03, MACE05, MSO01]. setting [VMA^{+}04]. Seventh [Woh01]. several [LS03a]. Severity [ZL06]. SEXTANT [SEHM06]. Shallow [SRC05]. shared [YBT01]. Shelf [LCS^{+}08]. Shortening [KC04]. side [DHOH03]. SIFT [WIK^{+}04]. signal [JWB02]. significance [IYY^{+}05]. SIGSOFT [DY08a]. Similarity [TCSH06]. simpler [PUT^{+}01]. Simplifying [ZH02]. Simulation [RCW08, SBMJ01, ACDD04, FSKM03, FRC^{+}02, GL05, GL00a, PDB00, SK01]. Simulation-Based [RCW08]. Simulation-verification [SBMJ01]. simulations [BGK^{+}02]. Site [ZLW07]. Size [Hat09, KZEL09, ZLX09, CFTV05, Dol00, EBGR01, EBG^{+}02, Eva03, HS01, KM04, LH03]. Size-Defect [KZEL09]. Skoll [PYM^{+}07]. SLA [LPRT07]. SLA-Driven [LPRT07]. slices [Ton03]. slicing [MACE05]. Sloan [Ber00]. Small [Bot05, PP05, ZS01]. small-memory [ZS01]. Smells [VDDR07]. smooth [Tur02]. Snapshots [KW07]. Society [Ano08-28, Ano04-39, Ano04-41, Ano04-44, Ano04-42, Ano04-40, Ano04-43, Ano06-28, Ano06-30, Ano06-29, Ano06-45, Ano06-44, Ano07a, Ano07b, Ano07-46, Ano07-47, Ano07-48, Ano07-45, Ano07-49, Ano08p, Ano08q, Ano08r, Ano09b, Ano09r]. Softly [BDHK06]. Software [AC07, ASM09, AR07, Ano03c, Ano04-29, Ano04x, Ano04w, Ano04y, Ano04-27, Ano04v, Ano04z, Ano07q, Ano08i, Ano09j, Ano09k, BJ09, BKM07, BLL06, CTM09, CFM08, CMRH09, CPY08, CH06, CMPS07, CYK^{+}05, CPR07, DXLN07, DY08a, DMJN08, DT04, DP09, DM06, DAB08, ER08, FGLW09, FK05, GHM08, Gri03, GN06, GR06, HRJ08, HGP09, HSD07, HJ08, HKL^{+}07, HKL05, HB02, HMHJ05, Hat09, HALM08, HB08, JKS04, JS07, KP09, KKJ08, KWZ08, KM04, KMCA06, KZEL09, KWG04, LW07a, LBJ06, LBMP08, LCS^{+}08, LW07b, MBFSV07, MB07, MLLF09, MN08, Moo09, MNS01, MJJ^{+}07, NH09, NTR09, OEGQ07, OFdL^{+}04, PD07, PJJ^{+}07, Rei06, Rot05, RMR04, SRK07, SKR08, SEHM06, SR06, SSCM06, SRS^{+}09, SEM01, TW07, VJB09, WS07, Woh01, YHJ08, YPK^{+}07, Zha08, vMW06, vO05, vdMR08]. software [AL03, ART03, AT05, ACDD04, ABF04, AS04, ABHL06, ABGM02, BZCJ02, BDIS04, BLD02, BT03, BNL05, Bou00, BW02, BEFL00, BMW02, Bro01, BP04, CA03, CDM02, CDM03, CS00, CCG^{+}04, Cha00, CCW00a, CCW00b, CT02, CMSB05, DKST05, DGR04, DPP00, DN02, Dol00, DIH00, Dyb05, EGK^{+}02, EB00, EBG^{+}02, ED05, FO00, GL05, GKMG02, GMK05, GKMS00, GFS05, HS03, HS01, HM03, HS02, HLT^{+}04, HLK03, HHC04, IYY^{+}05, JS02, JC04, JMO04, Jor05, KT00, Kel06, KBYC04, KHL01, KPP^{+}02, KPLJ03, KGR01, LI00, LS03a, LS03b, LLMZ06, LPSS00, MT00, MX05, Men02, MT04a, MD03, MMS01, MY04, MM06, MOJ05, MC02, MET02, MET03, MO03, MIB04, MSS05, Neu02, OWB05, PL03, PS04, PSE04, PSR05, PPW^{+}02, PV02, PU01, PP03, RM02, RL01]. software [RHD02, RL00, RDKN03, Sat03, SJLY00, Sel05, SK01, SG00, SP05, SV02, SHH^{+}05, SHP01, SC02, SC03, SK03, SBV01, TSJ00b, TSJ00a, TM05, TCKO02, TRL04, TGK^{+}04, Tur02, VDGD05, VR03, VMA^{+}04, WSY^{+}04, WS06b, WV00, WGC02, XY03, XS05, XS03, Xu03, XN06, ZWN^{+}06, ZZWD05]. Software-Based [LW07b]. software-fault [DGR04]. Solution [FAOW^{+}09, BBS02b]. solutions [PUT^{+}01, SA02]. Solving [HRJ08]. some [HLK03]. sorting [GL00a]. Source [CFM08, FWPG07, BEK^{+}04, DTB05, EKS03, GZ05, GFS05, HCM02a, HCM02b, HH06, KKI02, KT05, PSE04, PP05, RCM04, WH05, YMNCC04]. Space [dLM08b, BFSV04, HLP01, KLK00]. space-craft [HLP01]. spaces [YCP06]. spanning [MB03]. sparse [SC01]. SPARTACAS [MA04]. Spatial [GML05, HCC03]. spatial-temporal [HCC03]. Special [Ano03c, Ano04-29, Ano07q, Ano09j, Ano09k, BKM07, CTM09, DY08a, FGLW09, FKW06, GN06, GR06, HKL05, HMHJ05, HKT09, MN08, UBKW05, WW00, CGH02, TSJ00b, TSJ00a]. Specific [ACS09, CM09, KBYC04]. Specification [BNR06, DR06a, FMPR06, SYFP08, AH00, BDM01, BS00, BG00, FKGS04, FRC^{+}02, HL03, JS01, LH03, MN01, PD00, PP03]. Specification-Based [DR06a]. Specifications [PIM09, WM08, JDM01, Lev00, SD06b]. specify [HH06, TFD03]. Specifying [HVC09, MKL^{+}04]. spectra [XN05]. speed [HM03]. SPIN [BCY01, HLP01, Hol97, HB07]. splitting [GZ05]. SPNs [BPT00]. spreadsheet [BAvZ00, BSRR02, RBR06]. Spreadsheets [AE09]. spurious [SA02]. stability [Kel06]. stabilization [Gou02]. Stabilizing [DY08b]. Stable [KW07]. Stack [DS02, MM08a]. Staff [Ano04-44, Ano04-43, Ano04y]. staffing [ACDD04]. State [Ano05-29, APS07, BCH^{+}07, CGV09, Kni06, Kra07, Kra08c, Kra09b, LCS^{+}08, UBKW05, VSC09, dLM08b, AV01, BHS05, BDL04, BFSV04, CDM03, KLK00, Kni04d, Men02, SBMJ01, SD06b, ZyC03]. State-Based [UBKW05, BDL04, SD06b]. State-Density [CGV09]. state-of-the-art [Men02]. State-of-the-Practice [LCS^{+}08]. State-Space [dLM08b]. statecharts [CAB^{+}01]. STATEMATE [BHH^{+}09]. Static [KMS04, MGF07, MDDG07, SR06, SYFP08, Vic01, ZZ07, ZWN^{+}06]. station [MS00a]. Statistic [WS07]. Statistical [KKJ08, LFY^{+}06, JS02, KT01]. status [FK05]. steering [Vic01]. step [BSR04]. step-wise [BSR04]. Stochastic [DHS09, NPPW09, VSC09, BBS02b, BSV05, GH02, HT02, TFD03]. Stock [TW07]. stopping [Cha00]. storing [VZJ03]. Strategies [RDKN03, BLW03, ED05]. strategy [TL02]. streaming [WCL04]. Stress [KRM06]. strong [CW02a]. STRPN [HCC03]. Structural [HWM06, DKST05, GR01, ZX02]. Structure [DR06a, KLZR06, DS03, GL05, HCM02a, HCM02b, ZD04]. structure-based [GL05]. structured [BGM04, TA04]. structures [WSY^{+}04, Wan05, YBT01]. student [XC03]. Studies [JS07, KMT07, EMR02, HRRW01, MSS05, SV02]. Study [AC07, ASM09, CFM08, CNP08, DC06, DM08, DM06, KP09, KMCA06, LW07b, LZH08, MPL07, MLLF09, MM08b, NMJ09, RAW08, VJB09, YHJ08, ABGM02, BS00, DTB05, FSKM03, HDS06, HM03, HHC04, JC04, Kao03, Kel06, MO00, MRS02, PSE04, RCM04, SP05, SHP01, TGK^{+}04, Vok04, WV00, XY03]. Studying [MX05]. style [AS04, MMRM05]. style-aware [MMRM05]. SubCM [VMA^{+}04]. Subject [Ano01c, Ano02c, Ano02d, Ano03d]. subscribe [CCW03]. Success [MET02, Dyb05, GKMG02, MD03, MET03]. Suite [JG07, MM08a, JH03]. Suites [CDS08, OEGQ07, RCW08]. Supertotal [Bou00]. Support [EW06, HKL^{+}07, KMSL08, SRS^{+}09, XS07, CCW03, ERKF05, ECGN01, EW04, LHS03, MPS02, VR03]. Supported [BCH^{+}06, ACLdSS04]. supporting [BZCJ02, CPT03, DS03, DL05, SW02]. Survey [CZvD^{+}09, LCS^{+}08, BDIS04, DN02, Men02, MT04a, RL00, SHH^{+}05]. Symbolic [Wan05, CAB^{+}01]. symmetry [WSY^{+}04]. Symposium [ER08, HB02, Rot05, Woh01]. symptoms [LI00]. Synchronizability [FBS05]. synchronization [SW02]. Synchronous [MBFSV07, RP00]. Syntax [Bot05, HOM08, EW05]. Syntax-Aware [HOM08]. Synthesis [PJJ^{+}07, UKM03, UBC09, MS00a, SD06b]. Synthetic [KRM06]. System [ASM09, AGDS07, BNR06, CH06, CMPS07, ECDJ09, FMPR06, HALM08, KT09, dAFS09, AP03, BR03, CEM03, CS00, FO00, HP00, HW02b, KKI02, KL02, MP00, MN01, MC02, MPS02, PSJ00, RL01, Rav02, RGFR98, SA03, WCD^{+}05, YSW^{+}01, ZME06]. Systematic [CZvD^{+}09, HSD07, JS07, KBYC04, KMT07, LZH08, Pik06, Rus99, SCDP07, TYB06]. Systemic [CHMB08]. SysTems [FKW06, ABB^{+}00, AR07, Ano09j, Ano09k, BG09, BDHK06, CDS08, DY08b, Hat09, HKT09, Kou06, KRM06, LW07b, MPF08, MBGM09, MN08, PMT^{+}08, RCW08, SAG^{+}06, VSC09, WHY06, YPK^{+}07, ZME06, dLAF06, vMW06, BDM01, BZCJ02, BCMV03, BFGP03, BDB00, BW01, BFSV04, BSV05, BG00, Cha00, CW00, CJ03, CJMR07, CFTV05, ENDK02, FN01, GGV04, HA00, HK02, HS02, JWB02, JTMZ01, KT00, KLK00, KC04, Kho02, KCC04, KD00, KCLL03, LLH^{+}01, LC04, LRSF04, LZBQ04, MMRM05, MP00, MG00, MT04b, MG02, MM06, MB00b, Nix00, OWB05, PP02, PMM00, PSMK03, Rav02, RGRF01, Sel05, SP05, SP03, TCS05, Tur02, Vic01, WBY05, Wan05, WV00, XHD02, XC03, YBT01, WW00].

Table [Ano04-27, Ano04z]. tackle [AV01]. Tagging [SRS^{+}09]. Tainting [HOM08]. tale [BT03, SC02, SC03]. tamper [CT02]. tamper-proofing [CT02]. Task [SMD08, WS06b, MPS02, SHP01]. Tasks [HKL^{+}07, KMCA06, LCN08]. Taxonomy [DP09, DGR04]. TCTL [WHY06]. team [TCKO02]. teams [BH03, MY04]. technical [SJLY00]. Technically [Ano07-32, Ano08u, Ano08s, Ano08t]. Technique [KRM06, FKGS04, GH02, IR01, Neu02]. Techniques [AV01, DR06b, HKL^{+}07, MPL07, VJB09, DRW03, LPSS00, RBR06, SK01, SLB00, WH05]. Technology [HCC04a, HCC04b]. telecommunication [PSJ00]. Template [NAD03]. Temporal [GCD03, HVC09, HCC03, Kho02]. Terms [Ano08-28]. Test [CH06, CDS08, DR06b, EMR02, ECDJ09, JG07, JH03, LHH07, MBGM09, MPL07, MM08a, PJJ^{+}07, RCW08, VDDR07, AZ03, BLW03, BSRR02, CDM02, CDM03, CPT03, DKST05, DS00, EMT05, HRRW01, MPS01, MX05, MMS01, NFLJ06, RM02, RUCH01, TL02]. test-first [EMT05]. Test-suite [JH03]. Testability [HHH^{+}04]. testbed [vdHCHW02]. Testing [ABLN06, BSRR02, CH09, ER08, HB02, KRM06, LFY^{+}06, Rot05, RMR04, SSG^{+}07, SCDP07, VJB09, AW04, BSC03, BDL04, CJMR07, CYK^{+}05, DMM01, EFYB04, ERKF05, ED05, ENDK02, FMRW05, HCC04a, HCC04b, KT01, Kho02, KHCP00, KCT02, KWG04, LC06, LHS03, MB03, MIB04, Nta01, PBG04, RUCH01, Sat03, SH00, TL02, WV00, WGC02, XN05, ZyC03]. Testing-Based [LFY^{+}06]. tests [KCLL03]. Their [CMRH09, DYZ07]. them [RL00]. Theoretic [SRK07, AT05]. theoretical [RHD02, VZJ03]. Theory [HSD07, WHY06, ACL02, DH00]. theory-based [DH00]. Threat [XN06]. Threat-driven [XN06]. Three [OEGQ07, DRW03, FG02]. three-tier [FG02]. tier [BF00, FG02]. Time [AC07, BNR06, FRC^{+}02, FMPR06, HT02, NH09, VSC09, WHY06, YHJ08, ACLdSS04, BHHK03, BKO05, BFSV04, BSV05, EW06, ENDK02, HLT^{+}04, JTMZ01, KLK00, KC04, KCLL03, LLH^{+}01, LC04, LRSF04, LZBQ04, MN01, MMG03, PD00, PP02, PMM00, Pik06, Rav02, Rus99, SA03, Vic01, XHD02, ZS01]. time-dependent [Vic01]. time-triggered [Pik06, Rus99]. Timed [BDHK06, BFSV04, DHS09, DHQ^{+}08, ENDK02, MD00]. timing [HW02a, MKL^{+}04, Xu03]. token [KKI02]. token-based [KKI02]. tolerant [Pik06, Rus99]. Tool [APS07, BCH^{+}06, EW04, LHS03, SEHM06, ART03, BBS02a, EH01, GCD03, MM06, SMBK03, SA03, VMA^{+}04]. Tool-Equipped [APS07]. Tool-Supported [BCH^{+}06]. Toolkit [BGM04, ACLdSS04]. Tools [BKA^{+}07, ABGM02, CT02, DGR04, TSJ00b, TSJ00a]. tour [BHS05]. Trace [KT09, Egy03, JS01]. Traceability [CHMB08, ACC^{+}02, CHCC03, RJ01]. tracing [HDS06]. Tracking [MRV02, ZWL08]. training [PU01]. Tranquility [VEBD07]. Transactions [Ano04-29, Ano04x, Ano04w, Ano04y, Ano04-27, Ano04v, Ano04z, Ano09i, FG02]. transfer [ZyC03]. Transformation [XC03, Big04, HHH^{+}04, RR00, TYB06]. Transformation-based [XC03, Big04]. Transformations [KMS04, Mil02, Spi03]. transformer [MS00a]. transient [GH02, MMG03]. Transition [BHM06]. transitions [HT02]. transport [ZC06].

Tree [FWPG07, BFGP03]. Trees [KDM09, SM03a, ZyC03]. Trends [HGP09, PRT00b, PRT00a]. Triggered [SC06, Pik06, Rus99]. TSE [Ano04-56, Ano04-55, Ano04-59, Ano04-57, Ano04-60, Ano04-63, Ano04-61, Ano04-64, Ano04-62, Ano04-58, Ano05-61, Ano05-62, Ano05-56, Ano05-65, Ano05-55, Ano05-54, Ano05-60, Ano05-58, Ano05-63, Ano05-57, Ano05-59, Ano05-64, Ano06-56, Ano06-53, Ano06-52, Ano06-50, Ano06-55, Ano06-57, Ano06-49, Ano06-54, Ano06-48, Ano06-46, Ano06-47, Ano06-51, Ano07-53, Ano07-52, Ano07-54, Ano07-58, Ano07-60, Ano07-57, Ano07-51, Ano07-61, Ano07-59, Ano07-56, Ano07-50, Ano07-55, Ano08-33, Ano08-32, Ano08-30, Ano08-29, Ano08-31, Ano08-34, Ano09-28, Ano09-30, Ano09-27, Ano09y, Ano09-29, Ano09z, Kni04d, Kni06]. tune [EH01]. tuning [CW00, SA03]. TURTLE [ACLdSS04]. tutoring [XC03]. Two [Ano06-30, PULPT02, AL03, Ano06-28, Ano06-29, EL01, KT05]. Type [SHFJ08, DHOH03, WO00]. Type-Safe [SHFJ08]. Types [ML09].

ultra [WBY05]. ultra-high [WBY05]. UML [ACLdSS04, ABHL06, BLDYB05, BLL06, DAB08, EW04, FKGS04, GPHG^{+}03, Mil02, Mil07, PJJ^{+}07, VDGD05]. UML-based [BLDYB05, FKGS04]. Uncertainty [DXLN07, JG09, YHJ08, Jor04, Jor05]. understanding [DL05, PP05, SRC05, Ton03]. Unified [BHSLB03, THY03, HLK03, MP03, MB00b]. uniform [TM05, WMC01]. Unit [ECDJ09, VJB09]. units [MRV02]. Unpredication [SDA05]. unscheduling [SDA05]. unspeculation [SDA05]. Updates [VEBD07]. Usability [JMSS07, ZZZ07]. usage [KT01, TRW03]. usage-based [TRW03]. Use [DM08, DR06b, HSD07, LW07b, SRS^{+}09, BvDvET05, EBN02, IYY^{+}05, Lit00, NFLJ06, SP03]. usefulness [PULPT02]. User [LL03, SSG^{+}07, SMBK03, AP03, DLDvL05, ERKF05, MPS04, WIK^{+}04]. User-centric [LL03]. user-session [ERKF05]. User-Session-Based [SSG^{+}07]. user-visible [AP03]. users [WCD^{+}05]. Using [ABLN06, ABGM02, CDM03, CCF^{+}00, CHMB08, GZ05, HOM08, HFMN09, HH06, Kou06, KW07, MPF08, MB03, MIB04, OEGQ07, PRS04, PD07, PGM^{+}07, RCW08, SYFP08, Ton03, TCSH06, VSC09, XE03, ZD04, ZWL08, BVT03, BR03, BW01, BSV05, CFC01, GPHG^{+}03, GKMS00, HLP01, JTMZ01, Kel06, KM04, LI00, MR00, MS00a, MS00b, MPS01, Mil02, MM06, SK01, SM03a, SG00, SHP01, WSY^{+}04, XHD02, XN06, ZZZ07]. utility [GML05].

validate [CDM03]. Validating [EB00, HVC09]. Validation [MB00a, OEGQ07, AL03, ACLdSS04, Cha00, Dol00, GFS05]. validity [EBGR01, Eva03, MSS05]. value [XN05, ZWN^{+}06]. values [KT05]. Variability [ASM09, MYC05, WCD^{+}05]. variable [BBS02a, CDM03]. variables [BH04, CKB04]. VBR [MNH01]. vector [HS01]. vector-based [HS01]. Velocity [HFMN09]. verification [BSV05, CCG^{+}04, CJMR07, DS00, FRC^{+}02, HP00, HS02, HLT^{+}04, Pik06, Rus99, SBMJ01, TA04, Xu03, XN06]. verified [LRSF04]. Verifying [PIM09, EW04]. Verisim [BGK^{+}02]. version [ABGM02, WMC01, ZZWD05, ABGM02]. versus [AS04, KMT07, MOJ05]. VERTAF [HLT^{+}04]. via [BSC03, CLZ02, NPD01]. viewpoint [AAG^{+}04]. viewpoint-oriented [AAG^{+}04]. Views [RAW08, LD03]. Vigilance [LBJ06]. Virtual [CHMB08, JWB02]. visibility [VMA^{+}04]. visible [AP03]. Visual [Moo09, BCRW00, CDFG06, LD03, ZZC01]. Visualization [GHM08, BH04]. Visualizing [DLL09, DYZ07, EGK^{+}02]. visually [DL05]. Vulnerabilities [DVJP08]. Vulnerability [CCR07, TW07].

warehousing [LG03]. Warp [FRC^{+}02]. Warp-based [FRC^{+}02]. WASP [HOM08]. Watermarking [CT02]. way [SM03a]. weakest [CN00]. Web [BDFZ08, DVJP08, ERKF05, EA08, FBS05, HOM08, HVC09, KT01, MM08b, SSG^{+}07, TRL04, ZBN^{+}04, ZLW07]. weights [ATG^{+}06]. Well [TFD03]. Well-defined [TFD03]. Weyuker [ZX02, GR01, SJJ06]. WFMS [CDF01]. where [EH01]. white [BHS03]. Whole [BHSLB03, THY03]. Whole-Part [BHSLB03, THY03]. wireless [BCMS03, CCW03, MO03, NKL05, OBMP03, WCL04]. wise [BSR04]. Within [KMT07]. Within-Company [KMT07]. wizards [BCRW00]. Work [CMRH09]. workbenches [CDFG06]. workflow [HA00]. workflows [CCF^{+}00]. Working [vDS05]. Workload [KRM06, TRL04]. Workshop [CCW00b, CCW00a]. World [WW00]. Wp [ENDK02]. Wp-method [ENDK02]. wrappers [ESWH04].

X [BBG06, KKJ08]. X-FEDERATE [BBG06]. XML [KMS04, WCL04]. XP [HFMN09]. Xstream [WCL04].

Year [Ano06-30, Ano06-29, Ano06-28].

Z [MD00]. ZPL [CCL^{+}00].

References

Andrade:2004:MFV

[AAG^{+}04] J. Andrade, J. Ares, R. Garcia, J. Pazos, S. Rodriguez, and A. Silva. A methodological framework for viewpointoriented conceptual modeling. IEEE Transactions on Software Engineering, 30(5):282– 294, May 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1291832.

Adve:2000:PEE

[ABB^{+}00] Vikram S. Adve, Rajive L.

Bagrodia, James C. Browne, Ewa Deelman, Aditya Dube, Elias N. Houstis, John R. Rice, Rizos Sakellariou, David J. Sundaram-Stukel, Patricia J. Teller, and Mary K. Vernon. POEMS: End-to-end performance design of large parallel adaptive computational systems. IEEE Transactions on Software Engineering, 26(11):1027–1048, November 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=881716.

Arisholm:2004:DCM

[ABF04] E. Arisholm, L. C. Briand, and A. Foyen. Dynamic coupling measurement for objectoriented software. IEEE Transactions on Software Engineering, 30(8):491–506, August 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1316867.

Atkins:2002:UVC

[ABGM02] D. L. Atkins, T. Ball, T. L. Graves, and A. Mockus. Using version control data to evaluate the impact of software tools: a case study of the Version Editor. IEEE Transactions on Software Engineering, 28(7):625–637, July 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1019478.

Arisholm:2006:IUD

[ABHL06] E. Arisholm, L. C. Briand, S. E. Hove, and Y. Labiche. The impact of UML documentation on software maintenance: an experimental evaluation. IEEE Transactions on Software Engineering, 32 (6):365–381, June 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1650213.

Andrews:2006:UMA

[ABLN06] J. H. Andrews, L. C. Briand, Y. Labiche, and A. S. Namin. Using mutation analysis for assessing and comparing testing coverage criteria. IEEE Transactions on Software Engineering, 32(8):608–624, August 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1703390.

Agrawal:2007:SEQ

[AC07] M. Agrawal and K. Chari. Software effort, quality, and cycle time: A study of CMM level 5 projects. IEEE Transactions on Software Engineering, 33(3): 145–156, March 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4084133.

Antoniol:2002:RTL

[ACC^{+}02] G. Antoniol, G. Canfora, G. Casazza, A. De Lucia, and E. Merlo. Recovering traceability links between code and documentation. IEEE Transactions on Software Engineering, 28(10): 970–983, October 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1041053.

Antoniol:2004:ASN

[ACDD04] G. Antoniol, A. Cimitile, G. A.

Di Lucca, and M. Di Penta. Assessing staffing needs for a software maintenance project through queuing simulation. IEEE Transactions on Software Engineering, 30(1):43– 58, January 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265735.

Alencar:2002:LTI

[ACL02] P. S. C. Alencar, D. D. Cowan, and C. J. P. Lucena. A logical theory of interfaces and objects. IEEE Transactions on Software Engineering, 28 (6):548–575, June 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1010059.

Apvrille:2004:TRT

[ACLdSS04] L. Apvrille, J. P. Courtiat, C. Lohr, and P. de Saqui-Sannes. TURTLE: a realtime UML profile supported by a formal validation toolkit. IEEE Transactions on Software Engineering, 30(7):473– 487, July 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1318608.

Antkiewicz:2009:EFS

[ACS09] M. Antkiewicz, K. Czarnecki, and M. Stephan. Engineering of framework-specific modeling languages. IEEE Transactions on Software Engineering, 35(6):795–824, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4907004.

Abraham:2009:MOS

[AE09] R. Abraham and M. Erwig. Mutation operators for spreadsheets. IEEE Transactions on Software Engineering, 35(1):94–108, January/ February 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4609389.

Alur:2003:IMS

[AEY03] R. Alur, K. Etessami, and M. Yannakakis. Inference of message sequence charts. IEEE Transactions on Software Engineering, 29(7):623– 633, July 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1214326.

Antoniol:2006:FIE

[AG06] G. Antoniol and Y. G.

Guéhéneuc. Feature identification: An epidemiological metaphor. IEEE Transactions on Software Engineering, 32(9): 627–641, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1707664.

Arisholm:2007:EPP

[AGDS07] E. Arisholm, H. Gallis, T. Dyba, and Dag I. K. Sjøberg. Evaluating pair programming with respect to system complexity and programmer expertise. IEEE Transactions on Software Engineering, 33(2): 65–86, February 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4052584.

Atkinson:2009:FIM

[AGK09] C. Atkinson, M. Gutheil, and B. Kennel. A flexible infrastructure for multilevel language engineering. IEEE Transactions on Software Engineering, 35(6):742–755, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4907005.

Antoy:2000:ACI

[AH00] S. Antoy and D. Hamlet. Automatically checking an implementation against its formal specification. IEEE Transactions on Software Engineering, 26(1):55–69, January 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=825766.

Alshayeb:2003:EVO

[AL03] M. Alshayeb and Wei Li. An empirical validation of objectoriented metrics in two different iterative software processes.

IEEE Transactions on Software Engineering, 29(11):1043– 1049, November 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1245305.

Apel:2008:AFM

[ALS08] S. Apel, T. Leich, and G. Saake.

Aspectual feature modules. IEEE Transactions on Software Engineering, 34(2):162– 180, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4407729.

Andrews:2001:E

[And01] A. A. Andrews. Editorial.

IEEE Transactions on Software Engineering, 27(5):385– 386, May 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=922712.

Andrews:2002:ENE

[And02] A. A. Andrews. Editorial: New EIC introduction. IEEE Transactions on Software Engineering, 28(1):1, January 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=979984.

Anonymous:2000:RL

[Ano00] Anonymous. 1999 reviewers list. IEEE Transactions on Software Engineering, 26(1): 94–96, January 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=825768.

Anonymous:2001:RL

[Ano01a] Anonymous. 2000 reviewers list. IEEE Transactions on Software Engineering, 27(1): 94–96, January 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=895991.

Anonymous:2001:AI

[Ano01b] Anonymous. Author index. IEEE Transactions on Software Engineering, 27(12):1145– 1147, December 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=988712.

Anonymous:2001:SI

[Ano01c] Anonymous. Subject index. IEEE Transactions on Software Engineering, 27(12):1147– 1151, December 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=988713.

Anonymous:2002:AIa

[Ano02a] Anonymous. Author index.

IEEE Transactions on Software Engineering, 28(11):1194– 1196, November 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1049408.

Anonymous:2002:AIb

[Ano02b] Anonymous. Author index.

IEEE Transactions on Software Engineering, 28(12):1194– 1196, December 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1158291.

Anonymous:2002:SIa

[Ano02c] Anonymous. Subject index.

IEEE Transactions on Software Engineering, 28(11):1196– 1200, November 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1049409.

Anonymous:2002:SIb

[Ano02d] Anonymous. Subject index.

IEEE Transactions on Software Engineering, 28(12):1196– 1200, December 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1158292.

Anonymous:2003:AI

[Ano03a] Anonymous. Author index. IEEE Transactions on Software Engineering, 29(12):1135– 1137, December 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1265527.

Anonymous:2003:RL

[Ano03b] Anonymous. Reviewers list. IEEE Transactions on Software Engineering, 29(1):95– 96, January 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1166592.

Anonymous:2003:SSI

[Ano03c]  Anonymous.  Special sec-

tion — on the international conference on the foundations of software engineering. IEEE Transactions on Software Engineering, 29(10):0:1– 0:2, October 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1237167.

Anonymous:2003:SI

[Ano03d] Anonymous. Subject index. IEEE Transactions on Software Engineering, 29(12):1137– 1142, December 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265528.

Anonymous:2004:RL

[Ano04a] Anonymous. 2003 reviewers list. IEEE Transactions on Software Engineering, 30(1): 78–80, January 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265737.

Anonymous:2004:Ae

[Ano04b] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 30(10):709, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339282.

Anonymous:2004:Aj

[Ano04c] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 30(10):710, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339283.

Anonymous:2004:Al

[Ano04d] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 30(10):711, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1339284.

Anonymous:2004:Ad

[Ano04e] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 30(11):833, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1359774.

Anonymous:2004:Aa

[Ano04f] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 30(11):834, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1359775.

Anonymous:2004:Af

[Ano04g] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 30(11):835, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1359776.

Anonymous:2004:Ab

[Ano04h] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 30(11):836, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1359777.

Anonymous:2004:Ag

[Ano04i] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 30(11):837, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1359778.

Anonymous:2004:Ac

[Ano04j] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 30(11):838, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1359779.

Anonymous:2004:Ah

[Ano04k] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 30(11):839, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1359780.

Anonymous:2004:Ak

[Ano04l] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 30(11):840, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1359781.

Anonymous:2004:AI

[Ano04m] Anonymous. Annual index. IEEE Transactions on Software Engineering, 30(12):1084– 1096, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1377199.

Anonymous:2004:BCb

[Ano04n]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 30(5):c4, May 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1291839.

Anonymous:2004:BCh

[Ano04o]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 30(6):c4, June 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1321068.

Anonymous:2004:BCa

[Ano04p]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 30(7):c4, July 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1318611.

Anonymous:2004:BCc

[Ano04q]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 30(8):c4, August 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1316873.

Anonymous:2004:BCe

[Ano04r]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 30(9):c4, September 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1324650.

Anonymous:2004:BCg

[Ano04s]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 30(10):c4, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339287.

Anonymous:2004:BCd

[Ano04t]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 30(11):c4, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1359783.

Anonymous:2004:BCf

[Ano04u]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 30(12):c4, December 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1377201.

Anonymous:2004:ITSe

[Ano04v] Anonymous. IEEE Transactions on Software Engineering. IEEE Transactions on Software Engineering, 30(2): C2, February 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1265812.

Anonymous:2004:ITSb

[Ano04w] Anonymous. IEEE Transactions on Software Engineering. IEEE Transactions on Software Engineering, 30(3): 0:1, March 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1271167.

Anonymous:2004:ITSa

[Ano04x] Anonymous. IEEE Transactions on Software Engineering. IEEE Transactions on Software Engineering, 30(4): 0:1, April 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1274039.

Anonymous:2004:ITSc

[Ano04y] Anonymous. IEEE Transactions on Software Engineering — staff list. IEEE Transactions on Software Engineering, 30(1):0:2, January 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265730.

Anonymous:2004:ITSf

[Ano04z]  Anonymous.  IEEE Trans-

actions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, 30 (1):0:1, January 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265724.

Anonymous:2004:ITSd

[Ano04-27] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, 30(2):C1, February 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265811.

Anonymous:2004:CP

[Ano04-28] Anonymous. Call for papers. IEEE Transactions on Software Engineering, 30(7): 488, July 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1318609.

Anonymous:2004:CPI

[Ano04-29] Anonymous. Call for papers for the IEEE Transactions on Software Engineering special issue on mining software repositories. IEEE Transactions on Software Engineering, 30 (10):712, October 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1339285.

Anonymous:2004:ENA

[Ano04-30] Anonymous. Editorial: New AE introduction. IEEE Transactions on Software Engineering, 30(11):713–714, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1359766.

Anonymous:2004:FCd

[Ano04-31] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 30(5):c1, May 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1291829.

Anonymous:2004:FCb

[Ano04-32] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 30(6):c1, June 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1321056.

Anonymous:2004:FCg

[Ano04-33] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 30(7):c1, July 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1318602.

Anonymous:2004:FCh

[Ano04-34] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 30(8):c1, August 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1316864.

Anonymous:2004:FCa

[Ano04-35] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 30(9):c1, September 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1324642.

Anonymous:2004:FCe

[Ano04-36] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 30(10):c1, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339273.

Anonymous:2004:FCc

[Ano04-37] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 30(11):c1, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1359764.

Anonymous:2004:FCf

[Ano04-38] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 30(12):c1, December 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377182.

Anonymous:2004:ICSa

[Ano04-39] Anonymous. IEEE Computer Society — information. IEEE Transactions on Software Engineering, 30(4): 280, April 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1274048.

Anonymous:2004:ICSe

[Ano04-40] Anonymous. IEEE Computer Society information. IEEE Transactions on Software Engineering, 30(1):82, January 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265739.

Anonymous:2004:ICSb

[Ano04-41] Anonymous. IEEE Computer Society information. IEEE Transactions on Software Engineering, 30(3):208, March 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1271176.

Anonymous:2004:ICSd

[Ano04-42] Anonymous. IEEE Computer Society information to authors. IEEE Transactions on Software Engineering, 30 (2):142, February 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265820.

Anonymous:2004:ICSf

[Ano04-43] Anonymous. IEEE Computer Society staff list. IEEE Transactions on Software Engineering, 30(3):0:2, March 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1271168.

Anonymous:2004:ICSc

[Ano04-44] Anonymous. IEEE Computer Society’s — staff list. IEEE Transactions on Software Engineering, 30(4):0:2, April 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1274040.

Anonymous:2004:IAa

[Ano04-45] Anonymous. Information for authors. IEEE Transactions on Software Engineering, 30 (2):141, February 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1265819.

Anonymous:2004:IAb

[Ano04-46] Anonymous. Information for authors. IEEE Transactions on Software Engineering, 30 (3):207, March 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1271175.

Anonymous:2004:IFCc

[Ano04-47] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 30 (5):c2, May 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1291830.

Anonymous:2004:IFCa

[Ano04-48] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 30 (6):c2, June 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1321057.

Anonymous:2004:IFCh

[Ano04-49] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 30 (7):c2, July 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1318603.

Anonymous:2004:IFCe

[Ano04-50] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 30 (8):c2, August 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1316865.

Anonymous:2004:IFCf

[Ano04-51] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 30 (9):c2, September 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1324643.

Anonymous:2004:IFCd

[Ano04-52] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 30 (10):c2, October 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339274.

Anonymous:2004:IFCb

[Ano04-53] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 30 (11):c2, November 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1359765.

Anonymous:2004:IFCg

[Ano04-54] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 30 (12):c2, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377183.

Anonymous:2004:TIAb

[Ano04-55] Anonymous. TSE: Information for authors. IEEE Transactions on Software Engineering, 30(1):81, January 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265738.

Anonymous:2004:TIAa

[Ano04-56] Anonymous. TSE: Information for authors. IEEE Transactions on Software Engineering, 30(4):279, April 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1274047.

Anonymous:2004:TIAd

[Ano04-57] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 30(5):c3, May 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1291838.

Anonymous:2004:TIAj

[Ano04-58] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 30(6):c3, June 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1321067.

Anonymous:2004:TIAc

[Ano04-59] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 30(7):c3, July 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1318610.

Anonymous:2004:TIAe

[Ano04-60] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 30(8):c3, August 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1316872.

Anonymous:2004:TIAg

[Ano04-61] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 30(9):c3, September 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1324649.

Anonymous:2004:TIAi

[Ano04-62] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 30(10):c3, October 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1339286.

Anonymous:2004:TIAf

[Ano04-63] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 30(11):c3, November 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1359782.

Anonymous:2004:TIAh

[Ano04-64] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 30(12):c3, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377200.

Anonymous:2005:RL

[Ano05a] Anonymous. 2004 reviewers list. IEEE Transactions on Software Engineering, 31(1): 91–93, January 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1392722.

Anonymous:2005:Ag

[Ano05b] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 31(1):95, January 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1392724.

Anonymous:2005:Al

[Ano05c] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 31(1):96, January 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1392725.

Anonymous:2005:Am

[Ano05d] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 31(2):182, February 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1401932.

Anonymous:2005:Af

[Ano05e] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 31(2):183, February 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1401933.

Anonymous:2005:Ak

[Ano05f] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 31(2):184, February 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1401934.

Anonymous:2005:Ab

[Ano05g] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 31(4):356– 358, April 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1435355.

Anonymous:2005:Ah

[Ano05h] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 31(4):359, April 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1435356.

Anonymous:2005:Ad

[Ano05i] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 31(4):360, April 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1435357.

Anonymous:2005:Aj

[Ano05j] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 31(6):527, June 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1463234.

Anonymous:2005:Ae

[Ano05k] Anonymous. [advertisement].

IEEE Transactions on Software Engineering, 31(6):528, June 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1463235.

Anonymous:2005:Aa

[Ano05l] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 31(8):712, August 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1498774.

Anonymous:2005:Ac

[Ano05m] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 31(10):912, October 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1542072.

Anonymous:2005:Ai

[Ano05n] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 31(11):996, November 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1556557.

Anonymous:2005:BCh

[Ano05o]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(1):c4, January 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1392727.

Anonymous:2005:BCg

[Ano05p]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(2):c4, February 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1401936.

Anonymous:2005:BCb

[Ano05q]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(3):c4, March 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1423998.

Anonymous:2005:BCe

[Ano05r]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(4):c4, April 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1435359.

Anonymous:2005:BCi

[Ano05s]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(5):c4, May 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1438378.

Anonymous:2005:BCk

[Ano05t]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(6):c4, June 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1463237.

Anonymous:2005:BCd

[Ano05u]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(7):c4, July 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1492377.

Anonymous:2005:BCa

[Ano05v]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(8):c4, August 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1498776.

Anonymous:2005:BCf

[Ano05w]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(9):c4, September 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1514448.

Anonymous:2005:BCj

[Ano05x]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(10):c4, October 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1542074.

Anonymous:2005:BCl

[Ano05y]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(11):c4, November 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1556559.

Anonymous:2005:BCc

[Ano05z]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 31(12):c4, December 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1566610.

Anonymous:2005:CP

[Ano05-27] Anonymous. Call for papers. IEEE Transactions on Software Engineering, 31(1): 94, January 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1392723.

Anonymous:2005:EAE

[Ano05-28] Anonymous. Editorial: Associate Editor introduction and farewell. IEEE Transactions on Software Engineering, 31 (1):2, January 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1392716.

Anonymous:2005:ESJ

[Ano05-29] Anonymous. Editorial: State of the journal address. IEEE Transactions on Software Engineering, 31(1):1, January 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1392715.

Anonymous:2005:FCd

[Ano05-30] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 31(1):c1, January 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1392713.

Anonymous:2005:FCg

[Ano05-31] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 31(2):c1, February 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1401924.

Anonymous:2005:FCh

[Ano05-32] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 31(3):c1, March 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1423988.

Anonymous:2005:FCc

[Ano05-33] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 31(4):c1, April 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1435347.

Anonymous:2005:FCi

[Ano05-34] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 31(5):c1, May 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1438371.

Anonymous:2005:FCe

[Ano05-35] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 31(6):c1, June 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1463224.

Anonymous:2005:FCa

[Ano05-36] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 31(7):c1, July 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1492367.

Anonymous:2005:FCf

[Ano05-37] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 31(8):c1, August 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1498767.

Anonymous:2005:FCl

[Ano05-38] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 31(9):c1, September 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1514440.

Anonymous:2005:FCk

[Ano05-39] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 31(10):c1, October 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1542061.

Anonymous:2005:FCb

[Ano05-40] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 31(11):c1, November 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1556549.

Anonymous:2005:FCj

[Ano05-41] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 31(12):c1, December 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1566600.

Anonymous:2005:IFCh

[Ano05-42] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (1):c2, January 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1392714.

Anonymous:2005:IFCd

[Ano05-43] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (2):c2, February 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1401925.

Anonymous:2005:IFCe

[Ano05-44] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (3):c2, March 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1423989.

Anonymous:2005:IFCb

[Ano05-45] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (4):c2, April 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1435348.

Anonymous:2005:IFCj

[Ano05-46] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (5):c2, May 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1438372.

Anonymous:2005:IFCc

[Ano05-47] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (6):c2, June 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1463225.

Anonymous:2005:IFCg

[Ano05-48] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (7):c2, July 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1492368.

Anonymous:2005:IFCi

[Ano05-49] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (8):c2, August 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1498768.

Anonymous:2005:IFCl

[Ano05-50] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (9):c2, September 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1514441.

Anonymous:2005:IFCf

[Ano05-51] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (10):c2, October 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1542062.

Anonymous:2005:IFCa

[Ano05-52] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (11):c2, November 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1556550.

Anonymous:2005:IFCk

[Ano05-53] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 31 (12):c2, December 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1566601.

Anonymous:2005:TIAf

[Ano05-54] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(1):c3, January 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1392726.

Anonymous:2005:TIAe

[Ano05-55] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(2):c3, February 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1401935.

Anonymous:2005:TIAc

[Ano05-56] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(3):c3, March 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1423997.

Anonymous:2005:TIAj

[Ano05-57] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(4):c3, April 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1435358.

Anonymous:2005:TIAh

[Ano05-58] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(5):c3, May 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1438377.

Anonymous:2005:TIAk

[Ano05-59] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(6):c3, June 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1463236.

Anonymous:2005:TIAg

[Ano05-60] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(7):c3, July 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1492376.

Anonymous:2005:TIAa

[Ano05-61] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(8):c3, August 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1498775.

Anonymous:2005:TIAb

[Ano05-62] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(9):c3, September 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1514447.

Anonymous:2005:TIAi

[Ano05-63] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(10):c3, October 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1542073.

Anonymous:2005:TIAl

[Ano05-64] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(11):c3, November 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1556558.

Anonymous:2005:TIAd

[Ano05-65] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 31(12):c3, December 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1566609.

Anonymous:2006:AI

[Ano06a] Anonymous. 2005 annual index. IEEE Transactions on Software Engineering, 32(1): 54–64, January 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1583603.

Anonymous:2006:RL

[Ano06b] Anonymous. 2005 reviewers list. IEEE Transactions on Software Engineering, 32(1): 51–53, January 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1583602.

Anonymous:2006:A

[Ano06c] Anonymous. [advertisement]. IEEE Transactions on Software Engineering, 32(2):136, February 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1599421.

Anonymous:2006:BCb

[Ano06d]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(1):c4, January 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1583605.

Anonymous:2006:BCk

[Ano06e]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(2):c4, February 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1599423.

Anonymous:2006:BCj

[Ano06f]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(3):c4, March 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1610614.

Anonymous:2006:BCh

[Ano06g]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(4):c4, April 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1628975.

Anonymous:2006:BCa

[Ano06h]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(5):c4, May 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1642683.

Anonymous:2006:BCl

[Ano06i]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(6):c4, June 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1650220.

Anonymous:2006:BCe

[Ano06j]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(7):c4, July 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1677540.

Anonymous:2006:BCd

[Ano06k]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(8):c4, August 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1703392.

Anonymous:2006:BCi

[Ano06l]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(9):c4, September 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1707673.

Anonymous:2006:BCf

[Ano06m]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(10):c4, October 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1717476.

Anonymous:2006:BCc

[Ano06n]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(11):c4, November 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4015516.

Anonymous:2006:BCg

[Ano06o]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 32(12):c4, December 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4016577.

Anonymous:2006:FCb

[Ano06p]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(1):c1, January 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1583595.

Anonymous:2006:FCl

[Ano06q]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(2):c1, February 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1599415.

Anonymous:2006:FCi

[Ano06r]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(3):c1, March 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1610604.

Anonymous:2006:FCe

[Ano06s]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(4):c1, April 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1628967.

Anonymous:2006:FCh

[Ano06t]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(5):c1, May 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1642675.

Anonymous:2006:FCj

[Ano06u]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(6):c1, June 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1650210.

Anonymous:2006:FCf

[Ano06v]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(7):c1, July 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1677529.

Anonymous:2006:FCg

[Ano06w]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(8):c1, August 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1703382.

Anonymous:2006:FCd

[Ano06x]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(9):c1, September 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1707661.

Anonymous:2006:FCc

[Ano06y]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(10):c1, October 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1717468.

Anonymous:2006:FCa

[Ano06z]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 32(11):c1, November 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4015507.

Anonymous:2006:FCk

[Ano06-27] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 32(12):c1, December 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4016568.

Anonymous:2006:ICSa

[Ano06-28] Anonymous. IEEE Computer Society celebrates two 60-year anniversaries. IEEE Transactions on Software Engineering, 32(3):212, March 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1610612.

Anonymous:2006:ICSc

[Ano06-29] Anonymous. The IEEE Computer Society celebrates two 60-year anniversaries. IEEE Transactions on Software Engineering, 32(6):432, June 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1650218.

Anonymous:2006:ICSb

[Ano06-30] Anonymous. The IEEE Computer Society celebrates two 60-year anniversaries. IEEE Transactions on Software Engineering, 32(7):527, July 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1677537.

Anonymous:2006:I

[Ano06-31] Anonymous. In this issue. IEEE Transactions on Software Engineering, 32(11):928, November 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4015514.

Anonymous:2006:IFCb

[Ano06-32] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (1):c2, January 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1583596.

Anonymous:2006:IFCl

[Ano06-33] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (2):c2, February 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1599416.

Anonymous:2006:IFCf

[Ano06-34] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (3):c2, March 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1610605.

Anonymous:2006:IFCd

[Ano06-35] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (4):c2, April 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1628968.

Anonymous:2006:IFCk

[Ano06-36] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (5):c2, May 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1642676.

Anonymous:2006:IFCg

[Ano06-37] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (6):c2, June 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1650211.

Anonymous:2006:IFCj

[Ano06-38] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (7):c2, July 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1677530.

Anonymous:2006:IFCe

[Ano06-39] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (8):c2, August 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1703383.

Anonymous:2006:IFCi

[Ano06-40] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (9):c2, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1707662.

Anonymous:2006:IFCc

[Ano06-41] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (10):c2, October 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1717469.

Anonymous:2006:IFCa

[Ano06-42] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (11):c2, November 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4015508.

Anonymous:2006:IFCh

[Ano06-43] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 32 (12):c2, December 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4016569.

Anonymous:2006:JICb

[Ano06-44] Anonymous. Join the IEEE Computer Society! IEEE Transactions on Software Engineering, 32(7):528, July 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1677538.

Anonymous:2006:JICa

[Ano06-45] Anonymous. Join the IEEE Computer Society. IEEE Transactions on Software Engineering, 32(12):1006–1008, December 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4016575.

Anonymous:2006:TIAj

[Ano06-46] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(1):c3, January 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1583604.

Anonymous:2006:TIAk

[Ano06-47] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(2):c3, February 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1599422.

Anonymous:2006:TIAi

[Ano06-48] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(3):c3, March 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1610613.

Anonymous:2006:TIAg

[Ano06-49] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(4):c3, April 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1628974.

Anonymous:2006:TIAd

[Ano06-50] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(5):c3, May 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1642682.

Anonymous:2006:TIAl

[Ano06-51] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(6):c3, June 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1650219.

Anonymous:2006:TIAc

[Ano06-52] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(7):c3, July 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1677539.

Anonymous:2006:TIAb

[Ano06-53] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(8):c3, August 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1703391.

Anonymous:2006:TIAh

[Ano06-54] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(9):c3, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1707672.

Anonymous:2006:TIAe

[Ano06-55] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(10):c3, October 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1717475.

Anonymous:2006:TIAa

[Ano06-56] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(11):c3, November 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4015515.

Anonymous:2006:TIAf

[Ano06-57] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 32(12):c3, December 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4016576.

Anonymous:2007:AICa

[Ano07a] Anonymous. 180,000 aritlces in the IEEE Computer Society digital library [advertisement]. IEEE Transactions on Software Engineering, 33 (8):575, August 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4267027.

Anonymous:2007:AICb

[Ano07b] Anonymous. 180,000 aritlces in the IEEE Computer Society digital library [advertisement]. IEEE Transactions on Software Engineering, 33(11): 796, November 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4339234.

Anonymous:2007:AI

[Ano07c] Anonymous. 2006 annual index. IEEE Transactions on Software Engineering, 33(1): 57–63, January 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4027149.

Anonymous:2007:RL

[Ano07d] Anonymous. 2006 reviewers list. IEEE Transactions on Software Engineering, 33(1): 54–56, January 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4027148.

Anonymous:2007:BCj

[Ano07e]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(1):c4, January 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=4027152.

Anonymous:2007:BCg

[Ano07f]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(2):c4, February 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4052590.

Anonymous:2007:BCb

[Ano07g]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(3):c4, March 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4084139.

Anonymous:2007:BCa

[Ano07h]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(4):c4, April 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4123331.

Anonymous:2007:BCc

[Ano07i]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(5):c4, May 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4160975.

Anonymous:2007:BCk

[Ano07j]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(6):c4, June 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4181712.

Anonymous:2007:BCl

[Ano07k]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(7):c4, July 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4227831.

Anonymous:2007:BCf

[Ano07l]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(8):c4, August 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4267030.

Anonymous:2007:BCd

[Ano07m]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(9):c4, September 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4288199.

Anonymous:2007:BCi

[Ano07n]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(10):c4, October 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4302783.

Anonymous:2007:BCe

[Ano07o]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(11):–??, November 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4339236.

Anonymous:2007:BCh

[Ano07p]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 33(12):c4, December 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4375383.

Anonymous:2007:CPS

[Ano07q] Anonymous. Call-for-papers:

Special issue on software language engineering. IEEE Transactions on Software Engineering, 33(12):891, December 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4375380.

Anonymous:2007:ENA

[Ano07r] Anonymous. Editorial: New associate editors introduction. IEEE Transactions on Software Engineering, 33(10):641– 642, October 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4302776.

Anonymous:2007:FCk

[Ano07s]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 33(1):c1, January 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4027142.

Anonymous:2007:FCg

[Ano07t]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 33(2):c1, February 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4052582.

Anonymous:2007:FCi

[Ano07u]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 33(3):c1, March 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4084131.

Anonymous:2007:FCc

[Ano07v]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 33(4):c1, April 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4123322.

Anonymous:2007:FCl

[Ano07w]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 33(5):c1, May 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4160965.

Anonymous:2007:FCa

[Ano07x]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 33(6):c1, June 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4181705.

Anonymous:2007:FCd

[Ano07y]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 33(7):c1, July 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4227825.

Anonymous:2007:FCj

[Ano07z]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 33(8):c1, August 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4267020.

Anonymous:2007:FCe

[Ano07-27] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 33(9):c1, September 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4288190.

Anonymous:2007:FCf

[Ano07-28] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 33(10):c1, October 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4302774.

Anonymous:2007:FCb

[Ano07-29] Anonymous. [front cover]. IEEE Transactions on Software Engineering, 33(11):c1, November 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4339228.

Anonymous:2007:FCh

[Ano07-30] Anonymous. [front cover].

IEEE Transactions on Software Engineering, 33(12):c1, December 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4375376.

Anonymous:2007:I

[Ano07-31] Anonymous. In this issue. IEEE Transactions on Software Engineering, 33(1): 64, January 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4027150.

Anonymous:2007:IT

[Ano07-32] Anonymous. In this issue — technically. IEEE Transactions on Software Engineering, 33(4):269, April 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4123328.

Anonymous:2007:IFCk

[Ano07-33] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (1):c2, January 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4027143.

Anonymous:2007:IFCe

[Ano07-34] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (2):c2, February 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4052583.

Anonymous:2007:IFCj

[Ano07-35] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (3):c2, March 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4084132.

Anonymous:2007:IFCc

[Ano07-36] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (4):c2, April 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4123323.

Anonymous:2007:IFCl

[Ano07-37] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (5):c2, May 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4160966.

Anonymous:2007:IFCa

[Ano07-38] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (6):c2, June 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4181706.

Anonymous:2007:IFCh

[Ano07-39] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (7):c2, July 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4227826.

Anonymous:2007:IFCg

[Ano07-40] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (8):c2, August 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4267021.

Anonymous:2007:IFCi

[Ano07-41] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (9):c2, September 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4288191.

Anonymous:2007:IFCd

[Ano07-42] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (10):c2, October 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4302775.

Anonymous:2007:IFCb

[Ano07-43] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (11):c2, November 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4339229.

Anonymous:2007:IFCf

[Ano07-44] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 33 (12):c2, December 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4375377.

Anonymous:2007:JICd

[Ano07-45] Anonymous. Join the IEEE Computer Society! IEEE Transactions on Software Engineering, 33(2):142–144, February 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4052588.

Anonymous:2007:JICa

[Ano07-46] Anonymous. Join the IEEE Computer Society. IEEE Transactions on Software Engineering, 33(4):270–272, April 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4123329.

Anonymous:2007:JICb

[Ano07-47] Anonymous. Join the IEEE Computer Society! IEEE Transactions on Software Engineering, 33(5):366–368, May 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4160973.

Anonymous:2007:JICc

[Ano07-48] Anonymous. Join the IEEE Computer Society. IEEE Transactions on Software Engineering, 33(8):576, August 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4267028.

Anonymous:2007:JICe

[Ano07-49] Anonymous. Join the IEEE Computer Society. IEEE Transactions on Software Engineering, 33(12):892, December 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4375381.

Anonymous:2007:TIAk

[Ano07-50] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(1):c3, January 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4027151.

Anonymous:2007:TIAg

[Ano07-51] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(2):c3, February 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4052589.

Anonymous:2007:TIAb

[Ano07-52] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(3):c3, March 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4084138.

Anonymous:2007:TIAa

[Ano07-53] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(4):c3, April 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4123330.

Anonymous:2007:TIAc

[Ano07-54] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(5):c3, May 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4160974.

Anonymous:2007:TIAl

[Ano07-55] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(6):c3, June 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4181711.

Anonymous:2007:TIAj

[Ano07-56] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(7):c3, July 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4227830.

Anonymous:2007:TIAf

[Ano07-57] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(8):c3, August 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4267029.

Anonymous:2007:TIAd

[Ano07-58] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(9):c3, September 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4288198.

Anonymous:2007:TIAi

[Ano07-59] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(10):c3, October 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4302782.

Anonymous:2007:TIAe

[Ano07-60] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(11):c3, November 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4339235.

Anonymous:2007:TIAh

[Ano07-61] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 33(12):c3, December 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4375382.

Anonymous:2008:AI

[Ano08a]  Anonymous.  2007 annual

index. IEEE Transactions on Software Engineering, 34 (1):not in print, January/ February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4444349.

Anonymous:2008:RL

[Ano08b] Anonymous. 2007 reviewer’s list. IEEE Transactions on Software Engineering, 34(1): 154–156, January/February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4444344.

Anonymous:2008:BCc

[Ano08c]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 34(1):c4, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4444351.

Anonymous:2008:BCb

[Ano08d]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 34(2):c4, March/April 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4476758.

Anonymous:2008:BCd

[Ano08e]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 34(3):c4, May/June 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4536055.

Anonymous:2008:BCe

[Ano08f]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 34(4):c4, July/August 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4585927.

Anonymous:2008:BCa

[Ano08g]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 34(5):c4, September/October 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4636878.

Anonymous:2008:BCf

[Ano08h]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 34(6):c4, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4709495.

Anonymous:2008:CPS

[Ano08i]  Anonymous.  Call for pa-

pers: Search-based optimization for software engineering. IEEE Transactions on Software Engineering, 34(2):304, March/April 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4476756.

Anonymous:2008:FCe

[Ano08j]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 34(1):c1, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4444340.

Anonymous:2008:FCd

[Ano08k]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 34(2):c1, March/April 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4476752.

Anonymous:2008:FCc

[Ano08l]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 34(3):c1, May/June 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4536052.

Anonymous:2008:FCa

[Ano08m]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 34(4):c1, July/August 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4585923.

Anonymous:2008:FCb

[Ano08n]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 34(5):c1, September/October 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4636874.

Anonymous:2008:FCf

[Ano08o]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 34(6):c1, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4709488.

Anonymous:2008:ICSa

[Ano08p] Anonymous. IEEE Computer Society 2009 membership application. IEEE Transactions on Software Engineering, 34(6):860–862, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4709491.

Anonymous:2008:ICSb

[Ano08q] Anonymous. IEEE Computer Society career center. IEEE Transactions on Software Engineering, 34(6):863, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4709492.

Anonymous:2008:ICSc

[Ano08r]  Anonymous.  IEEE Com-

puter Society CSDP [advertisement]. IEEE Transactions on Software Engineering, 34(6):864, November/ December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4709493.

Anonymous:2008:ITb

[Ano08s] Anonymous. In this issue — technically. IEEE Transactions on Software Engineering, 34(1):157, January/ February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4444345.

Anonymous:2008:ITc

[Ano08t] Anonymous. In this issue — technically. IEEE Transactions on Software Engineering, 34(1):158, January/ February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4444346.

Anonymous:2008:ITa

[Ano08u] Anonymous. In this issue — technically. IEEE Transactions on Software Engineering, 34(1):159, January/ February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4444347.

Anonymous:2008:IFCc

[Ano08v]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 34(1):c2, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4444341.

Anonymous:2008:IFCb

[Ano08w]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 34(2): c2, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4476753.

Anonymous:2008:IFCe

[Ano08x]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 34 (3):c2, May/June 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4536053.

Anonymous:2008:IFCd

[Ano08y]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 34(4): c2, July/August 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4585924.

Anonymous:2008:IFCa

[Ano08z]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 34(5):c2, September/October 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4636875.

Anonymous:2008:IFCf

[Ano08-27] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, 34 (6):c2, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4709489.

Anonymous:2008:SJS

[Ano08-28] Anonymous. Society journals seek editors in chief for 2009-2010 terms. IEEE Transactions on Software Engineering, 34(1):160, January/ February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4444348.

Anonymous:2008:TIAd

[Ano08-29] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 34(1):c3, January/ February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4444350.

Anonymous:2008:TIAc

[Ano08-30] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 34(2): c3, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4476757.

Anonymous:2008:TIAe

[Ano08-31] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 34(3):c3, May/June 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4536054.

Anonymous:2008:TIAb

[Ano08-32] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 34(4): c3, July/August 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4585926.

Anonymous:2008:TIAa

[Ano08-33] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 34(5):c3, September/October 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4636877.

Anonymous:2008:TIAf

[Ano08-34] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 34(6):c3, November/ December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4709494.

Anonymous:2009:RL

[Ano09a]  Anonymous.  2008 review-

ers list. IEEE Transactions on Software Engineering, 35(1):138–141, January/ February 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4771849.

Anonymous:2009:GRJ

[Ano09b] Anonymous. 7 great reasons for joining the IEEE Computer Society [advertisement]. IEEE Transactions on Software Engineering, 35(6):880, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=5353441.

Anonymous:2009:BCf

[Ano09c]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 35(1):c4, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4771852.

Anonymous:2009:BCc

[Ano09d]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 35(2):c4, March/April 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4809713.

Anonymous:2009:BCe

[Ano09e]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 35(3):c4, May/June 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5061642.

Anonymous:2009:BCd

[Ano09f]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 35(4):c4, July/August 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5186365.

Anonymous:2009:BCb

[Ano09g]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 35(5):c4, September/October 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5275154.

Anonymous:2009:BCa

[Ano09h]  Anonymous.  [back cover].

IEEE Transactions on Software Engineering, 35(6):c4, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5353443.

Anonymous:2009:CPN

[Ano09i] Anonymous. Call for papers for new IEEE Transactions on Affective Computing. IEEE Transactions on Software Engineering, 35(6):879, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5353440.

Anonymous:2009:CPSa

[Ano09j] Anonymous. Call for papers for special issue on software services and service-based systems. IEEE Transactions on Software Engineering, 35(4): 592, July/August 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5186363.

Anonymous:2009:CPSb

[Ano09k]  Anonymous.  Call for pa-

pers: Special issue on software services and servicebased systems. IEEE Transactions on Software Engineering, 35(5):736, September/October 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5275152.

Anonymous:2009:FCf

[Ano09l]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 35(1):c1, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4771846.

Anonymous:2009:FCa

[Ano09m]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 35(2):c1, March/April 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4809709.

Anonymous:2009:FCc

[Ano09n]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 35(3):c1, May/June 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5061639.

Anonymous:2009:FCb

[Ano09o]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 35(4):c1, July/August 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5186359.

Anonymous:2009:FCd

[Ano09p]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 35(5):c1, September/October 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=5275150.

Anonymous:2009:FCe

[Ano09q]  Anonymous.  [front cover].

IEEE Transactions on Software Engineering, 35(6):c1, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=5353436.

Anonymous:2009:ICS

[Ano09r]  Anonymous.  IEEE Com-

puter Society 2009 membership application. IEEE Transactions on Software Engineering, 35(1):142–144, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4771850.

Anonymous:2009:IFCf

[Ano09s]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 35(1):c2, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=4771847.

Anonymous:2009:IFCa

[Ano09t]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 35(2): c2, March/April 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4809710.

Anonymous:2009:IFCc

[Ano09u]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 35 (3):c2, May/June 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5061640.

Anonymous:2009:IFCb

[Ano09v]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 35(4): c2, July/August 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5186360.

Anonymous:2009:IFCd

[Ano09w]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 35(5):c2, September/October 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5275151.

Anonymous:2009:IFCe

[Ano09x]  Anonymous.  [inside front

cover]. IEEE Transactions on Software Engineering, 35 (6):c2, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5353437.

Anonymous:2009:TIAd

[Ano09y] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 35(1):c3, January/ February 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4771851.

Anonymous:2009:TIAf

[Ano09z] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 35(2): c3, March/April 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4809712.

Anonymous:2009:TIAc

[Ano09-27] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 35(3):c3, May/June 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5061641.

Anonymous:2009:TIAa

[Ano09-28] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 35(4): c3, July/August 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5186364.

Anonymous:2009:TIAe

[Ano09-29] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 35(5):c3, September/October 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5275153.

Anonymous:2009:TIAb

[Ano09-30] Anonymous. TSE information for authors. IEEE Transactions on Software Engineering, 35(6):c3, November/ December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5353442.

Anton:2003:FPE

[AP03] A. I. Anton and C. Potts. Functional paleontology: the evolution of user-visible system services. IEEE Transactions on Software Engineering, 29(2): 151–166, February 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1178053.

Ardagna:2007:ASC

[AP07] D. Ardagna and B. Pernici. Adaptive service composition in flexible processes. IEEE Transactions on Software Engineering, 33(6):369– 384, June 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4181707.

Attiogbe:2007:FTE

[APS07] C. Attiogbe, P. Poizat, and G. Salaun. A formal and tool-equipped approach for the integration of state diagrams and formal datatypes. IEEE Transactions on Software Engineering, 33(3):157– 170, March 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4084134.

Andersson:2007:RQA

[AR07] C. Andersson and P. Runeson. A replicated quantitative analysis of fault distributions in complex software systems. IEEE Transactions on Software Engineering, 33 (5):273–286, May 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4160967.

Anderson:2003:DIF

[ART03] P. Anderson, T. Reps, and T. Teitelbaum. Design and implementation of a finegrained software inspection tool. IEEE Transactions on Software Engineering, 29(8): 721–733, August 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1223646.

Arisholm:2004:EED

[AS04] E. Arisholm and Dag I. K.

Sjøberg. Evaluating the effect of a delegated versus centralized control style on the maintainability of object-oriented software. IEEE Transactions on Software Engineering, 30(8): 521–534, August 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1316869.

Anda:2009:VRS

[ASM09] B. C. D. Anda, Dag I. K.

Sjøberg, and A. Mockus. Variability and reproducibility in software engineering: A study of four companies that developed the same system. IEEE Transactions on Software Engineering, 35(3):407– 429, May/June 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4693714.

Andritsos:2005:ITS

[AT05] P. Andritsos and V. Tzerpos. Information-theoretic software clustering. IEEE Transactions on Software Engineering, 31(2): 150–165, February 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1401930.

Auer:2006:OPF

[ATG^{+}06] M. Auer, A. Trendowicz, B. Graser, E. Haunschmid, and S. Biffl. Optimal project feature weights in analogy-based cost estimation: improvement and limitations. IEEE Transactions on Software Engineering, 32(2): 83–92, February 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1599418.

Alagar:2001:TTS

[AV01] S. Alagar and S. Venkatesan. Techniques to tackle state explosion in global predicate detection. IEEE Transactions on Software Engineering, 27(8): 704–714, August 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=940566.

Avritzer:2004:RMP

[AW04] A. Avritzer and E. J. Weyuker.

The role of modeling in the performance testing of ecommerce applications. IEEE Transactions on Software Engineering, 30(12):1072–1083, December 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377198.

Andrews:2003:GTR

[AZ03] J. H. Andrews and Yingjun Zhang. General test result checking with log file analysis. IEEE Transactions on Software Engineering, 29(7): 634–648, July 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1214327.

Breaux:2008:ARR

[BA08] T. Breaux and A. Antón. Analyzing regulatory rules for privacy and security requirements. IEEE Transactions on Software Engineering, 34(1):5–20, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4359472.

Baker:2007:FCD

[Bak07] B. S. Baker. Finding clones with dup: Analysis of an experiment. IEEE Transactions on Software Engineering, 33(9): 608–621, September 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4288194.

Burnett:2000:EHS

[BAvZ00] M. Burnett, A. Agrawal, and P. van Zee. Exception handling in the spreadsheet paradigm. IEEE Transactions on Software Engineering, 26(10):923– 942, October 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=879817.

Briand:2001:CEE

[BBD01] L. C. Briand, C. Bunse, and J. W. Daly. A controlled experiment for evaluating quality guidelines on the maintainability of object-oriented designs. IEEE Transactions on Software Engineering, 27 (6):513–530, June 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=926174.

Bhatti:2006:XFP

[BBG06] R. Bhatti, E. Bertino, and A. Ghafoor. X-FEDERATE: a policy engineering framework for federated access management. IEEE Transactions on Software Engineering, 32 (5):330–346, May 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1642680.

Baik:2002:DCC

[BBS02a] Jongmoon Baik, B. Boehm, and B. M. Steece. Disaggregating and calibrating the CASE tool variable in COCOMO II. IEEE Transactions on Software Engineering, 28(11):1009– 1022, November 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1049401.

Balbo:2002:PFS

[BBS02b] G. Balbo, S. C. Bruell, and M. Sereno. Product form solution for generalized stochastic Petri nets. IEEE Transactions on Software Engineering, 28(10):915– 932, October 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1041049.

Bruns:2003:SPA

[BC03] G. Bruns and S. Chandra.

Searching for points-to analysis. IEEE Transactions on Software Engineering, 29(10):

883–897, October 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1237170.

Ball:2004:EAI

[BCH04] T. Ball, M. Chechik, and D. Hoffman. Editorial: AE introduction. IEEE Transactions on Software Engineering, 30(1):1–2, January 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1265731.

Binkley:2006:TSR

[BCH^{+}06] D. Binkley, M. Ceccato, M. Harman, F. Ricca, and P. Tonella. Tool-supported refactoring of existing objectoriented code into aspects. IEEE Transactions on Software Engineering, 32(9):698– 717, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1707668.

Baier:2007:MCM

[BCH^{+}07] C. Baier, L. Cloth, B. R. Haverkort, M. Kuntz, and M. Siegle. Model checking Markov chains with actions and state labels. IEEE Transactions on Software Engineering, 33(4):209–224, April 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4123324.

Bellavista:2003:CAM

[BCMS03] P. Bellavista, A. Corradi, R. Montanari, and C. Stefanelli. Context-aware middleware for resource management in the wireless internet. IEEE Transactions on Software Engineering, 29(12):1086– 1099, December 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265523.

Bianchi:2003:IRL

[BCMV03] A. Bianchi, D. Caivano, V. Marengo, and G. Visaggio. Iterative reengineering of legacy systems. IEEE Transactions on Software Engineering, 29(3):225– 241, March 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1183932.

Batory:2000:DWV

[BCRW00] D. Batory, Gang Chen, E. Robertson, and Tao Wang. Design wizards and visual programming environments for GenVoca generators. IEEE Transactions on Software Engineering, 26(5):441–452, May 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=846301.

Bang:2001:CMC

[BCY01] Ki-Seok Bang, Jin-Young Choi, and Chuck Yoo. Comments on “The Model Checker SPIN”. IEEE Transactions on Software Engineering, 27(6):573– 576, June 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=926177. See [Hol97].

Bansiya:2002:HMO

[BD02] J. Bansiya and C. G. Davis. A hierarchical model for objectoriented design quality assessment. IEEE Transactions on Software Engineering, 28(1): 4–17, January 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=979986.

Bailey:2003:ADD

[BD03] M. W. Bailey and J. W. Davidson. Automatic detection and diagnosis of faults in generated code for procedure calls. IEEE Transactions on Software Engineering, 29(11):1031– 1042, November 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1245304.

Bowman:2000:SPS

[BDB00] H. Bowman, J. Derrick, and E. Brinksma. Selected papers from the second IFIP int’l conference on formal methods for open object based distributed systems, 1997. IEEE Transactions on Software Engineering, 26(7):577–578, July 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=859528.

Bartoletti:2008:SBD

[BDFZ08] M. Bartoletti, P. Degano, G. L. Ferrari, and R. Zunino. Semantics-based design for secure Web services. IEEE Transactions on Software Engineering, 34(1):33–49, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4359467.

Bohnenkamp:2006:MCM

[BDHK06] H. Bohnenkamp, P. R. D’Argenio, H. Hermanns, and J. P. Katoen. MODEST: A compositional modeling formalism for hard and softly timed systems. IEEE Transactions on Software Engineering, 32(10): 812–830, October 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1717473.

Balsamo:2004:MBP

[BDIS04] S. Balsamo, A. Di Marco, P. Inverardi, and M. Simeoni. Model-based performance prediction in software development: a survey. IEEE Transactions on Software Engineering, 30(5):295–310, May 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1291833.

Briand:2004:AIS

[BDL04] L. C. Briand, M. Di Penta, and Y. Labiche. Assessing and improving state-based class testing: a series of experiments. IEEE Transactions on Software Engineering, 30(11):770– 793, November 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1359770.

Babaoglu:2001:GCP

[BDM01] O. Babaoglu, R. Davoli, and A. Montresor. Group communication in partitionable systems: specification and algorithms. IEEE Transactions on Software Engineering, 27 (4):308–336, April 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=917522.

Briand:2000:CEC

[BEFL00] L. C. Briand, K. El Emam, B. G. Freimut, and O. Laitenberger. A comprehensive evaluation of capture-recapture models for estimating software defect content. IEEE Transactions on Software Engineering, 26(6):518–540, June 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=852741.

Bojic:2004:ALF

[BEK^{+}04] D. Bojic, T. Eisenbarth, R. Koschke, D. Simon, and D. Velasevic. Addendum to “Locating features in source code”. IEEE Transactions on Software Engineering, 30(2): 140–??, February 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265818.  See

[EKS03].

Bergin:2000:SRP

[Ber00] T. Bergin. Sloan research project. IEEE Transactions on Software Engineering, 26 (5):478, May 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=846303.

Baden:2000:PMD

[BF00] S. B. Baden and S. J. Fink.

A programming methodology for dual-tier multicomputers. IEEE Transactions on Software Engineering, 26(3):212– 226, March 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=842948.

Bobbio:2003:PFT

[BFGP03] A. Bobbio, G. Franceschinis, R. Gaeta, and L. Portinale. Parametric fault tree for the dependability analysis of redundant systems and its high-level Petri net semantics. IEEE Transactions on Software Engineering, 29(3): 270–287, March 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1183940.

Bucci:2004:TSS

[BFSV04] G. Bucci, A. Fedeli, L. Sassoli, and E. Vicario. Timed state space analysis of real-time preemptive systems. IEEE Transactions on Software Engineering, 30(2):97–111, February 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1265815.

Buchs:2000:FSF

[BG00] D. Buchs and N. Guelfi. A formal specification framework for object-oriented distributed systems. IEEE Transactions on Software Engineering, 26 (7):635–652, July 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=859532.

Benaroch:2009:IEO

[BG09] M. Benaroch and J. Goldstein.

An integrative economic optimization approach to systems development risk management. IEEE Transactions on Software Engineering, 35(5):638–653, September/October 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4815277.

Bhargavan:2002:VFA

[BGK^{+}02] K. Bhargavan, C. A. Gunter, Moonjoo Kim, Insup Lee, D. Obradovic, O. Sokolsky, and M. Viswanathan. Verisim: formal analysis of network simulations. IEEE Transactions on Software Engineering, 28(2): 129–145, February 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=988495.

Bederson:2004:TDI

[BGM04] B. B. Bederson, J. Grosjean, and J. Meyer. Toolkit design for interactive structured graphics. IEEE Transactions on Software Engineering, 30(8):

535–546, August 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1316870.

Biffl:2003:IDD

[BH03] S. Biffl and M. Halling. Investigating the defect detection effectiveness and cost benefit of nominal inspection teams. IEEE Transactions on Software Engineering, 29(5):385– 397, May 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1199069.

Binkley:2004:AVP

[BH04] D. Binkley and M. Harman. Analysis and visualization of predicate dependence on formal parameters and global variables. IEEE Transactions on Software Engineering, 30(11): 715–735, November 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1359767.

Bode:2009:CDE

[BHH^{+}09] E. Bode, M. Herbstritt, H. Hermanns, S. Johr, T. Peikenkamp, R. Pulungan, J. Rakow, R. Wimmer, and B. Becker. Compositional dependability evaluation for STATEMATE. IEEE Transactions on Software Engineering, 35(2):274– 292, March/April 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4711060.

Baier:2003:MCA

[BHHK03] C. Baier, B. Haverkort, H. Hermanns, and J. P. Katoen. Model-checking algorithms for continuous-time Markov chains. IEEE Transactions on Software Engineering, 29(6):524– 541, June 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1205180.

Barak:2006:IHS

[BHM06] D. Barak, D. Harel, and R. Marelly. InterPlay: Horizontal scale-up and transition to design in scenario-based programming. IEEE Transactions on Software Engineering, 32 (7):467–485, July 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1677533.

Barbier:2003:CAB

[BHS03] F. Barbier and B. Henderson-Sellers. Controversies about the black and white diamonds. IEEE Transactions on Software Engineering, 29(11):1056– ??, November 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1245308.

Bontemps:2005:LSC

[BHS05] Y. Bontemps, P. Heymans, and P. Y. Schobbens. From live sequence charts to state machines and back: a guided tour. IEEE Transactions on Software Engineering, 31(12):999– 1014, December 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1566603.

Barbier:2003:FWP

[BHSLB03] F. Barbier, B. Henderson-Sellers, A. Le Parc-Lacayrelle, and J. M. Bruel. Formalization of the whole-part relationship in the unified modeling language. IEEE Transactions on Software Engineering, 29 (5):459–470, May 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1199074.

Biggerstaff:2004:NAT

[Big04] T. J. Biggerstaff. A new architecture for transformationbased generators. IEEE Transactions on Software Engineering, 30(12):1036–1054, December 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1377196.

Basit:2009:DMA

[BJ09] H. A. Basit and S. Jarzabek. A data mining approach for detecting higherlevel clones in software. IEEE Transactions on Software Engineering, 35(4):497–514, July/ August 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4796208.

Bellon:2007:CEC

[BKA^{+}07] S. Bellon, R. Koschke, G. Antoniol, J. Krinke, and E. Merlo. Comparison and evaluation of clone detection tools. IEEE Transactions on Software Engineering, 33(9):577–591, September 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4288192.

Binkley:2007:GEI

[BKM07] D. Binkley, R. Koschke, and S. Mancoridis. Guest Editors’ introduction to the special section from the international conference on software maintenance and evolution. IEEE Transactions on Software Engineering, 33(12): 797–798, December 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4375378.

Braberman:2005:SMA

[BKO05] V. Braberman, N. Kicillof, and A. Olivero. A scenariomatching approach to the description and model checking of real-time properties. IEEE Transactions on Software Engineering, 31(12):1028– 1041, December 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1566605.

Barat:2002:RIS

[BLD02] F. Barat, R. Lauwereins, and G. Deconinck. Reconfigurable instruction set processors from a hardware/software perspective. IEEE Transactions on Software Engineering, 28(9): 847–862, September 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1033225.

Briand:2005:EIF

[BLDYB05] L. C. Briand, Y. Labiche, M. Di Penta, and H. Yan-Bondoc. An experimental investigation of formality in UML-based development. IEEE Transactions on Software Engineering, 31(10): 833–849, October 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1542066.

Beydoun:2009:FGM

[BLHS^{+}09] G. Beydoun, G. Low, B. Henderson-Sellers, H. Mouratidis, J. J. Gomez-Sanz, J. Pavon, and C. Gonzalez-Perez. FAML: A generic metamodel for MAS development. IEEE Transactions on Software Engineering, 35(6):841–863, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4967615.

Briand:2006:TRE

[BLL06] L. C. Briand, Y. Labiche, and J. Leduc. Toward the reverse engineering of UML sequence diagrams for distributed Java software. IEEE Transactions on Software Engineering, 32(9): 642–663, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1707665.

Briand:2003:IGB

[BLW03] L. C. Briand, Y. Labiche, and Yihong Wang. An investigation of graph-based class integration test order strategies. IEEE Transactions on Software Engineering, 29(7): 594–607, July 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1214324.

Buhr:2000:AEH

[BM00] P. A. Buhr and W. Y. R. Mok. Advanced exception handling mechanisms. IEEE Transactions on Software Engineering, 26(9):820– 836, September 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=877844.

Briand:2002:OPG

[BMB02] L. C. Briand, S. Morasca, and V. R. Basili. An operational process for goaldriven definition of measures. IEEE Transactions on Software Engineering, 28(12):1106– 1125, December 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1158285.

Briand:2002:AAF

[BMW02] L. C. Briand, W. L. Melo, and J. Wust. Assessing the applicability of fault-proneness models across object-oriented software projects. IEEE Transactions on Software Engineering, 28(7):706–720, July 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1019484.

Beyer:2005:ERC

[BNL05] D. Beyer, A. Noack, and C. Lewerentz. Efficient relational calculation for software analysis. IEEE Transactions on Software Engineering, 31(2): 137–149, February 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1401929.

Bellini:2006:RCI

[BNR06] P. Bellini, P. Nesi, and D. Rogai. Reply to comments on “An Interval Logic for Real-Time System Specification”. IEEE Transactions on Software Engineering, 32(6):428– 431, June 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1650217.  See

[MN01, FMPR06].

Bolognesi:2000:TCO

[Bol00] T. Bolognesi. Toward constraintobject-oriented development. IEEE Transactions on Software Engineering, 26(7):594– 616, July 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=859530.

Botting:2005:SET

[Bot05] R. J. Botting. Small errors in “Toward Formalizing Domain Modeling Semantics in Language Syntax”. IEEE Transactions on Software Engineering, 31(10):911, October 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1542071.  See

[EW05].

Boute:2000:SFD

[Bou00] R. Boute. Supertotal function definition in mathematics and software engineering. IEEE Transactions on Software Engineering, 26(7):662– 672, July 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=859534.

Brylow:2004:DAI

[BP04] D. Brylow and J. Palsberg. Deadline analysis of interrupt-driven software. IEEE Transactions on Software Engineering, 30(10):634– 655, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339276.

Bobbio:2000:MFI

[BPT00] A. Bobbio, A. Puliafito, and M. Tekel. A modeling framework to implement preemption policies in non-Markovian SPNs. IEEE Transactions on Software Engineering, 26(1):

36–54, January 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=825765.

Berling:2003:EEM

[BR03] T. Berling and P. Runeson. Efficient evaluation of multifactor dependent system performance using fractional factorial design. IEEE Transactions on Software Engineering, 29(9): 769–781, September 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1232283.

Braunl:2000:PIA

[Bra00]  T. Braunl.  Parallaxis-III:

architecture-independent data parallel processing. IEEE Transactions on Software Engineering, 26(3):227–243, March 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=842949.

Broy:2001:TMF

[Bro01] M. Broy. Toward a mathematical foundation of software engineering methods. IEEE Transactions on Software Engineering, 27(1):42– 57, January 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=895987.

Borger:2000:PMS

[BS00] E. Borger and W. Schulte. A practical method for specification and analysis of exception handling — a Java/JVM case study. IEEE Transactions on Software Engineering, 26(9): 872–887, September 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=877847.

Boland:2003:CPR

[BSC03] P. J. Boland, H. Singh, and B. Cukic. Comparing partition and random testing via majorization and Schur functions. IEEE Transactions on Software Engineering, 29(1): 88–94, January 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1166591.

Ben-Shaul:2001:DAD

[BSHL01] I. Ben-Shaul, O. Holder, and B. Lavva. Dynamic adaptation and deployment of distributed components in hadas. IEEE Transactions on Software Engineering, 27(9):769– 787, September 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=950315.

Batory:2004:SSW

[BSR04] D. Batory, J. N. Sarvela, and A. Rauschmayer. Scaling step-wise refinement. IEEE Transactions on Software Engineering, 30(6):355–371, June 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1321059.

Burnett:2002:THS

[BSRR02] M. Burnett, A. Sheretov, Bing Ren, and G. Rothermel. Testing homogeneous spreadsheet grids with the “what you see is what you test” methodology. IEEE Transactions on Software Engineering, 28(6): 576–594, June 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1010060.

Bucci:2005:CVP

[BSV05] G. Bucci, L. Sassoli, and E. Vicario. Correctness verification and performance analysis of real-time systems using stochastic preemptive time Petri nets. IEEE Transactions on Software Engineering, 31(11):913– 927, November 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1556551.

Berry:2003:CFM

[BT03] D. M. Berry and W. F. Tichy. Comments on “Formal methods application: an empirical tale of software development”. IEEE Transactions on Software Engineering, 29 (6):567–571, June 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1205183. See [SC02]

and response [SC03].

Bruntink:2005:UCD

[BvDvET05] M. Bruntink, A. van Deursen, R. van Engelen, and T. Tourwe. On the use of clone detection for identifying crosscutting concern code. IEEE Transactions on Software Engineering, 31(10):804–818, October 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1542064.

Bandi:2003:PMP

[BVT03] R. K. Bandi, V. K. Vaishnavi, and D. E. Turk. Predicting maintenance performance using object-oriented design complexity metrics. IEEE Transactions on Software Engineering, 29(1):77– 87, January 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1166590.

Briand:2001:MDE

[BW01] L. C. Briand and J. Wust.

Modeling development effort in object-oriented systems using design properties. IEEE Transactions on Software Engineering, 27(11):963– 986, November 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=965338.

Bratthall:2002:IPD

[BW02] L. Bratthall and C. Wohlin. Is it possible to decorate graphical software design and architecture models with qualitative information? — An experiment. IEEE Transactions on Software Engineering, 28(12):1181– 1193, December 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1158290.

Baumstark:2005:RSI

[BW05] L. B. Baumstark and L. M.

Wills. Retargeting sequential image-processing programs for data parallel execution. IEEE Transactions on Software Engineering, 31(2):116– 136, February 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1401928.

Baghdadi:2002:CPE

[BZCJ02] A. Baghdadi, N. E. Zergainoh, W. O. Cesario, and A. A. Jerraya. Combining a performance estimation methodology with a hardware/software codesign flow supporting multiprocessor systems. IEEE Transactions on Software Engineering, 28(9): 822–831, September 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1033223.

Canfora:2003:GEI

[CA03] G. Canfora and A. A. Andrews. Guest Editors’ introduction: 2001 international conference on software maintenance. IEEE Transactions on Software Engineering, 29 (3):193–194, March 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1183926.

Chan:2001:OSM

[CAB^{+}01] W. Chan, R. J. Anderson, P. Beame, D. H. Jones, D. Notkin, and W. E. Warner. Optimizing symbolic model checking for statecharts. IEEE Transactions on Software Engineering, 27(2):170–190, February 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=908961.

Camp:2001:AGP

[Cam01] L. J. Camp. An atomicitygenerating protocol for anonymous currencies. IEEE Transactions on Software Engineering, 27(3):272–278, March 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=910862.

Casale:2009:CEC

[Cas09]  G. Casale.  CoMoM: Effi-

cient class-oriented evaluation of multiclass performance models. IEEE Transactions on Software Engineering, 35(2):162– 177, March/April 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4641939.

Chan:2003:MRM

[CC03] A. T. S. Chan and Siu-Nam Chuang. MobiPADS: a reflective middleware for contextaware mobile computing. IEEE Transactions on Software Engineering, 29(12):1072–1085, December 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265522.

Chuang:2008:DQA

[CC08] S. N. Chuang and A. T. S.

Chan. Dynamic QoS adaptation for mobile middleware. IEEE Transactions on Software Engineering, 34(6): 738–752, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4547430.

Casati:2000:UPD

[CCF^{+}00] F. Casati, S. Castano, M. Fugini, I. Mirbel, and B. Pernici. Using patterns to design rules in workflows. IEEE Transactions on Software Engineering, 26(8):760– 785, August 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=879813.

Chaki:2004:MVS

[CCG^{+}04] S. Chaki, E. M. Clarke, A. Groce, S. Jha, and H. Veith. Modular verification of software components in C. IEEE Transactions on Software Engineering, 30(6): 388–402, June 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1321061.

Chamberlain:2000:ZMI

[CCL^{+}00] B. L. Chamberlain, Sung-Eun Choi, C. Lewis, C. Lin, L. Snyder, and W. D. Weathersby. ZPL: a machine independent programming language for parallel computers.

IEEE Transactions on Software Engineering, 26(3):197– 211, March 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=842947.

Cavusoglu:2007:EVD

[CCR07] H. Cavusoglu, H. Cavusoglu, and S. Raghunathan. Efficiency of vulnerability disclosure mechanisms to disseminate vulnerability knowledge. IEEE Transactions on Software Engineering, 33(3):171– 185, March 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4084135.

Cheng:2000:GEI

[CCW00a] A. M. K. Cheng, P. Clements, and M. Woodside. Guest Editors’ introduction: workshop on software and performance. IEEE Transactions on Software Engineering, 26(11):1025– 1026, November 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=881715.

Cheng:2000:WSP

[CCW00b] A. M. K. Cheng, P. Clements, and M. Woodside. Workshop on software and performance [Guest Editors’ introduction]. IEEE Transactions on Software Engineering, 26 (12):1121, December 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=888626.

Caporuscio:2003:DES

[CCW03] M. Caporuscio, A. Carzaniga, and A. L. Wolf. Design and evaluation of a support service for mobile, wireless publish/subscribe applications. IEEE Transactions on Software Engineering, 29(12):1059– 1071, December 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1265521.

Cugola:2001:JEB

[CDF01] G. Cugola, E. Di Nitto, and A. Fuggetta. The JEDI eventbased infrastructure and its application to the development of the OPSS WFMS. IEEE Transactions on Software Engineering, 27(9):827– 850, September 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=950318.

Costagliola:2006:CMC

[CDFG06] G. Costagliola, V. Deufemia, F. Ferrucci, and C. Gravino. Constructing meta-CASE workbenches by exploiting visual language generators. IEEE Transactions on Software Engineering, 32(3):156–175, March 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1610608.

Cangussu:2002:FMS

[CDM02] J. W. Cangussu, R. A. De-Carlo, and A. P. Mathur. A formal model of the software test process. IEEE Transactions on Software Engineering, 28(8): 782–796, August 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1027800.

Cangussu:2003:USA

[CDM03] J. W. Cangussu, R. A. De-Carlo, and A. P. Mathur. Using sensitivity analysis to validate a state variable model of the software test process. IEEE Transactions on Software Engineering, 29(5):430– 443, May 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1199072.

Cysneiros:2004:NRE

[CdPL04] L. M. Cysneiros and J. C. S.

do Prado Leite. Nonfunctional requirements: from elicitation to conceptual models. IEEE Transactions on Software Engineering, 30(5):

328–350, May 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1291835.

Cohen:2008:CIT

[CDS08] M. B. Cohen, M. B. Dwyer, and J. Shi. Constructing interaction test suites for highlyconfigurable systems in the presence of constraints: A greedy approach. IEEE Transactions on Software Engineering, 34(5):633–650, September/October 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4564473.

Capra:2003:CCA

[CEM03] L. Capra, W. Emmerich, and C. Mascolo. CARISMA: context-aware reflective middleware system for mobile applications. IEEE Transactions on Software Engineering, 29(10):929–945, October 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1237173.

Chen:2001:RDU

[CFC01] Shyh-Kwei Chen, W. K. Fuchs, and Jen-Yao Chung. Reversible debugging using program instrumentation. IEEE Transactions on Software Engineering, 27(8):715–727, August 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=940726.

Capra:2008:ESR

[CFM08] E. Capra, C. Francalanci, and F. Merlo. An empirical study on the relationship between software design quality, development effort and governance in open source projects. IEEE Transactions on Software Engineering, 34(6):765–782, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4599582.

Canal:2003:ARC

[CFP^{+}03] C. Canal, L. Fuentes, E. Pimentel, J. M. Troya, and A. Vallecillo. Adding roles to CORBA objects. IEEE Transactions on Software Engineering, 29(3):242–260, March 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1183935.

Costagliola:2005:CPA

[CFTV05] G. Costagliola, F. Ferrucci, G. Tortora, and G. Vitiello. Class point: an approach for the size estimation of object-oriented systems. IEEE Transactions on Software Engineering, 31(1): 52–74, January 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1392720.

Chechik:2001:AAC

[CG01] M. Chechik and J. Gannon. Automatic analysis of consistency between requirements and designs. IEEE Transactions on Software Engineering, 27(7):651–672, July 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=935856. See [JY06].

Ciardo:2002:ISS

[CGH02] G. Ciardo, R. German, and B. R. Haverkort. Introduction to the special section on Petri nets and performance models. IEEE Transactions on Software Engineering, 28(10): 913–914, October 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1041048.

Cortellessa:2005:MBP

[CGPA^{+}05] V. Cortellessa, K. Goseva-Popstojanova, Kalaivani Appukkutty, A. R. Guedem, A. Hassan, R. Elnaggar, W. Abdelmoez, and H. H. Ammar. Model-based performance risk analysis. IEEE Transactions on Software Engineering, 31(1):3–20, January 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1392717.

Carnevali:2009:SDF

[CGV09] L. Carnevali, L. Grassi, and E. Vicario. State-density functions over DBM domains in the analysis of non-Markovian models. IEEE Transactions on Software Engineering, 35(2): 178–194, March/April 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4711059.

Chari:2006:STP

[CH06] K. Chari and A. Hevner. System test planning of software: An optimization approach. IEEE Transactions on Software Engineering, 32 (7):503–509, July 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1677535.

Chittimalli:2009:RCI

[CH09] P. K. Chittimalli and M. J.

Harrold. Recomputing coverage information to assist regression testing. IEEE Transactions on Software Engineering, 35(4):452–469, July/ August 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4760153.

Chavez:2000:DAS

[Cha00] T. Chavez. A decision-analytic stopping rule for validation of commercial software systems. IEEE Transactions on Software Engineering, 26(9):907– 918, September 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=877849.

Chatzigeorgiou:2003:MAO

[Cha03] A. Chatzigeorgiou. Mathematical assessment of objectoriented design quality. IEEE Transactions on Software Engineering, 29(11):1050–1053, November 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1245306.

Cleland-Huang:2003:EBT

[CHCC03] J. Cleland-Huang, C. K. Chang, and M. Christensen. Event-based traceability for managing evolutionary change. IEEE Transactions on Software Engineering, 29(9):796– 810, September 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1232285.

Cleland-Huang:2008:GCT

[CHMB08] J. Cleland-Huang, W. Marrero, and B. Berenbach. Goalcentric traceability: Using virtual plumblines to maintain critical systemic qualities. IEEE Transactions on Software Engineering, 34(5):685–699, September/October 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4553719.

Chen:2003:COS

[CJ03] J. C. Chen and H. C. Jiau.

Comments on “On object systems and behavioral inheritance”. IEEE Transactions on Software Engineering, 29 (6):576–??, June 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1205185.  See

[HK02].

Constant:2007:IFV

[CJMR07] C. Constant, T. Jéron, H. Marchand, and V. Rusu. Integrating formal verification and conformance testing for reactive systems. IEEE Transactions on Software Engineering, 33(8): 558–574, August 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4267026.

Chae:2004:ICM

[CKB04] Heung Seok Chae, Yong Rae Kwon, and Doo Hwan Bae. Improving cohesion metrics for classes by considering dependent instance variables. IEEE Transactions on Software Engineering, 30(11):826– 832, November 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1359773.

Cabri:2002:EMA

[CLZ02] G. Cabri, L. Leonardi, and F. Zambonelli. Engineering mobile agent applications via context-dependent coordination. IEEE Transactions on Software Engineering, 28(11):1039–1055, November 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1049403.

Cuadrado:2009:MBA

[CM09] J. Sanchez Cuadrado and J. G. Molina. A modelbased approach to families of embedded domain-specific languages. IEEE Transactions on Software Engineering, 35(6):825–840, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=4782971.

Concas:2007:PLL

[CMPS07] G. Concas, M. Marchesi, S. Pinna, and N. Serra. Powerlaws in a large object-oriented software system. IEEE Transactions on Software Engineering, 33(10):687–708, October 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4302780.

Cataldo:2009:SDW

[CMRH09] M. Cataldo, A. Mockus, J. A.

Roberts, and J. D. Herbsleb. Software dependencies, work dependencies, and their impact on failures. IEEE Transactions on Software Engineering, 35(6):864–878, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5166450.

Cubranic:2005:HPM

[CMSB05] D. Cubranic, G. C. Murphy, J. Singer, and K. S. Booth. Hipikat: a project memory for software development. IEEE Transactions on Software Engineering, 31 (6):446–465, June 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1463229.

Cavalcanti:2000:WPS

[CN00] A. Cavalcanti and D. A. Naumann. A weakest precondition semantics for refinement of object-oriented programs. IEEE Transactions on Software Engineering, 26(8): 713–728, August 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=879810.

Carver:2008:IEB

[CNP08] J. C. Carver, N. Nagappan, and A. Page. The impact of educational background on the effectiveness of requirements inspections: An empirical study. IEEE Transactions on Software Engineering, 34(6):800–812, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4564472.

Cortellessa:2007:ISM

[CPR07] V. Cortellessa, P. Pierini, and D. Rossi. Integrating software models and platform models for performance analysis. IEEE Transactions on Software Engineering, 33(6):385– 401, June 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4181708.

Canal:2008:MBA

[CPS08] C. Canal, P. Poizat, and G. Salaún. Model-based adaptation of behavioral mismatching components. IEEE Transactions on Software Engineering, 34(4):546–563, July/ August 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4527252.

Chen:2003:CRF

[CPT03] T. Y. Chen, Pak-Lok Poon, and T. H. Tse. A choice relation framework for supporting category-partition test case generation. IEEE Transactions on Software Engineering, 29 (7):577–593, July 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1214323.

Chang:2008:DNC

[CPY08] R. Y. Chang, A. Podgurski, and J. Yang. Discovering neglected conditions in software by mining dependence graphs. IEEE Transactions on Software Engineering, 34(5):579–596, September/October 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4492791.

Chatterjee:2001:CPA

[CRL01] R. Chatterjee, B. G. Ryder, and W. A. Landi. Complexity of points-to analysis of Java in the presence of exceptions. IEEE Transactions on Software Engineering, 27 (6):481–512, June 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=926173.

Cartwright:2000:EIO

[CS00] M. Cartwright and M. Shepperd. An empirical investigation of an object-oriented software system. IEEE Transactions on Software Engineering, 26(8):786–796, August 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=879814.

Collberg:2002:WTP

[CT02] C. S. Collberg and C. Thomborson. Watermarking, tamperproofing, and obfuscation — tools for software protection. IEEE Transactions on Software Engineering, 28(8):735– 746, August 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1027797.

Canfora:2009:GEI

[CTM09] G. Canfora, L. Tahvildari, and H. A. Muller. Guest Editors’ introduction to the special section from the international conference on software maintenance. IEEE Transactions on Software Engineering, 35(4):450–451, July/ August 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5186362.

Chen:2000:ATP

[CW00] Ying Chen and M. Winslett.

Automated tuning of parallel I/O systems: an approach to portable I/O performance for scientific applications. IEEE Transactions on Software Engineering, 26 (4):362–383, April 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=844494.

Chen:2002:EDO

[CW02a] Loon-Been Chen and I-Chen Wu. An efficient distributed online algorithm to detect strong conjunctive predicates. IEEE Transactions on Software Engineering, 28(11):1077– 1084, November 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1049405.

Cheng:2002:FID

[CW02b] B. H. C. Cheng and E. Y. Wang. Formalizing and integrating the dynamic model for object-oriented modeling. IEEE Transactions on Software Engineering, 28(8):747– 762, August 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1027798.

Coppit:2005:SAB

[CYK^{+}05] D. Coppit, Jinlin Yang, S. Khurshid, Wei Le, and K. Sullivan. Software assurance by bounded exhaustive testing. IEEE Transactions on Software Engineering, 31(4): 328–339, April 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1435353.

Cornelissen:2009:SSP

[CZvD^{+}09] B. Cornelissen, A. Zaidman, A. van Deursen, L. Moonen, and R. Koschke. A systematic survey of program comprehension through dynamic analysis. IEEE Transactions on Software Engineering, 35(5):684–702, September/October 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4815280.

Dzidek:2008:REE

[DAB08] W. J. Dzidek, E. Arisholm, and L. C. Briand. A realistic empirical evaluation of the costs and benefits of UML in software maintenance. IEEE Transactions on Software Engineering, 34(3):407– 432, May/June 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4459340.

deAlfaro:2009:LBS

[dAFS09] L. de Alfaro, M. Faella, and M. Stoelinga. Linear and branching system metrics. IEEE Transactions on Software Engineering, 35(2):258– 273, March/April 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4721438.

Diaz:2001:CP

[DAP01a] P. Diaz, I. Aedo, and F. Panetsos. Call for papers. IEEE Transactions on Software Engineering, 27(8):766–767, August 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=940729.

Diaz:2001:MDB

[DAP01b] P. Diaz, I. Aedo, and F. Panetsos. Modeling the dynamic behavior of hypermedia applications. IEEE Transactions on Software Engineering, 27 (6):550–572, June 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=926176.

diBlasio:2000:CFA

[dBFT00] P. di Blasio, K. Fisher, and C. Talcott. A control-flow analysis for a calculus of concurrent objects. IEEE Transactions on Software Engineering, 26(7):617–634, July 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=859531.

Damian:2006:ESC

[DC06] D. Damian and J. Chisan. An empirical study of the complex relationships between requirements engineering processes and other processes that lead to payoffs in productivity, quality, and risk management. IEEE Transactions on Software Engineering, 32 (7):433–453, July 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1677531.

Deavours:2002:MFI

[DCC^{+}02] D. D. Deavours, G. Clark, T. Courtney, D. Daly, S. Derisavi, J. M. Doyle, W. H.

Sanders, and P. G. Webster. The mobius framework and its implementation. IEEE Transactions on Software Engineering, 28(10):956– 969, October 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1041052.

DeLine:2001:APM

[DeL01a] R. DeLine. Avoiding packaging mismatch with flexible packaging. IEEE Transactions on Software Engineering, 27(2): 124–143, February 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=908958. See correc-

tions [DeL01b].

DeLine:2001:CAP

[DeL01b]  R. DeLine.  Corrections

to “Avoiding packaging mismatch with flexible packaging”. IEEE Transactions on Software Engineering, 27(6): 577, June 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=926178. See [DeL01a].

Delgado:2004:TCR

[DGR04] N. Delgado, A. Q. Gates, and S. Roach. A taxonomy and catalog of runtime software-fault monitoring tools.

IEEE Transactions on Software Engineering, 30(12):859– 872, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377185.

DeLoach:2000:TBR

[DH00] S. A. DeLoach and T. C. Hartrum. A theory-based representation for object-oriented domain models. IEEE Transactions on Software Engineering, 26(6):500–517, June 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=852740.

Dolado:2003:EII

[DHOH03] J. J. Dolado, M. Harman, M. C. Otero, and L. Hu. An empirical investigation of the influence of a type of side effects on program comprehension. IEEE Transactions on Software Engineering, 29(7): 665–670, July 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1214329.

Dong:2008:TAP

[DHQ^{+}08] J. S. Dong, P. Hao, S. Qin, J. Sun, and W. Yi. Timed automata patterns. IEEE Transactions on Software Engineering, 34(6):844–859, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4586397.

Donatelli:2009:MCT

[DHS09] S. Donatelli, S. Haddad, and J. Sproston. Model checking timed and stochastic properties with CSL^{TA}. IEEE Transactions on Software Engineering, 35(2):224– 240, March/April 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4721440.

Drummond:2000:LAA

[DIH00] C. G. Drummond, D. Ionescu, and R. C. Holte. A learning agent that assists the browsing of software libraries. IEEE Transactions on Software Engineering, 26(12):1179– 1196, December 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=888631.

deJonge:2005:BLC

[dJ05] M. de Jonge. Build-level components. IEEE Transactions on Software Engineering, 31 (7):588–600, July 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1492373.

Darcy:2005:SCS

[DKST05] D. P. Darcy, C. F. Kemerer, S. A. Slaughter, and J. E. Tomayko. The structural complexity of software an experimental test. IEEE Transactions on Software Engineering, 31(11):982– 995, November 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1556556.

Ducasse:2005:CBV

[DL05] S. Ducasse and M. Lanza. The class blueprint: visually supporting the understanding of glasses. IEEE Transactions on Software Engineering, 31 (1):75–90, January 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1392721.

deLeon:2006:HID

[dLAF06] D. Conte de Leon and J. Alves-Foss. Hidden implementation dependencies in high assurance and critical computing systems. IEEE Transactions on Software Engineering, 32(10): 790–811, October 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1717472.

Damas:2005:GAB

[DLDvL05] C. Damas, B. Lambeau, P. Dupont, and A. van Lamsweerde. Generating annotated behavior models from end-user scenarios. IEEE Transactions on Software Engineering, 31(12):1056–1073, December 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1566607.

DAmbros:2009:VCC

[DLL09] M. D’Ambros, M. Lanza, and M. Lungu. Visualizing cochange information with the evolution radar. IEEE Transactions on Software Engineering, 35(5):720–735, September/October 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4815274.

Damian:2008:NMM

[DLM08a] D. Damian, F. Lanubile, and T. Mallardo. On the need for mixed media in distributed requirements negotiations. IEEE Transactions on Software Engineering, 34(1):116–132, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4378346.

dAmorim:2008:DEE

[dLM08b] M. d’Amorim, S. Lauterburg, and D. Marinov. Delta execution for efficient state-space exploration of object-oriented programs. IEEE Transactions on Software Engineering, 34(5):597–613, September/October 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4528965.

Duraes:2006:ESF

[DM06] J. A. Duraes and H. S. Madeira. Emulation of software faults: A field data study and a practical approach. IEEE Transactions on Software Engineering, 32(11):849– 867, November 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4015509.

Dingsoyr:2008:IEP

[DM08] T. Dingsoyr and N. B. Moe. The impact of employee participation on the use of an electronic process guide: A longitudinal case study. IEEE Transactions on Software Engineering, 34(2):212– 225, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4384504.

Desai:2005:IPD

[DMCS05] N. Desai, A. U. Mallya, A. K. Chopra, and M. P. Singh. Interaction protocols as design abstractions for business processes. IEEE Transactions on Software Engineering, 31(12):1015–1027, December 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1566604.

Dig:2008:ESM

[DMJN08] D. Dig, K. Manzoor, R. E.

Johnson, and T. N. Nguyen. Effective software merging in the presence of object-oriented refactorings. IEEE Transactions on Software Engineering, 34(3):321–335, May/June 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4509441.

Delamaro:2001:IMA

[DMM01] M. E. Delamaro, J. C. Maidonado, and A. P. Mathur. Interface mutation: an approach for integration testing. IEEE Transactions on Software Engineering, 27(3): 228–247, March 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=910859.

Dobrica:2002:SSA

[DN02] L. Dobrica and E. Niemela. A survey on software architecture analysis methods. IEEE Transactions on Software Engineering, 28(7):638– 653, July 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1019479.

Dolado:2000:VCB

[Dol00] J. J. Dolado. A validation of the component-based method for software size estimation. IEEE Transactions on Software Engineering, 26(10):1006– 1021, October 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=879821.

Ducasse:2009:SAR

[DP09] S. Ducasse and D. Pollet. Software architecture reconstruction: A process-oriented taxonomy. IEEE Transactions on Software Engineering, 35(4):573–591, July/ August 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4815276.

Devanbu:2000:GEI

[DPP00] P. T. Devanbu, D. E. Perry, and J. S. Poulin. Guest Editors introduction: next generation software reuse. IEEE Transactions on Software Engineering, 26(5):423–424, May 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=846299.

Demsky:2006:GDR

[DR06a] B. Demsky and M. C. Rinard. Goal-directed reasoning for specification-based data structure repair. IEEE Transactions on Software Engineering, 32(12):931–951, December 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4016571.

Do:2006:UMF

[DR06b] Hyunsook Do and G. Rothermel. On the use of mutation faults in empirical assessments of test case prioritization techniques. IEEE Transactions on Software Engineering, 32(9): 733–752, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1707670.

Demsky:2009:AEH

[DR09] B. Demsky and M. Rinard. Automatic extraction of heap reference properties in object-oriented programs. IEEE Transactions on Software Engineering, 35(3):305– 324, May/June 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4693716.

Dunsmore:2003:DET

[DRW03] A. Dunsmore, M. Roper, and M. Wood. The development and evaluation of three diverse techniques for object-oriented code inspection. IEEE Transactions on Software Engineering, 29(8): 677–686, August 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1223643.

Devanbu:2000:CVT

[DS00] P. T. Devanbu and S. G. Stubblebine. Cryptographic verification of test coverage claims. IEEE Transactions on Software Engineering, 26(2):178– 192, February 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=841116.

Devanbu:2002:SQI

[DS02] P. T. Devanbu and S. G. Stubblebine. Stack and queue integrity on hostile platforms. IEEE Transactions on Software Engineering, 28(1):100– 108, January 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=979991.

Dillon:2003:IGC

[DS03] L. K. Dillon and R. E. K.

Stirewalt. Inference graphs: a computational structure supporting generation of customizable and correct analysis components. IEEE Transactions on Software Engineering, 29(2): 133–150, February 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1178052.

Dillon:2004:GEI

[DT04] L. K. Dillon and W. F.

Tichy. Guest Editors’ introduction: 2003 international conference on software engineering. IEEE Transactions on Software Engineering, 30 (6):353–354, June 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1321058.

Dinh-Trong:2005:FPR

[DTB05] T. T. Dinh-Trong and J. M.

Bieman. The FreeBSD project: a replication case study of open source development. IEEE Transactions on Software Engineering, 31 (6):481–494, June 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1463231.

Desmet:2008:PPA

[DVJP08] L. Desmet, P. Verbaeten, W. Joosen, and F. Piessens. Provable protection against Web application vulnerabilities related to session data dependencies. IEEE Transactions on Software Engineering, 34(1):50–64, January/ February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4359468.

Dai:2007:UAS

[DXLN07] Y. S. Dai, M. Xie, Q. Long, and S. H. Ng. Uncertainty analysis in software reliability modeling by Bayesian analysis with maximum-entropy principle. IEEE Transactions on Software Engineering, 33(11): 781–795, November 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4339233.

Devanbu:2008:ISS

[DY08a] P. Devanbu and M. Young. Introduction to the special section on the ACM SIG-SOFT Foundations of Software Engineering Conference. IEEE Transactions on Software Engineering, 34(4):433, July/August 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4585925.

Dolev:2008:TSS

[DY08b] S. Dolev and R. Yagel. Towards self-stabilizing operating systems. IEEE Transactions on Software Engineering, 34(4):564–576, July/ August 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4553720.

Dyba:2005:EIK

[Dyb05] T. Dyba. An empirical investigation of the key factors for success in software process improvement. IEEE Transactions on Software Engineering, 31(5):410–424, May 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1438376.

Dong:2007:VDP

[DYZ07] J. Dong, S. Yang, and K. Zhang. Visualizing design patterns in their applications and compositions. IEEE Transactions on Software Engineering, 33(7):433– 453, July 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4227827.

Escalona:2008:NMD

[EA08] M. J. Escalona and G. Aragón. NDT. A model-driven approach for Web requirements. IEEE Transactions on Software Engineering, 34(3):377– 390, May/June 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4497213.

ElEmam:2000:VII

[EB00] K. El Emam and A. Birk. Validating the ISO/IEC 15504 measure of software requirements analysis process capability. IEEE Transactions on Software Engineering, 26(6): 541–566, June 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=852742.

ElEmam:2002:OCS

[EBG^{+}02] K. El Emam, S. Benlarbi, N. Goel, W. Melo, H. Lounis, and S. N. Rai. The optimal class size for object-oriented software. IEEE Transactions on Software Engineering, 28 (5):494–509, May 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1000452.

ElEmam:2001:CEC

[EBGR01] K. El Emam, S. Benlarbi, N. Goel, and S. N. Rai. The confounding effect of class size on the validity of object-oriented metrics. IEEE Transactions on Software Engineering, 27(7):630– 650, July 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=935855. See com-

ments [Eva03].

Ernst:2002:EAC

[EBN02] M. D. Ernst, G. J. Badros, and D. Notkin. An empirical analysis of c preprocessor use. IEEE Transactions on Software Engineering, 28(12):1146– 1170, December 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1158288.

Elbaum:2009:CRD

[ECDJ09] S. Elbaum, H. N. Chin, M. B.

Dwyer, and M. Jorde. Carving and replaying differential unit test cases from system test cases. IEEE Transactions on Software Engineering, 35(1):29–45, January/ February 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4711061.

Ernst:2001:DDL

[ECGN01] M. D. Ernst, J. Cockrell, W. G. Griswold, and D. Notkin. Dynamically discovering likely program invariants to support program evolution. IEEE Transactions on Software Engineering, 27(2):99–123, February 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=908957.

Elbaum:2005:PDS

[ED05] S. Elbaum and M. Diep. Profiling deployed software: assessing strategies and testing opportunities. IEEE Transactions on Software Engineering, 31(4):312–327, April 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1435352.

El-Fakih:2004:FBI

[EFYB04] K. El-Fakih, N. Yevtushenko, and G. Bochmann. FSM-based incremental conformance testing methods. IEEE Transactions on Software Engineering, 30(7):425–436, July 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1318604.

Eick:2001:DCD

[EGK^{+}01] S. G. Eick, T. L. Graves, A. F. Karr, J. S. Marron, and A. Mockus. Does code decay? Assessing the evidence from change management data. IEEE Transactions on Software Engineering, 27 (1):1–12, January 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=895984.

Eick:2002:VSC

[EGK^{+}02] S. G. Eick, T. L. Graves, A. F. Karr, A. Mockus, and P. Schuster. Visualizing software changes. IEEE Transactions on Software Engineering, 28(4):396–412, April 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=995435.

Egyed:2003:SDA

[Egy03] A. Egyed. A scenario-driven approach to trace dependency analysis. IEEE Transactions on Software Engineering, 29(2): 116–132, February 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1178051.

Eom:2001:THT

[EH01] Hyeonsang Eom and J. K.

Hollingsworth. A tool to help tune where computation is performed. IEEE Transactions on Software Engineering, 27 (7):618–629, July 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=935854.

Eisenbarth:2003:LFS

[EKS03] T. Eisenbarth, R. Koschke, and D. Simon. Locating features in source code. IEEE Transactions on Software Engineering, 29(3):210– 224, March 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1183929. See adden-

dum [BEK^{+}04].

ElEmam:2001:ECR

[EL01] K. El Emam and O. Laitenberger. Evaluating capturerecapture models with two inspectors. IEEE Transactions on Software Engineering, 27(9): 851–864, September 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=950319.

Elbaum:2002:TCP

[EMR02] S. Elbaum, A. G. Malishevsky, and G. Rothermel. Test case prioritization: a family of empirical studies. IEEE Transactions on Software Engineering, 28(2):159– 182, February 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=988497.

Erdogmus:2005:ETF

[EMT05] H. Erdogmus, M. Morisio, and M. Torchiano. On the effectiveness of the test-first approach to programming. IEEE Transactions on Software Engineering, 31(3):226– 237, March 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1423994.

En-Nouaary:2002:TWM

[ENDK02] A. En-Nouaary, R. Dssouli, and F. Khendek. Timed wp-method: testing realtime systems. IEEE Transactions on Software Engineering, 28(11):1023–1038, November 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1049402.

Elbaum:2008:GEI

[ER08] S. Elbaum and D. S. Rosenblum. Guest Editors’ introduction: International symposium on software testing and analysis. IEEE Transactions on Software Engineering, 34(5):577–578, September/October 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4636876.

Elbaum:2005:LUS

[ERKF05] S. Elbaum, G. Rothermel, S. Karre, and M. Fisher II. Leveraging user-session data to support Web application testing. IEEE Transactions on Software Engineering, 31(3): 187–202, March 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1423991.

Edwards:2004:CCW

[ESWH04] S. H. Edwards, M. Sitaraman, B. W. Weide, and E. Hollingsworth. Contractchecking wrappers for C++ classes. IEEE Transactions on Software Engineering, 30(11): 794–810, November 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1359771.

Evanco:2003:CCE

[Eva03] W. M. Evanco. Comments on “The confounding effect of class size on the validity of object-oriented metrics”. IEEE Transactions on Software Engineering, 29(7): 670–672, July 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?

arnumber=1214331.  See

[EBGR01].

Eshuis:2004:TSV

[EW04] R. Eshuis and R. Wieringa.

Tool support for verifying UML activity diagrams. IEEE Transactions on Software Engineering, 30(7):437–447, July 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1318605.

Evermann:2005:TFD

[EW05] J. Evermann and Y. Wand. Toward formalizing domain modeling semantics in language syntax. IEEE Transactions on Software Engineering, 31 (1):21–37, January 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1392718. See errata

[OFdL^{+}04].

Egyed:2006:SMD

[EW06] A. Egyed and D. S. Wile. Support for managing design-time decisions. IEEE Transactions on Software Engineering, 32 (5):299–314, May 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1642678.

Eaddy:2008:DCC

[EZS^{+}08] M. Eaddy, T. Zimmermann, K. D. Sherwood, V. Garg, G. C. Murphy, N. Nagappan, and A. V. Aho. Do crosscutting concerns cause defects? IEEE Transactions on Software Engineering, 34(4):497– 515, July/August 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4527257.

Franks:2009:EMS

[FAOW^{+}09] G. Franks, T. Al-Omari, M. Woodside, O. Das, and S. Derisavi. Enhanced modeling and solution of layered queueing networks. IEEE Transactions on Software Engineering, 35(2):148–161, March/ April 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4620121.

Fu:2005:SCA

[FBS05] X. Fu, T. Bultan, and J. Su. Synchronizability of conversations among Web services. IEEE Transactions on Software Engineering, 31(12):1042– 1055, December 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1566606.

Freimut:2005:DIC

[FBV05] B. Freimut, L. C. Briand, and F. Vollei. Determining inspection cost-effectiveness by combining project data and expert opinion. IEEE Transactions on Software Engineering, 31(12):1074–1092, December 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1566608.

Forax:2004:RIJ

[FDR04] R. Forax, E. Duris, and G. Roussel. A reflective implementation of Java multimethods. IEEE Transactions on Software Engineering, 30(12):1055–1071, December 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377197.

Fetzer:2004:ADM

[FFH04] C. Fetzer, P. Felber, and K. Hogstedt. Automatic detection and masking of nonatomic exception handling. IEEE Transactions on Software Engineering, 30(8):547– 560, August 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1316871.

Flanagan:2005:EPA

[FFQ05] C. Flanagan, S. N. Freund, and S. Qadeer. Exploiting purity for atomicity. IEEE Transactions on Software Engineering, 31(4):275–291, April 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1435350.

Frolund:2002:TEE

[FG02] S. Frolund and R. Guerraoui. e-Transactions: end-to-end reliability for three-tier architectures. IEEE Transactions on Software Engineering, 28 (4):378–395, April 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=995430.

Favre:2009:GEI

[FGLW09] J. M. Favre, D. Gasević, R. Lammel, and A. Winter. Guest Editors’ introduction to the special section on software language engineering. IEEE Transactions on Software Engineering, 35(6):737–741, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=5353438.

Frakes:2005:SRR

[FK05] W. B. Frakes and Kyo Kang.

Software reuse research: status and future. IEEE Transactions on Software Engineering, 31(7):529–536, July 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1492369.

France:2004:UBP

[FKGS04] R. B. France, D. K. Kim, Sudipto Ghosh, and E. Song. A UML-based pattern specification technique. IEEE Transactions on Software Engineering, 30(3):193–206, March 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1271174.

Franceschinis:2006:GEI

[FKW06] G. Franceschinis, J. P. Katoen, and M. Woodside. Guest Editors’ introduction to the special section on the First International Conference on the Quantitative Evaluation of SysTems (QEST). IEEE Transactions on Software Engineering, 32(8): 529–530, August 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1703384.

Furia:2006:CIL

[FMPR06] C. A. Furia, A. Morzenti, M. Pradella, and M. G. Rossi. Comments on “An Interval Logic for Real-Time System Specification”. IEEE Transactions on Software Engineering, 32(6):424–427, June 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1650216.  See

[MN01] and reply [BNR06].

Fu:2005:RTJ

[FMRW05] C. Fu, A. Milanova, B. G. Ryder, and D. G. Wonnacott. Robustness testing of Java server applications. IEEE Transactions on Software Engineering, 31(4):292–311, April 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1435351.

Fioravanti:2001:EPM

[FN01] F. Fioravanti and P. Nesi. Estimation and prediction metrics for adaptive maintenance effort of object-oriented systems. IEEE Transactions on Software Engineering, 27(12):1062– 1084, December 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=988708.

Fenton:2000:QAF

[FO00] N. E. Fenton and N. Ohlsson. Quantitative analysis of faults and failures in a complex software system. IEEE Transactions on Software Engineering, 26(8):797– 814, August 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=879815.

Frey:2002:FSV

[FRC^{+}02] P. Frey, R. Radhakrishnan, H. W. Carter, P. A. Wilsey, and P. Alexander. A formal specification and verification framework for time warp-based parallel simulation. IEEE Transactions on Software Engineering, 28(1): 58–78, January 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=979989.

Foss:2003:SSM

[FSKM03] T. Foss, E. Stensrud, B. Kitchenham, and I. Myrtveit. A simulation study of the model evaluation criterion MMRE. IEEE Transactions on Software Engineering, 29(11):985– 995, November 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1245300.

Fluri:2007:CDT

[FWPG07] B. Fluri, M. Wuersch, M. Pinzger, and H. Gall. Change distilling: Tree differencing for finegrained source code change extraction. IEEE Transactions on Software Engineering, 33(11):

725–743, November 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4339230.

Gueheneuc:2008:DMA

[GA08] Y. G. Guéhéneuc and G. Antoniol. DeMIMA: A multilayered approach for design pattern identification. IEEE Transactions on Software Engineering, 34(5):667–684, September/October 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4564471.

Gurfinkel:2003:TLQ

[GCD03] A. Gurfinkel, M. Chechik, and B. Devereux. Temporal logic query checking: a tool for model exploration. IEEE Transactions on Software Engineering, 29(10):898– 914, October 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1237171.

Gyimothy:2005:EVO

[GFS05] T. Gyimothy, R. Ferenc, and I. Siket. Empirical validation of object-oriented metrics on open source software for fault prediction. IEEE Transactions on Software Engineering, 31(10): 897–910, October 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1542070.

Glasser:2004:ACM

[GGV04] U. Glasser, Y. Gurevich, and M. Veanes. Abstract communication model for distributed systems. IEEE Transactions on Software Engineering, 30 (7):458–472, July 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1318607.

Gribaudo:2002:FSP

[GH02] M. Gribaudo and A. Horvath. Fluid stochastic Petri nets augmented with flush-out arcs: a transient analysis technique. IEEE Transactions on Software Engineering, 28(10): 944–955, October 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1041051.

Gallagher:2008:SAV

[GHM08] K. Gallagher, A. Hatch, and M. Munro. Software architecture visualization: An evaluation framework and its application. IEEE Transactions on Software Engineering, 34(2):260– 270, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4378398.

Gilmore:2001:EAA

[GHR01] S. Gilmore, J. Hillston, and M. Ribaudo. An efficient algorithm for aggregating PEPA models. IEEE Transactions on Software Engineering, 27 (5):449–464, May 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=922715.

Gopal:2002:MPS

[GKMG02] A. Gopal, M. S. Krishnan, T. Mukhopadhyay, and D. R. Goldenson. Measurement programs in software development: determinants of success. IEEE Transactions on Software Engineering, 28(9):863– 875, September 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1033226.

Graves:2000:PFI

[GKMS00] T. L. Graves, A. F. Karr, J. S. Marron, and H. Siy. Predicting fault incidence using software change history. IEEE Transactions on Software Engineering, 26(7):653–661, July 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=859533.

Grammatikakis:2000:PQS

[GL00a] M. D. Grammatikakis and S. Liesche. Priority queues and sorting methods for parallel simulation. IEEE Transactions on Software Engineering, 26(5):401–422, May 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=846298.

Gu:2000:EIA

[GL00b] Junjie Gu and Zhiyuan Li.

Efficient interprocedural array data-flow analysis for automatic program parallelization. IEEE Transactions on Software Engineering, 26(3): 244–261, March 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=842950.

Gokhale:2005:SAS

[GL05] S. S. Gokhale and Michael Rung-Tsong Lyu. A simulation approach to structurebased software reliability analysis. IEEE Transactions on Software Engineering, 31(8): 643–656, August 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1498770.

Gopal:2005:IIF

[GMK05] A. Gopal, T. Mukhopadhyay, and M. S. Krishnan. The impact of institutional forces on software metrics programs. IEEE Transactions on Software Engineering, 31(8):679– 694, August 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1498772.

Gold:2005:SCM

[GML05] N. E. Gold, A. M. Mohan, and P. J. Layzell. Spatial complexity metrics: an investigation of utility. IEEE Transactions on Software Engineering, 31 (3):203–212, March 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1423992.

Griswold:2006:GEI

[GN06] W. G. Griswold and B. Nuseibeh. Guest Editors’ introduction to the special section on the international conference on software engineering. IEEE Transactions on Software Engineering, 32(12): 929–930, December 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4016570.

Gorlatch:2000:TFB

[Gor00] S. Gorlatch. Toward formallybased design of message passing programs. IEEE Transactions on Software Engineering, 26(3): 276–288, March 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=842952.

Gouda:2002:MS

[Gou02] M. G. Gouda. Multiphase stabilization. IEEE Transactions on Software Engineering, 28(2): 201–208, February 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=988499.

Goseva-Popstojanova:2003:ALR

[GPHG^{+}03] K. Goseva-Popstojanova, A. Hassan, A. Guedem, W. Abdelmoez, D. E. M. Nassar, H. Ammar, and A. Mili. Architectural-level risk analysis using UML. IEEE Transactions on Software Engineering, 29(10):946–960, October 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1237174.

Gursaran:2001:AWP

[GR01] Gursaran and G. Roy. On the applicability of Weyuker Property 9 to object-oriented structural inheritance complexity metrics. IEEE Transactions on Software Engineering, 27(4):381–384, April 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=917526. See [ZX02].

Gyimothy:2006:GEI

[GR06] T. Gyimothy and V. Rajlich. Guest Editors’ introduction to the special issue on the international conference on software maintenance and evolution. IEEE Transactions on Software Engineering, 32(9): 625–626, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1707663.

Griswold:2003:GEI

[Gri03] W. G. Griswold. Guest Editor’s introduction: 2002 Conference on the Foundations of Software Engineering. IEEE Transactions on Software Engineering, 29(10):865, October 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1237168.

Guerraoui:2001:GCS

[GS01] R. Guerraoui and A. Schiper. The generic consensus service. IEEE Transactions on Software Engineering, 27(1): 29–41, January 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=895986.

Gregoriades:2005:SBA

[GS05] A. Gregoriades and A. Sutcliffe. Scenario-based assessment of nonfunctional requirements. IEEE Transactions on Software Engineering, 31 (5):392–409, May 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1438375.

Godfrey:2005:UOA

[GZ05] M. W. Godfrey and L. Zou.

Using origin analysis to detect merging and splitting of source code entities. IEEE Transactions on Software Engineering, 31(2):166–181, February 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1401931.

Hagen:2000:EHW

[HA00] C. Hagen and G. Alonso.

Exception handling in workflow management systems. IEEE Transactions on Software Engineering, 26(10):943– 958, October 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=879818.

Heitmeyer:2008:AFM

[HALM08] C. Heitmeyer, M. Archer, E. Leonard, and J. McLean. Applying formal methods to a certifiably secure software system. IEEE Transactions on Software Engineering, 34(1):82–98, January/ February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4407731.

Hatton:2009:PLD

[Hat09] L. Hatton. Power-law distributions of component size in general software systems. IEEE Transactions on Software Engineering, 35(4):566– 572, July/August 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4711063.

Harrold:2002:GEI

[HB02] M. J. Harrold and A. Bertolino. Guest Editors’ introduction: 2000 international symposium on software testing and analysis. IEEE Transactions on Software Engineering, 28(2): 113–114, February 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=988493.

Holzmann:2007:DME

[HB07] G. J. Holzmann and D. Bosnacki. The design of a multicore extension of the SPIN model checker. IEEE Transactions on Software Engineering, 33(10):

659–674, October 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4302778.

Hughes:2008:IGM

[HB08] G. Hughes and T. Bultan. Interface grammars for modular software model checking. IEEE Transactions on Software Engineering, 34(5):614–632, September/October 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4609388.

Hsu:2003:SPN

[HCC03] Ping-Yu Hsu, Yuan-Bin Chang, and Yen-Liang Chen. STRPN: a Petri-net approach for modeling spatial-temporal relations between moving multimedia objects. IEEE Transactions on Software Engineering, 29 (1):63–76, January 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1166589.

Hwang:2004:CTT

[HCC04a] Gwan-Hwan Hwang, Sheng-Jen Chang, and Huey-Der Chu. Correction to “Technology for testing nondeterministic client/server database applications”. IEEE Transactions on Software Engineering, 30(4):278, April 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1274046.  See

[HCC04b].

Hwang:2004:TTN

[HCC04b] Gwan-Hwan Hwang, Sheng-Jen Chang, and Huey-Der Chu. Technology for testing nondeterministic client/ server database applications. IEEE Transactions on Software Engineering, 30(1):59– 77, January 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1265736. See com-

ments [HCC04a].

Haider:2008:EDB

[HCC^{+}08] S. W. Haider, J. W. Cangussu, K. M. L. Cooper, R. Dantu, and S. Haider. Estimation of defects based on defect decay model: ED^{3}M. IEEE Transactions on Software Engineering, 34(3):336– 356, May/June 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4492790.

Hendrix:2002:CEC

[HCM02a] D. Hendrix, J. H. Cross, and S. Maghsoodloo. Corrections to “The effectiveness of control structure diagrams in source code comprehension activities”. IEEE Transactions on Software Engineering, 28(6):624, June 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1010064.  See

[HCM02b].

Hendrix:2002:ECS

[HCM02b] D. Hendrix, J. H. Cross, and S. Maghsoodloo. The effectiveness of control structure diagrams in source code comprehension activities. IEEE Transactions on Software Engineering, 28(5):463–477, May 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1000450.

See corrections [HCM02a].

Hayes:2006:ACL

[HDS06] J. H. Hayes, A. Dekhtyar, and S. K. Sundaram. Advancing candidate link generation for requirements tracing: the study of methods. IEEE Transactions on Software Engineering, 32(1):4–19, January 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1583599.

Hearty:2009:PPV

[HFMN09] P. Hearty, N. Fenton, D. Marquez, and M. Neil. Predicting project velocity in XP using a learning dynamic Bayesian network model. IEEE Transactions on Software Engineering, 35(1):124–137, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4624275.

Hamill:2009:CTS

[HGP09] M. Hamill and K. Goseva-Popstojanova. Common trends in software fault and failure data. IEEE Transactions on Software Engineering, 35(4):484–496, July/ August 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4760152.

Hou:2006:USS

[HH06] D. Hou and H. J. Hoover. Using SCL to specify and check design intent in source code. IEEE Transactions on Software Engineering, 32(6): 404–423, June 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1650215.

Hungerford:2004:RSD

[HHC04] B. C. Hungerford, A. R. Hevner, and R. W. Collins. Reviewing software diagrams: a cognitive study. IEEE Transactions on Software Engineering, 30(2):82–96, February 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1265814.

Harman:2004:TT

[HHH^{+}04] M. Harman, L. Hu, R. Hierons, J. Wegener, H. Sthamer, A. Baresel, and M. Roper. Testability transformation. IEEE Transactions on Software Engineering, 30(1):3–16, January 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1265732.

Hannay:2008:RDA

[HJ08] J. Hannay and M. Jørgensen.

The role of deliberate artificial design elements in software engineering experiments. IEEE Transactions on Software Engineering, 34(2):242– 259, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4453833.

Harel:2002:OSB

[HK02] D. Harel and O. Kupferman. On object systems and behavioral inheritance. IEEE Transactions on Software Engineering, 28(9):889– 903, September 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1033228. See com-

ments [CJ03].

Han:2009:CGP

[HKB09] T. Han, J. P. Katoen, and D. Berteun. Counterexample generation in probabilistic model checking. IEEE Transactions on Software Engineering, 35(2):241–257, March/April 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4770111.

Harman:2005:GES

[HKL05] M. Harman, B. Korel, and P. K. Linos. Guest editorial: Special issue on software maintenance and evolution. IEEE Transactions on Software Engineering, 31(10): 801–803, October 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1542063.

Haran:2007:TCE

[HKL^{+}07] M. Haran, A. Karr, M. Last, A. Orso, A. A. Porter, A. Sanil, and S. Fouche. Techniques for classifying executions of deployed software to support software engineering tasks. IEEE Transactions on Software Engineering, 33(5):287– 304, May 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4160968.

Hillston:2009:GEI

[HKT09] J. Hillston, M. Kwiatkowska, and M. Telek. Guest Editors’ introduction to the special issue on quantitative evaluation of computer systems. IEEE Transactions on Software Engineering, 35(2):145– 147, March/April 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4809711.

Huang:2003:MMC

[HL03] Sun-Jen Huang and R. Lai.

Measuring the maintainability of a communication protocol based on its formal specification. IEEE Transactions on Software Engineering, 29 (4):327–344, April 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1191797.

Huang:2003:USS

[HLK03] Chin-Yu Huang, M. R. Lyu, and Sy-Yen Kuo. A unified scheme of some nonhomogeneous Poisson process models for software reliability estimation. IEEE Transactions on Software Engineering, 29(3): 261–269, March 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1183936.

Haley:2008:SRE

[HLMN08] C. Haley, R. Laney, J. Moffett, and B. Nuseibeh. Security requirements engineering: A framework for representation and analysis. IEEE Transactions on Software Engineering, 34(1):133–153, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4359475.

Havelund:2001:FAS

[HLP01] K. Havelund, M. Lowry, and J. Penix. Formal analysis of a space-craft controller using SPIN. IEEE Transactions on Software Engineering, 27(8): 749–765, August 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=940728.

Hsiung:2004:VAF

[HLT^{+}04] Pao-Ann Hsiung, Shang-Wei Lin, Chih-Hao Tseng, Trong-Yen Lee, Jin-Ming Fu, and Win-Bin See. VERTAF: an application framework for the design and verification of embedded real-time software. IEEE Transactions on Software Engineering, 30(10):656– 674, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339277.

Herbsleb:2003:ESS

[HM03] J. D. Herbsleb and A. Mockus.

An empirical study of speed and communication in globally distributed software development. IEEE Transactions on Software Engineering, 29 (6):481–494, June 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1205177.

Hassan:2005:GEI

[HMHJ05] A. E. Hassan, A. Mockus, R. C.

Holt, and P. M. Johnson. Guest Editor’s introduction: Special issue on mining software repositories. IEEE Transactions on Software Engineering, 31 (6):426–428, June 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1463227.

Holzmann:1997:MCS

[Hol97] G. J. Holzmann. The model checker SPIN. IEEE Transactions on Software Engineering, 23(5):279–295, May 1997. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?

                                        arnumber=588521. See com-

ments [BCY01].

Halfond:2008:WPW

[HOM08] W. Halfond, A. Orso, and P. Manolios. WASP: Protecting Web applications using positive tainting and syntaxaware evaluation. IEEE Transactions on Software Engineering, 34(1):65–81, January/ February 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4359474.

Haxthausen:2000:FDV

[HP00] A. E. Haxthausen and J. Peleska. Formal development and verification of a distributed railway control system. IEEE Transactions on Software Engineering, 26(8): 687–701, August 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=879808.

Henkel:2007:DDJ

[HRD07] J. Henkel, C. Reichenbach, and A. Diwan. Discovering documentation for Java container classes. IEEE Transactions on Software Engineering, 33(8): 526–543, August 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?

arnumber=4267024. See errata

[HRD08].

Henkel:2008:EDD

[HRD08] J. Henkel, C. Reichenbach, and A. Diwan. Errata for “Discovering Documentation for Java Container Classes” [aug 07 526–543]. IEEE Transactions on Software Engineering, 34(2): 303, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4476755.  See

[HRD07].

Hall:2008:POS

[HRJ08] J. Hall, L. Rapanotti, and M. Jackson. Problem oriented software engineering: Solving the package router control problem. IEEE Transactions on Software Engineering, 34(2):226–241, March/April 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4384506.

Harrold:2001:ESP

[HRRW01] M. J. Harrold, D. Rosenblum, G. Rothermel, and E. Weyuker. Empirical studies of a prediction model for regression test selection. IEEE Transactions on Software Engineering, 27 (3):248–263, March 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=910860.

Hastings:2001:VBA

[HS01] T. E. Hastings and A. S. M. Sajeev. A vector-based approach to software size measurement and effort estimation. IEEE Transactions on Software Engineering, 27(4): 337–350, April 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=917523.

Holzmann:2002:AVM

[HS02] G. J. Holzmann and M. H. Smith. An automated verification method for distributed systems software based on model extraction. IEEE Transactions on Software Engineering, 28 (4):364–377, April 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=995426.

Harrold:2003:GEI

[HS03] M. J. Harrold and W. Schafer. Guest Editors’ introduction: 2001 international conference on software engineering. IEEE Transactions on Software Engineering, 29(2):97– 98, February 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1178047.

Hannay:2007:SRT

[HSD07] J. E. Hannay, Dag I. K.

Sjøberg, and T. Dyba. A systematic review of theory use in software engineering experiments. IEEE Transactions on Software Engineering, 33(2): 87–107, February 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4052585.

Horvath:2002:TDA

[HT02] A. Horvath and M. Telek.

Time domain analysis of non-Markovian stochastic Petri nets with PRI transitions. IEEE Transactions on Software Engineering, 28(10):933– 943, October 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1041050.

Halle:2009:SVD

[HVC09] S. Halle, R. Villemaire, and O. Cherkaoui. Specifying and validating data-aware temporal Web service properties. IEEE Transactions on Software Engineering, 35(5):669–683, September/October 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4907003.

Healy:2002:ADE

[HW02a] C. A. Healy and D. B. Whalley. Automatic detection and exploitation of branch constraints for timing analysis. IEEE Transactions on Software Engineering, 28(8):763– 781, August 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1027799.

Hrischuk:2002:LCR

[HW02b] C. E. Hrischuk and C. M. Woodside. Logical clock requirements for reverse engineering scenarios from a distributed system. IEEE Transactions on Software Engineering, 28 (4):321–339, April 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=995416.

Holmes:2006:ASC

[HWM06] R. Holmes, R. J. Walker, and G. C. Murphy. Approximate structural context matching: An approach to recommend relevant examples. IEEE Transactions on Software Engineering, 32(12):952– 970, December 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4016572.

Iyer:2001:ATE

[IR01] S. Iyer and S. Ramesh. Apportioning: a technique for efficient reachability analysis of concurrent objectoriented programs. IEEE Transactions on Software Engineering, 27(11):1037–1056, November 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=965343.

Inoue:2005:RSS

[IYY^{+}05] K. Inoue, R. Yokomori, T. Yamamoto, M. Matsushita, and S. Kusumoto. Ranking significance of software components based on use relations. IEEE Transactions on Software Engineering, 31(3):213– 225, March 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1423993.

Jorgensen:2004:ESS

[JC04] M. Jorgensen and G. J. Carelius. An empirical study of software project bidding. IEEE Transactions on Software Engineering, 30(12):953– 969, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377191.

Jilani:2001:DAM

[JDM01] L. L. Jilani, J. Desharnais, and A. Mili. Defining and applying measures of distance between specifications. IEEE Transactions on Software Engineering, 27(8):673– 703, August 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=940565.

Jeffrey:2007:IFD

[JG07] D. Jeffrey and N. Gupta. Improving fault detection capability by selectively retaining test cases during test suite reduction. IEEE Transactions on Software Engineering, 33(2): 108–123, February 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4052586.

Jorgensen:2009:ILL

[JG09] M. Jørgensen and T. M. Gruschke. The impact of lessonslearned sessions on effort estimation and uncertainty assessments. IEEE Transactions on Software Engineering, 35(3): 368–383, May/June 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4752843.

Jones:2003:TSR

[JH03] J. A. Jones and M. J. Harrold. Test-suite reduction and prioritization for modified condition/decision coverage. IEEE Transactions on Software Engineering, 29(3): 195–209, March 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1183927.

Jiau:2004:CSP

[JKS04] H. C. Jiau, Chia Hung Kao, and Kuo-Feng Ssu. Comments on “Software process representation and analysis for framework instantiation”. IEEE Transactions on Software Engineering, 30(10):707– ??, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339280.  See

[OFdL^{+}04].

Jorgensen:2004:RSE

[JMO04] M. Jorgensen and K. Molokken-Ostvold. Reasons for software effort estimation error: impact of respondent role, information collection approach, and data analysis method. IEEE Transactions on Software Engineering, 30(12):993– 1007, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1377193.

Juristo:2007:GEU

[JMSS07] N. Juristo, A. Moreno, and M. I. Sanchez-Segura. Guidelines for eliciting usability functionalities. IEEE Transactions on Software Engineering, 33(11):744–758, November 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4339231.

Jorgensen:2004:RAE

[Jor04] M. Jorgensen. Realism in assessment of effort estimation uncertainty: it matters how you ask. IEEE Transactions on Software Engineering, 30 (4):209–217, April 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1274041.

Jorgensen:2005:EBG

[Jor05] M. Jorgensen. Evidence-based guidelines for assessment of software development cost uncertainty. IEEE Transactions on Software Engineering, 31(11):942–954, November 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1556553.

Julien:2006:EFR

[JR06] C. Julien and G. C. Roman. EgoSpaces: facilitating rapid development of context-aware mobile applications. IEEE Transactions on Software Engineering, 32(5): 281–298, May 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1642677.

Janicki:2001:FTA

[JS01] R. Janicki and E. Sekerinski. Foundations of the trace assertion method of module interface specification. IEEE Transactions on Software Engineering, 27(7):577– 598, July 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=935852.

Jalote:2002:OCL

[JS02] P. Jalote and A. Saxena. Optimum control limits for employing statistical process control in software process. IEEE Transactions on Software Engineering, 28(12):1126–1134, December 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1158286.

Jorgensen:2007:SRS

[JS07] M. Jorgensen and M. Shepperd. A systematic review of software development cost estimation studies. IEEE Transactions on Software Engineering, 33 (1):33–53, January 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4027147.

Juan:2001:RMR

[JTMZ01] E. Y. T. Juan, J. J. P. Tsai, T. Murata, and Yi Zhou. Reduction methods for real-time systems using delay time Petri nets. IEEE Transactions on Software Engineering, 27(5): 422–448, May 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=922714.

Jackson:2001:LEO

[JW01] D. Jackson and A. Waingold. Lightweight extraction of object models from bytecode. IEEE Transactions on Software Engineering, 27(2):156– 169, February 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=908960.

Janka:2002:VBM

[JWB02] R. S. Janka, L. M. Wills, and L. B. Baumstark. Virtual benchmarking and model continuity in prototyping embedded multiprocessor signal processing systems. IEEE Transactions on Software Engineering, 28(9):832–846, September 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1033224.

Jiau:2006:CAA

[JY06] Hewijin Christine Jiau and Dung-Feng Yu. Comments on “Automatic analysis of consistency between requirements and designs”. IEEE Transactions on Software Engineering, 32(4):279–280, April 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1628973.  See

[CG01].

Kao:2003:CQP

[Kao03] Chia Hung Kao. Comments on “Quality, productivity, and learning in framework-based development: an exploratory case study”. IEEE Transactions on Software Engineering, 29(3):288, March 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1183941.  See

[MRS02].

Kim:2004:SRA

[KBYC04] Sung Kim, F. B. Bastani, I-Ling Yen, and I. R. Chen. Systematic reliability analysis of a class of application-specific embedded software framework. IEEE Transactions on Software Engineering, 30(4):218– 230, April 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1274042.

Kang:2004:SMT

[KC04] J. A. Kang and A. M. K. Cheng. Shortening matching time in OPS5 production systems. IEEE Transactions on Software Engineering, 30 (7):448–457, July 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1318606.

Konrad:2004:OAP

[KCC04] S. Konrad, B. H. C. Cheng, and L. A. Campbell. Object analysis patterns for embedded systems. IEEE Transactions on Software Engineering, 30(12): 970–992, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1377192.

Kuo:2003:EOS

[KCLL03] Tei-Wei Kuo, Li-Pin Chang, Yu-Hua Liu, and Kwei-Jay Lin. Efficient online schedulability tests for real-time systems. IEEE Transactions on Software Engineering, 29(8): 734–751, August 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1223647.

Koppol:2002:IIT

[KCT02] P. V. Koppol, R. H. Carver, and Kuo-Chung Tai. Incremental integration testing of concurrent programs. IEEE Transactions on Software Engineering, 28(6):607– 623, June 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1010062.

Koopman:2000:EHE

[KD00] P. Koopman and J. DeVale.

The exception handling effectiveness of POSIX operating systems. IEEE Transactions on Software Engineering, 26(9): 837–848, September 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=877845.

Kraft:2009:GRP

[KDM09] N. A. Kraft, E. B. Duffy, and B. A. Malloy. Grammar recovery from parse trees and metrics-guided grammar refactoring. IEEE Transactions on Software Engineering, 35(6):780–794, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=5278661.

Kelly:2006:SDC

[Kel06] D. Kelly. A study of design characteristics in evolving software using stability as a criterion. IEEE Transactions on Software Engineering, 32 (5):315–329, May 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1642679.

Kramer:2001:GEI

[KGR01] J. Kramer, D. Garlan, and D. S. Rosenblum. Guest Editors’ introduction: 1999 international conference on software engineering. IEEE Transactions on Software Engineering, 27(2): 97–98, February 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=908956.

King:2000:PMC

[KHCP00] S. King, J. Hammond, R. Chapman, and A. Pryor. Is proof more cost-effective than testing? IEEE Transactions on Software Engineering, 26(8): 675–686, August 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=879807.

Kitchenham:2001:MSM

[KHL01] B. A. Kitchenham, R. T.

Hughes, and S. G. Linkman. Modeling software measurement data. IEEE Transactions on Software Engineering, 27(9): 788–804, September 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=950316.

Khoumsi:2002:TAT

[Kho02]  A. Khoumsi.  A tempo-

ral approach for testing distributed systems. IEEE Transactions on Software Engineering, 28(11):1085–1103, November 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1049406.

Kamiya:2002:CMT

[KKI02] T. Kamiya, S. Kusumoto, and K. Inoue. CCFinder: a multilinguistic token-based code clone detection system for large scale source code. IEEE Transactions on Software Engineering, 28(7):654– 670, July 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1019480.

Keung:2008:AXP

[KKJ08] J. W. Keung, B. A. Kitchenham, and D. R. Jeffery. Analogy-X: Providing statistical inference to analogybased software cost estimation. IEEE Transactions on Software Engineering, 34(4):471– 484, July/August 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4527255.

Kordon:2002:IRS

[KL02] F. Kordon and Luqi. An introduction to rapid system prototyping. IEEE Transactions on Software Engineering, 28(9): 817–821, September 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1033222.

Kang:2000:ESS

[KLK00] I. Kang, I. Lee, and Young-Si Kim. An efficient state space generation for the analysis of real-time systems. IEEE Transactions on Software Engineering, 26(5):453– 477, May 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=846302.

Kuncak:2006:MPA

[KLZR06] V. Kuncak, P. Lam, K. Zee, and M. C. Rinard. Modular pluggable analyses for data structure consistency. IEEE Transactions on Software Engineering, 32(12):988– 1005, December 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4016574.

Kitchenham:2004:SPM

[KM04] B. Kitchenham and E. Mendes.

Software productivity measurement using multiple size measures. IEEE Transactions on Software Engineering, 30(12):1023–1035, December 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377195.

Ko:2006:ESH

[KMCA06] A. J. Ko, B. A. Myers, M. J.

Coblenz, and H. H. Aung. An exploratory study of how developers seek, relate, and collect relevant information during software maintenance tasks. IEEE Transactions on Software Engineering, 32(12):971– 987, December 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4016573.

Kirkegaard:2004:SAX

[KMS04] C. Kirkegaard, A. Moller, and M. I. Schwartzbach. Static analysis of XML transformations in Java. IEEE Transactions on Software Engineering, 30(3):181–192, March 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1271173.

Kistijantoro:2008:EAS

[KMSL08] A. I. Kistijantoro, G. Morgan, S. K. Shrivastava, and M. C. Little. Enhancing an application server to support available components. IEEE Transactions on Software Engineering, 34(4):531– 545, July/August 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4528966.

Kitchenham:2007:CVW

[KMT07] B. A. Kitchenham, E. Mendes, and G. H. Travassos. Cross versus within-company cost estimation studies: A systematic review. IEEE Transactions on Software Engineering, 33(5):316–329, May 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4160970.

Knight:2002:E

[Kni02] J. Knight. Editorial. IEEE Transactions on Software Engineering, 28(1):3, January 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=979985.

Knight:2003:ENA

[Kni03] J. Knight. Editorial: New AEIC introduction and AE farewell. IEEE Transactions on Software Engineering, 29 (11):961, November 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1245297.

Knight:2004:EAF

[Kni04a] J. Knight. Editorial: AE farewell. IEEE Transactions on Software Engineering, 30 (5):281, May 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1291831.

Knight:2004:EAIb

[Kni04b] J. Knight. Editorial: AE introduction. IEEE Transactions on Software Engineering, 30(8): 489–490, August 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1316866.

Knight:2004:EAIa

[Kni04c] J. Knight. Editorial: AE introduction and farewell. IEEE Transactions on Software Engineering, 30(10):633, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1339275.

Knight:2004:EST

[Kni04d] J. Knight. Editorial: the state of TSE. IEEE Transactions on Software Engineering, 30 (2):81, February 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1265813.

Knight:2005:ENAa

[Kni05a] J. Knight. Editorial: New Associate Editor introduction. IEEE Transactions on Software Engineering, 31(6): 425, June 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1463226.

Knight:2005:ENAb

[Kni05b] J. Knight. Editorial: New associate editors introduction. IEEE Transactions on Software Engineering, 31(3):185– 186, March 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1423990.

Knight:2006:ENE

[Kni06] J. Knight. Editorial: A new Editor-in-chief and the state of TSE. IEEE Transactions on Software Engineering, 32 (1):1, January 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1583597.

Kounev:2006:PME

[Kou06] S. Kounev. Performance modeling and evaluation of distributed component-based systems using queueing Petri nets. IEEE Transactions on Software Engineering, 32(7):486– 502, July 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1677534.

Kemerer:2009:IDC

[KP09] C. F. Kemerer and M. C.

Paulk. The impact of design and code reviews on software quality: An empirical study based on PSP data. IEEE Transactions on Software Engineering, 35(4):534– 550, July/August 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4815279.

Kitchenham:2003:MSB

[KPLJ03] B. Kitchenham, L. M. Pickard, S. Linkman, and P. W. Jones. Modeling software bidding risks. IEEE Transactions on Software Engineering, 29(6): 542–554, June 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1205181.

Kitchenham:2002:PGE

[KPP^{+}02] B. A. Kitchenham, S. L. Pfleeger, L. M. Pickard, P. W. Jones, D. C. Hoaglin, K. El Emam, and J. Rosenberg. Preliminary guidelines for empirical research in software engineering. IEEE Transactions on Software Engineering, 28(8): 721–734, August 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1027796.

Kramer:2006:EMN

[Kra06a] J. Kramer. Editorial: A message from the new Editorin-Chief. IEEE Transactions on Software Engineering, 32 (1):2–3, January 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1583598.

Kramer:2006:ENAb

[Kra06b] J. Kramer. Editorial: New associate editors introduction. IEEE Transactions on Software Engineering, 32(3):137– 139, March 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1610606.

Kramer:2006:ENAa

[Kra06c] J. Kramer. Editorial: New associate editors introduction. IEEE Transactions on Software Engineering, 32(10):769– 770, October 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1717470.

Kramer:2007:ESJ

[Kra07] J. Kramer. Editorial: State of the journal. IEEE Transactions on Software Engineering, 33(1):1, January 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4027144.

Kramer:2008:ENAa

[Kra08a] J. Kramer. Editorial: New Associate Editor introduction. IEEE Transactions on Software Engineering, 34(2):161, March/April 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4476754.

Kramer:2008:ENAb

[Kra08b] J. Kramer. Editorial: New associate editors introduction. IEEE Transactions on Software Engineering, 34(6): 721–722, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4709490.

Kramer:2008:SJA

[Kra08c] J. Kramer. State of the journal address. IEEE Transactions on Software Engineering, 34(1):1– 2, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4444342.

Kramer:2009:ENA

[Kra09a] J. Kramer. Editorial: New associate editors introduction. IEEE Transactions on Software Engineering, 35(4):449, July/August 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=5186361.

Kramer:2009:ESJ

[Kra09b] J. Kramer. Editorial: State of the journal address. IEEE Transactions on Software Engineering, 35(1):1, January/ February 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4771848.

Krishnamurthy:2006:SWG

[KRM06] D. Krishnamurthy, J. A. Rolia, and S. Majumdar. A synthetic workload generation technique for stress testing session-based systems. IEEE Transactions on Software Engineering, 32(11): 868–882, November 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4015510.

Kadamuddi:2000:CAP

[KT00] D. Kadamuddi and J. J. P.

Tsai. Clustering algorithm for parallelizing software systems in multiprocessors environment. IEEE Transactions on Software Engineering, 26 (4):340–361, April 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=844493.

Kallepalli:2001:MMU

[KT01] C. Kallepalli and J. Tian.

Measuring and modeling usage and reliability for statistical Web testing. IEEE Transactions on Software Engineering, 27(11):1023–1036, November 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=965342.

Koru:2005:CHC

[KT05] A. G. Koru and J. Tian. Comparing high-change modules and modules with the highest measurement values in two large-scale open-source products. IEEE Transactions on Software Engineering, 31(8): 625–642, August 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1498769.

Kemper:2009:ATA

[KT09] P. Kemper and C. Tepper. Automated trace analysis of discrete-event system models. IEEE Transactions on Software Engineering, 35(2):195– 208, March/April 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4620122.

Kshemkalyani:2007:DAS

[KW07] A. Kshemkalyani and B. Wu. Detecting arbitrary stable properties using efficient snapshots. IEEE Transactions on Software Engineering, 33(5): 330–346, May 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4160971.

Kuhn:2004:SFI

[KWG04] D. R. Kuhn, D. R. Wallace, and A. M. Gallo. Software fault interactions and implications for software testing. IEEE Transactions on Software Engineering, 30(6):418– 421, June 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1321063.

Kim:2008:CSC

[KWZ08] S. Kim, E. J. Whitehead, Jr., and Y. Zhang. Classifying software changes: Clean or buggy? IEEE Transactions on Software Engineering, 34(2):181– 196, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4408585.

Koru:2009:IFF

[KZEL09] A. G. Koru, D. Zhang, K. El Emam, and H. Liu. An investigation into the functional form of the size-defect relationship for software modules. IEEE Transactions on Software Engineering, 35(2):293– 304, March/April 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4693715.

LeTraon:2006:DCI

[LBJ06] Y. Le Traon, B. Baudry, and J. M. Jezequel. Design by contract to improve software vigilance. IEEE Transactions on Software Engineering, 32(8): 571–586, August 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1703388.

Lessmann:2008:BCM

[LBMP08] S. Lessmann, B. Baesens, C. Mues, and S. Pietsch. Benchmarking classification models for software defect prediction: A proposed framework and novel findings. IEEE Transactions on Software Engineering, 34(4):485– 496, July/August 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4527256.

Lee:2004:ORT

[LC04] Y. H. Lee and A. M. K. Cheng. Optimizing real-time equational rule-based systems. IEEE Transactions on Software Engineering, 30(2):112– 125, February 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1265816.

Lei:2006:RTC

[LC06] Y. Lei and R. H. Carver.

Reachability testing of concurrent programs. IEEE Transactions on Software Engineering, 32(6):382–403, June 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1650214.

Lui:2008:EPP

[LCN08] K. M. Lui, K. C. C. Chan, and J. Nosek. The effect of pairs in program design tasks. IEEE Transactions on Software Engineering, 34(2):197– 211, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4378344.

Li:2008:SPS

[LCS^{+}08] J. Li, R. Conradi, O. P.

Slyngstad, M. Torchiano, M. Morisio, and C. Bunse. A state-of-the-practice survey of risk management in development with off-the-shelf software components. IEEE Transactions on Software Engineering, 34(2):271– 286, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4459339.

Lanza:2003:PVL

[LD03] M. Lanza and S. Ducasse. Polymetric views — a lightweight visual approach to reverse engineering. IEEE Transactions on Software Engineering, 29(9): 782–795, September 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1232284.

Lee:2004:FCD

[Lee04]  Soojung Lee.  Fast, cen-

tralized detection and resolution of distributed deadlocks in the generalized model. IEEE Transactions on Software Engineering, 30(9):561– 573, September 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1324644.

Laitenberger:2001:IRQ

[LEH01] O. Laitenberger, K. El Emam, and T. G. Harbich. An internally replicated quasiexperimental comparison of checklist and perspective based reading of code documents. IEEE Transactions on Software Engineering, 27(5):387– 421, May 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=922713.

Leveson:2000:ISA

[Lev00] N. G. Leveson. Intent specifications: an approach to building human-centered specifications. IEEE Transactions on Software Engineering, 26(1): 15–35, January 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=825764.

Liu:2006:SDH

[LFY^{+}06] Chao Liu, Long Fei, Xifeng Yan, Jiawei Han, and S. P. Midkiff. Statistical debugging: A hypothesis testing-based approach. IEEE Transactions on Software Engineering, 32(10): 831–848, October 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1717474.

Little:2003:PII

[LG03] R. G. Little and M. L. Gibson. Perceived influences on implementing data warehousing. IEEE Transactions on Software Engineering, 29(4): 290–296, April 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1191794.

Liu:2005:DLP

[LGF05] Y. Liu, I. Gorton, and A. Fekete. Design-level performance prediction of component-based applications. IEEE Transactions on Software Engineering, 31(11):928– 941, November 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1556552.

Lai:2003:MES

[LH03] R. Lai and Sun-Jen Huang. A model for estimating the size of a formal communication protocol specification and its implementation. IEEE Transactions on Software Engineering, 29 (1):46–62, January 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1166588.

Li:2007:SAR

[LHH07] Z. Li, M. Harman, and R. M. Hierons. Search algorithms for regression test case prioritization. IEEE Transactions on Software Engineering, 33 (4):225–237, April 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4123325.

Long:2003:TST

[LHS03] B. Long, D. Hoffman, and P. Strooper. Tool support for testing concurrent Java components. IEEE Transactions on Software Engineering, 29 (6):555–566, June 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1205182.

Lee:2000:DRS

[LI00] I. Lee and R. K. Iyer. Diagnosing rediscovered software problems using symptoms. IEEE Transactions on Software Engineering, 26(2): 113–127, February 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=841113.

Littlewood:2000:UPD

[Lit00] B. Littlewood. The use of proof in diversity arguments. IEEE Transactions on Software Engineering, 26(10):1022– 1023, October 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=879822.

Lakhotia:2005:MDO

[LKV05] Arun Lakhotia, Eric Uday Kumar, and M. Venable. A method for detecting obfuscated calls in malicious binaries. IEEE Transactions on Software Engineering, 31(11): 955–968, November 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1556554.

Lotz:2000:FSM

[LKW00] V. Lotz, V. Kessler, and G. H. Walter. A formal security model for microprocessor hardware. IEEE Transactions on Software Engineering, 26(8): 702–712, August 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=879809.

Lum:2003:UCC

[LL03] Wai Yip Lum and F. C. M. Lau. User-centric content negotiation for effective adaptation service in mobile computing. IEEE Transactions on Software Engineering, 29(12):1100– 1111, December 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265524.

Lee:2001:BCR

[LLH^{+}01] Chang-Gun Lee, Kwangpo Lee, Joosun Hahn, Yang-Min Seo, Sang Lyul Min, Rhan Ha, Seongsoo Hong, Chang Yun Park, Minsuk Lee, and Chong Sang Kim. Bounding cache-related preemption delay for real-time systems. IEEE Transactions on Software Engineering, 27(9):805– 826, September 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=950317.

Li:2006:CMF

[LLMZ06] Z. Li, S. Lu, S. Myagmar, and Y. Zhou. CP-Miner: finding copy-paste and related bugs in large-scale software code. IEEE Transactions on Software Engineering, 32(3):176– 192, March 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1610609.

Lutz:2004:EAS

[LM04] R. R. Lutz and I. C. Mikulski.

Empirical analysis of safetycritical anomalies during operations. IEEE Transactions on Software Engineering, 30 (3):172–180, March 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1271171.

Lodi:2007:SDC

[LPRT07] G. Lodi, F. Panzieri, D. Rossi, and E. Turrini. SLAdriven clustering of QoSaware application servers. IEEE Transactions on Software Engineering, 33(3):186– 197, March 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4084136.

Littlewood:2000:MEC

[LPSS00] B. Littlewood, P. T. Popov, L. Strigini, and N. Shryane. Modeling the effects of combining diverse software fault detection techniques. IEEE Transactions on Software Engineering, 26(12):1157–1167, December 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=888629.

Litoiu:2000:DPR

[LRS00] M. Litoiu, J. Rolia, and G. Serazzi. Designing process replication and activation: a quantitative approach. IEEE Transactions on Software Engineering, 26(12):1168– 1178, December 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=888630.

Li:2004:FVA

[LRSF04] P. Li, Binoy Ravindran, S. Suhaib, and S. Feizabadi. A formally verified application-level framework for real-time scheduling on POSIX real-time operating systems. IEEE Transactions on Software Engineering, 30(9):613– 629, September 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1324648.

Lei:2003:ESN

[LS03a] Skylar Lei and M. R. Smith.

Evaluation of several nonparametric bootstrap methods to estimate confidence intervals for software metrics. IEEE Transactions on Software Engineering, 29(11):996– 1004, November 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1245301.

Li:2003:RSE

[LS03b] Ming Li and C. S. Smidts. A ranking of software engineering measures based on expert opinion. IEEE Transactions on Software Engineering, 29(9): 811–824, September 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1232286.

Lau:2007:SCM

[LW07a] K. K. Lau and Z. Wang.

Software component models. IEEE Transactions on Software Engineering, 33(10):709– 724, October 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4302781.

Littlewood:2007:UMA

[LW07b] B. Littlewood and D. Wright. The use of multilegged arguments to increase confidence in safety claims for softwarebased systems: A study based on a BBN analysis of an idealized example. IEEE Transactions on Software Engineering, 33(5):347–365, May 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4160972.

Luqi:2004:DDD

[LZBQ04] Luqi, L. Zhang, V. Berzins, and Y. Qiao. Documentation driven development for complex real-time systems. IEEE Transactions on Software Engineering, 30(12):936– 952, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377190.

Liu:2008:SSF

[LZH08] C. Liu, X. Zhang, and J. Han. A systematic study of failure proximity. IEEE Transactions on Software Engineering, 34(6):826–843, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4589219.

Morel:2004:SAC

[MA04] B. Morel and P. Alexander. SPARTACAS: automating component reuse and adaptation. IEEE Transactions on Software Engineering, 30(9): 587–600, September 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1324646.

Mock:2005:PSD

[MACE05] M. Mock, D. C. Atkinson, C. Chambers, and S. J. Eggers. Program slicing with dynamic points-to sets. IEEE Transactions on Software Engineering, 31(8):657– 678, August 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1498771.

Mendonca:2000:VAI

[MB00a] M. G. Mendonca and V. R.

Basili. Validation of an approach for improving existing measurement frameworks. IEEE Transactions on Software Engineering, 26(6):484– 499, June 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=852739.

Murata:2000:HIH

[MB00b] T. Murata and A. Borgida.

Handling of irregularities in human centered systems: a unified framework for data and processes. IEEE Transactions on Software Engineering, 26(10):959–977, October 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=879819.

Marre:2003:USS

[MB03] M. Marre and A. Bertolino. Using spanning sets for coverage testing. IEEE Transactions on Software Engineering, 29(11): 974–984, November 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1245299.

Maqbool:2007:HCS

[MB07] O. Maqbool and H. Babri. Hierarchical clustering for software architecture recovery. IEEE Transactions on Software Engineering, 33(11):759– 780, November 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4339232.

Mangeruca:2007:SPD

[MBFSV07] L. Mangeruca, M. Baleani, A. Ferrari, and A. Sangiovanni-Vincentelli. Semantics-preserving design of embedded control software from synchronous models. IEEE Transactions on Software Engineering, 33(8): 497–509, August 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4267022.

Masood:2009:SET

[MBGM09] A. Masood, R. Bhatti, A. Ghafoor, and A. P. Mathur. Scalable and effective test generation for role-based access control systems. IEEE Transactions on Software Engineering, 35(5):654–668, September/October 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4967616.

Mookerjee:2002:DCP

[MC02] V. S. Mookerjee and I. R. Chiang. A dynamic coordination policy for software system construction. IEEE Transactions on Software Engineering, 28 (7):684–694, July 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1019482.

Menzies:2006:SBP

[MCHL06] T. Menzies, Z. Chen, J. Hihn, and K. Lum. Selecting best practices for effort estimation. IEEE Transactions on Software Engineering, 32(11):883– 895, November 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4015511.

Mahony:2000:TCO

[MD00] B. Mahony and Jin Song Dong. Timed communicating object z. IEEE Transactions on Software Engineering, 26(2):150– 177, February 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=841115.

Menzies:2003:MSF

[MD03] T. Menzies and J. S. Di Stefano. More success and failure factors in software reuse. IEEE Transactions on Software Engineering, 29(5):474– 477, May 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1199076. See com-

ments [MET03].

Menzies:2007:PPR

[MDDG07] T. Menzies, A. Dekhtyar, J. Distefano, and J. Greenwald. Problems with precision: A response to “Comments on ‘Data Mining Static Code Attributes to Learn Defect Predictors’ ”. IEEE Transactions on Software Engineering, 33(9): 637–640, September 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4288197.  See

[MGF07, ZZ07].

Mens:2002:SAS

[Men02]  T. Mens.  A state-of-the-

art survey on software merging. IEEE Transactions on Software Engineering, 28(5): 449–462, May 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1000449.

Morisio:2002:SFF

[MET02] M. Morisio, M. Ezran, and C. Tully. Success and failure factors in software reuse. IEEE Transactions on Software Engineering, 28(4):340– 357, April 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=995420.

Morisio:2003:CMS

[MET03] M. Morisio, M. Ezran, and C. Tully. Comments on “More success and failure factors in software reuse”. IEEE Transactions on Software Engineering, 29(5):478–??, May 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1199077.  See

[MD03].

Menasce:2000:MDP

[MG00] D. A. Menasce and H. Gomaa. A method for design and performance modeling of client/server systems. IEEE Transactions on Software Engineering, 26(11):1066– 1085, November 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=881718.

Mills:2002:KBA

[MG02] K. L. Mills and H. Gomaa. Knowledge-based automation of a design method for concurrent systems. IEEE Transactions on Software Engineering, 28(3):228–255, March 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=991319.

Menzies:2007:DMS

[MGF07] T. Menzies, J. Greenwald, and A. Frank. Data mining static code attributes to learn defect predictors. IEEE Transactions on Software Engineering, 33 (1):2–13, January 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4027145. See com-

ments [ZZ07] and response [MDDG07].

Muccini:2004:USA

[MIB04] H. Muccini, P. Inverardi, and A. Bertolino. Using software architecture for code testing. IEEE Transactions on Software Engineering, 30(3): 160–171, March 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1271170.

Milicev:2002:AMT

[Mil02]  D. Milicev.  Automatic

model transformations using extended UML object diagrams in modeling environments. IEEE Transactions on Software Engineering, 28 (4):413–431, April 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=995438.

Milicev:2007:SAA

[Mil07] D. Milicev. On the semantics of associations and association ends in UML. IEEE Transactions on Software Engineering, 33(4):238–251, April 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4123326.

Miner:2006:SGC

[Min06]  A. S. Miner.  Saturation

for a general class of models. IEEE Transactions on Software Engineering, 32(8): 559–570, August 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1703387.

Mitchell:2005:RRC

[Mit05] B. Mitchell. Resolving race conditions in asynchronous partial order scenarios. IEEE Transactions on Software Engineering, 31(9):767–784, September 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1514445.

Mitchell:2008:CCC

[Mit08] B. Mitchell. Characterizing communication channel deadlocks in sequence diagrams. IEEE Transactions on Software Engineering, 34(3):305– 320, May/June 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4509440.

Murtojarvi:2007:DPN

[MJJ^{+}07] M. Murtojarvi, J. Jarvinen, M. Johnsson, T. Leipala, and O. S. Nevalainen. Determining the proper number and price of software licenses. IEEE Transactions on Software Engineering, 33(5):305– 315, May 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4160969.

Mok:2004:STC

[MKL^{+}04] A. K. Mok, P. Konana, Guangtian Liu, Chan-Gun Lee, and Honguk Woo. Specifying timing constraints and composite events: an application in the design of electronic brokerages. IEEE Transactions on Software Engineering, 30(12): 841–858, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377184.

Mantyla:2009:WTD

[ML09] M. V. Mäntylä and C. Lassenius. What types of defects are really discovered in code reviews? IEEE Transactions on Software Engineering, 35(3): 430–448, May/June 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4604671.

Mattsson:2009:LMD

[MLLF09] A. Mattsson, B. Lundell, B. Lings, and B. Fitzgerald. Linking model-driven development and software architecture: A case study. IEEE Transactions on Software Engineering, 35(1):83–93, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4657364.

Mitchell:2006:AMS

[MM06] B. S. Mitchell and S. Mancoridis. On the automatic modularization of software systems using the bunch tool. IEEE Transactions on Software Engineering, 32(3):193– 208, March 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1610610.

McMaster:2008:CSC

[MM08a] S. McMaster and A. Memon. Call-stack coverage for GUI test suite reduction. IEEE Transactions on Software Engineering, 34(1):99–115, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4378345.

Mendes:2008:BNM

[MM08b] E. Mendes and N. Mosley. Bayesian network models for Web effort prediction: A comparative study. IEEE Transactions on Software Engineering, 34(6):723–737, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4589218.

Mosse:2003:NRT

[MMG03] D. Mosse, R. Melhem, and Sunondo Ghosh. A nonpreemptive real-time scheduler with recovery from transient faults and its implementation. IEEE Transactions on Software Engineering, 29(8): 752–767, August 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1223648.

Malek:2005:SAA

[MMRM05] S. Malek, M. Mikic-Rakic, and N. Medvidovic. A styleaware architectural middleware for resource-constrained, distributed systems. IEEE Transactions on Software Engineering, 31(3):256–272, March 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1423996.

Michael:2001:GST

[MMS01] C. C. Michael, G. McGraw, and M. A. Schatz. Generating software test data by evolution. IEEE Transactions on Software Engineering, 27(12):1085– 1110, December 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=988709.

Mattolini:2001:ILR

[MN01] R. Mattolini and P. Nesi. An interval logic for realtime system specification. IEEE Transactions on Software Engineering, 27(3):208– 227, March 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=910858. See com-

ments [FMPR06] and reply [BNR06].

McDaniel:2008:GEI

[MN08] P. McDaniel and B. Nuseibeh. Guest Editors’ introduction: Special section on software engineering for secure systems. IEEE Transactions on Software Engineering, 34(1):3–4, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4444343.

Makaroff:2001:DIV

[MNH01] D. Makaroff, G. Neufeld, and N. Hutchinson. Design and implementation of a VBR continuous media file server. IEEE Transactions on Software Engineering, 27(1):13– 28, January 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=895985.

Murphy:2001:SRM

[MNS01] G. C. Murphy, D. Notkin, and K. J. Sullivan. Software reflexion models: bridging the gap between design and implementation. IEEE Transactions on Software Engineering, 27(4):364–380, April 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=917525.

Maxion:2000:EEH

[MO00] R. A. Maxion and R. T. Olszewski. Eliminating exception handling errors with dependability cases: a comparative, empirical study. IEEE Transactions on Software Engineering, 26(9):888– 906, September 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=877848.

Morisio:2003:GEI

[MO03] M. Morisio and M. Oivo.

Guest Editors’ introduction: software engineering for the wireless Internet. IEEE Transactions on Software Engineering, 29(12):1057–1058, December 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265520.

Molokken-Ostvold:2005:CSP

[MOJ05] K. Molokken-Ostvold and M. Jorgensen. A comparison of software project overruns — flexible versus sequential development models. IEEE Transactions on Software Engineering, 31(9):754– 766, September 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1514444.

Moody:2009:PNT

[Moo09] D. Moody. The “physics” of notations: Toward a scientific basis for constructing visual notations in software engineering. IEEE Transactions on Software Engineering, 35(6):756–779, November/December 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=5353439.

Marongiu:2000:AMS

[MP00] A. Marongiu and P. Palazzari. Automatic mapping of system of N-dimensional affine recurrence equations (SARE) onto distributed memory parallel systems. IEEE Transactions on Software Engineering, 26(3): 262–275, March 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=842951.

Manzoni:2003:IER

[MP03] L. V. Manzoni and R. T. Price.

Identifying extensions required by RUP (rational unified process) to comply with CMM (capability maturity model) levels 2 and 3. IEEE Transactions on Software Engineering, 29(2): 181–192, February 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1178058.

Marcus:2008:UCC

[MPF08] A. Marcus, D. Poshyvanyk, and R. Ferenc. Using the conceptual cohesion of classes for fault prediction in object-oriented systems. IEEE Transactions on Software Engineering, 34(2):287– 300, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4384505.

Masri:2007:EST

[MPL07] W. Masri, A. Podgurski, and D. Leon. An empirical study of test case filtering techniques based on exercising information flows. IEEE Transactions on Software Engineering, 33 (7):454–477, July 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4227828.

Memon:2001:HGT

[MPS01] A. M. Memon, M. E. Pollack, and M. L. Soffa. Hierarchical GUI test case generation using automated planning. IEEE Transactions on Software Engineering, 27(2): 144–155, February 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=908959.

Mori:2002:CSD

[MPS02] G. Mori, F. Paterno, and C. Santoro. CTTE: support for developing and analyzing task models for interactive system design. IEEE Transactions on Software Engineering, 28(8): 797–813, August 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1027801.

Mori:2004:DDM

[MPS04] G. Mori, F. Paterno, and C. Santoro. Design and development of multidevice user interfaces through multiple logical descriptions. IEEE Transactions on Software Engineering, 30(8):507–520, August 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1316868.

Malkhi:2000:SEJ

[MR00] D. Malkhi and M. K. Reiter. Secure execution of Java applets using a remote playground. IEEE Transactions on Software Engineering, 26(12):1197–1209, December 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=888632.

Morisio:2002:QPL

[MRS02] M. Morisio, D. Romano, and I. Stamelos. Quality, productivity, and learning in framework-based development: an exploratory case study. IEEE Transactions on Software Engineering, 28(9):876– 888, September 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1033227.  See

[Kao03].

Murphy:2002:TMU

[MRV02] A. L. Murphy, G. C. Roman, and G. Varghese. Tracking mobile units for dependable message delivery. IEEE Transactions on Software Engineering, 28(5):433–448, May 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1000448.

Marchand:2000:IDP

[MS00a] H. Marchand and M. Samaan. Incremental design of a power transformer station controller using a controller synthesis methodology. IEEE Transactions on Software Engineering, 26(8):729–741, August 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=879811.

Maruyama:2000:ACG

[MS00b] K. Maruyama and K. I. Shima. An automatic class generation mechanism by using method integration. IEEE Transactions on Software Engineering, 26 (5):425–440, May 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=846300.

Myrtveit:2001:ADS

[MSO01] I. Myrtveit, E. Stensrud, and U. H. Olsson. Analyzing data sets with missing data: an empirical evaluation of imputation methods and likelihood-based methods. IEEE Transactions on Software Engineering, 27(11):999– 1013, November 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=965340.

Myrtveit:2005:RVC

[MSS05] I. Myrtveit, E. Stensrud, and M. Shepperd. Reliability and validity in comparative studies of software prediction models. IEEE Transactions on Software Engineering, 31(5): 380–391, May 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1438374.

Medvidovic:2000:CCF

[MT00] N. Medvidovic and R. N. Taylor. A classification and comparison framework for software architecture description languages. IEEE Transactions on Software Engineering, 26 (1):70–93, January 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=825767.

Mens:2004:SSR

[MT04a] T. Mens and T. Tourwe. A survey of software refactoring. IEEE Transactions on Software Engineering, 30(2): 126–139, February 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265817.

Miller:2004:GMP

[MT04b] R. Miller and A. Tripathi.

The guardian model and primitives for exception handling in distributed systems. IEEE Transactions on Software Engineering, 30(12):1008–1022, December 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377194.

Memon:2005:SFD

[MX05] A. M. Memon and Q. Xie. Studying the fault-detection effectiveness of GUI test cases for rapidly evolving software. IEEE Transactions on Software Engineering, 31(10):884– 896, October 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1542069.

Miller:2004:CBM

[MY04] J. Miller and Zhichao Yin. A cognitive-based mechanism for constructing software inspection teams. IEEE Transactions on Software Engineering, 30(11):811–825, November 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1359772.

Moon:2005:ADD

[MYC05] Mikyeong Moon, Keunhyuk Yeom, and Heung Seok Chae. An approach to developing domain requirements as a core asset based on commonality and variability analysis in a product line. IEEE Transactions on Software Engineering, 31 (7):551–569, July 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1492371.

Niu:2003:TSM

[NAD03] Jianwei Niu, J. M. Atlee, and N. A. Day. Template semantics for model-based notations. IEEE Transactions on Software Engineering, 29(10): 866–882, October 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1237169.

Neumann:2002:ENN

[Neu02] D. E. Neumann. An enhanced neural network technique for software risk analysis. IEEE Transactions on Software Engineering, 28(9):904– 912, September 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1033229.

Nebut:2006:ATG

[NFLJ06] C. Nebut, F. Fleurey, Y. Le Traon, and J. M. Jezequel. Automatic test generation: a use case driven approach. IEEE Transactions on Software Engineering, 32(3):140– 155, March 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1610607.

Nan:2009:IBS

[NH09] N. Nan and D. E. Harter. Impact of budget and schedule pressure on software development cycle time and effort. IEEE Transactions on Software Engineering, 35(5):624–637, September/October 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4815275.

Nixon:2000:MPR

[Nix00] B. A. Nixon. Management of performance requirements for information systems. IEEE Transactions on Software Engineering, 26(12):1122–1146, December 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=888627.

Niemela:2005:TAK

[NKL05] E. Niemela, J. Kalaoja, and P. Lago. Toward an architectural knowledge base for wireless service engineering. IEEE Transactions on Software Engineering, 31(5):361– 379, May 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1438373.

Napier:2009:CPP

[NMJ09] N. P. Napier, L. Mathiassen, and R. D. Johnson. Combining perceptions and prescriptions in requirements engineering process assessment: An industrial case study. IEEE Transactions on Software Engineering, 35(5):593–606, September/October 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4967614.

Nottegar:2001:PEM

[NPD01] C. Nottegar, C. Priami, and P. Degano. Performance evaluation of mobile processes via abstract machines. IEEE Transactions on Software Engineering, 27(10):867– 889, October 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=962559.

Nyland:2000:DMD

[NPGM00] L. S. Nyland, J. F. Prins, A. Goldberg, and P. H. Mills. A design methodology for data-parallel applications. IEEE Transactions on Software Engineering, 26(4): 293–314, April 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=844491.

Norman:2009:MCP

[NPPW09] G. Norman, C. Palamidessi, D. Parker, and P. Wu. Model checking probabilistic and stochastic extensions of the π-calculus. IEEE Transactions on Software Engineering, 35(2):209–223, March/April 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4626962.

Ntafos:2001:CRP

[Nta01] S. C. Ntafos. On comparisons of random, partition, and proportional partition testing. IEEE Transactions on Software Engineering, 27(10): 949–960, October 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=962563.

Ngo-The:2009:ORA

[NTR09] A. Ngo-The and G. Ruhe. Optimized resource allocation for software release planning. IEEE Transactions on Software Engineering, 35(1):109–123, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4641940.

Oliveira:2004:RJA

[OAF^{+}04] T. C. Oliveira, P. S. C. Alencar, I. M. Filho, C. J. P. de Lucena, and D. D. Cowan. Response to jiau et al.’s comments. IEEE Transactions on Software Engineering, 30(10): 708–??, October 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339281.

Ocampo:2003:TRP

[OBMP03] A. Ocampo, D. Boggio, J. Munch, and G. Palladino. Toward a reference process for developing wireless Internet services. IEEE Transactions on Software Engineering, 29(12):1122– 1134, December 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265526.

Olague:2007:EVT

[OEGQ07] H. M. Olague, L. H. Etzkorn, S. Gholston, and S. Quattlebaum. Empirical validation of three software metrics suites to predict fault-proneness of object-oriented classes developed using highly iterative or agile software development processes. IEEE Transactions on Software Engineering, 33 (6):402–419, June 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4181709.

Oliveira:2004:SPR

[OFdL^{+}04] T. C. Oliveira, I. M. Filho, C. J. P. de Lucena, P. S. C. Alencar, and D. D. Cowan. Software process representation and analysis for framework instantiation. IEEE Transactions on Software Engineering, 30(3): 145–159, March 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1271169. See com-

ments [JKS04].

Offutt:2002:CP

[OJ02] J. Offutt and J. Jezequel. Call for papers. IEEE Transactions on Software Engineering, 28(4):432, April 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=995443.

Ostrand:2005:PLN

[OWB05] T. J. Ostrand, E. J. Weyuker, and R. M. Bell. Predicting the location and number of faults in large software systems. IEEE Transactions on Software Engineering, 31(4): 340–355, April 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1435354.

Petrenko:2004:CCE

[PBG04] A. Petrenko, S. Boroday, and R. Groz. Confirming configurations in EFSM testing. IEEE Transactions on Software Engineering, 30(1):29– 42, January 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265734.

Peleg:2000:MMP

[PD00] M. Peleg and D. Dori. The model multiplicity problem: experimenting with real-time specification methods. IEEE Transactions on Software Engineering, 26(8):742–759, August 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=879812.

Pai:2007:EAS

[PD07] G. J. Pai and J. Bechta Dugan.

Empirical analysis of software fault content and fault proneness using Bayesian methods. IEEE Transactions on Software Engineering, 33(10):675– 686, October 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4302779.

Prakash:2000:APS

[PDB00] S. Prakash, E. Deelman, and R. Bagrodia. Asynchronous parallel simulation of parallel programs. IEEE Transactions on Software Engineering, 26 (5):385–400, May 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=846297.

Poshyvanyk:2007:FLU

[PGM^{+}07] D. Poshyvanyk, Y. G. Guéhéneuc, A. Marcus, G. Antoniol, and V. Rajlich. Feature location using probabilistic ranking of methods based on execution scenarios and information retrieval. IEEE Transactions on Software Engineering, 33 (6):420–432, June 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4181710.

Pike:2006:NIA

[Pik06] L. Pike. A note on inconsistent axioms in Rushby’s “Systematic formal verification for fault-tolerant time-triggered algorithms”. IEEE Transactions on Software Engineering, 32 (5):347–348, May 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1642681.

Pelliccione:2009:CFD

[PIM09] P. Pelliccione, P. Inverardi, and H. Muccini. CHARMY: A framework for designing and verifying architectural specifications. IEEE Transactions on Software Engineering, 35(3): 325–346, May/June 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4711062.

Pickin:2007:TSU

[PJJ^{+}07] S. Pickin, C. Jard, T. Jeron, J. M. Jezequel, and Y. Le Traon. Test synthesis from UML models of distributed software. IEEE Transactions on Software Engineering, 33 (4):252–269, April 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4123327.

Parnas:2003:RIS

[PL03] D. L. Parnas and M. Lawford. The role of inspection in software quality assurance. IEEE Transactions on Software Engineering, 29(8): 674–676, August 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1223642.

Pietro:2000:GES

[PMM00] P. S. Pietro, A. Morzenti, and S. Morasca. Generation of execution sequences for modular time critical systems. IEEE Transactions on Software Engineering, 26(2):

128–149, February 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=841114.

Parsons:2008:EIC

[PMT^{+}08] T. Parsons, A. Mos, M. Trofin, T. Gschwind, and J. Murphy. Extracting interactions in component-based systems. IEEE Transactions on Software Engineering, 34(6):783–799, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4599581.

Peters:2002:RBM

[PP02] D. K. Peters and D. L. Parnas. Requirements-based monitors for real-time systems. IEEE Transactions on Software Engineering, 28(2):146– 158, February 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=988496.

Prowell:2003:FSB

[PP03] S. J. Prowell and J. H. Poore. Foundations of sequence-based software specification. IEEE Transactions on Software Engineering, 29(5):417–429, May 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1199071.

Purushothaman:2005:TUR

[PP05] R. Purushothaman and D. E.

Perry. Toward understanding the rhetoric of small source code changes. IEEE Transactions on Software Engineering, 31(6):511–526, June 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1463233.

Perry:2002:RII

[PPW^{+}02] D. E. Perry, A. Porter, M. W.

Wade, L. G. Votta, and J. Perpich. Reducing inspection interval in large-scale software development. IEEE Transactions on Software Engineering, 28(7):695–705, July 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1019483.

Padberg:2004:UML

[PRS04] F. Padberg, T. Ragg, and R. Schoknecht. Using machine learning for estimating the defect content after an inspection. IEEE Transactions on Software Engineering, 30 (1):17–28, January 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1265733.

Perry:2000:CTEb

[PRT00a] D. E. Perry, A. Romanovsky, and A. Tripathi. Current trends in exception handling. IEEE Transactions on Software Engineering, 26(9):817– 819, September 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=877843.

Perry:2000:CTEa

[PRT00b] D. E. Perry, A. Romanovsky, and A. Tripathi. Current trends in exception handling. IEEE Transactions on Software Engineering, 26(10):921– 922, October 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=879816.

Parsons:2004:CHS

[PS04] J. Parsons and C. Saunders. Cognitive heuristics in software engineering applying and extending anchoring and adjustment to artifact reuse. IEEE Transactions on Software Engineering, 30(12):873– 888, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1377186.

Paulson:2004:ESO

[PSE04] J. W. Paulson, G. Succi, and A. Eberlein. An empirical study of open-source and closed-source software products. IEEE Transactions on Software Engineering, 30(4): 246–256, April 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1274044.

Petriu:2000:ABP

[PSJ00] D. Petriu, C. Shousha, and A. Jalnapurkar. Architecturebased performance analysis applied to a telecommunication system. IEEE Transactions on Software Engineering, 26(11):1049–1065, November 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=881717.

Popov:2003:EBR

[PSMK03] P. Popov, L. Strigini, J. May, and S. Kuball. Estimating bounds on the reliability of diverse systems. IEEE Transactions on Software Engineering, 29(4):345–359, April 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1191798.

Pendharkar:2005:PMP

[PSR05] P. C. Pendharkar, G. H. Subramanian, and J. A. Rodger. A probabilistic model for predicting software development effort. IEEE Transactions on Software Engineering, 31 (7):615–624, July 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1492375.

Prechelt:2001:EME

[PU01] L. Prechelt and B. Unger. An experiment measuring the effects of personal software process (psp) training. IEEE Transactions on Software Engineering, 27(5):465– 472, May 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=922716.

Prechelt:2002:TCE

[PULPT02] L. Prechelt, B. Unger-Lamprecht, M. Philippsen, and W. F. Tichy. Two controlled experiments assessing the usefulness of design pattern documentation in program maintenance. IEEE Transactions on Software Engineering, 28 (6):595–606, June 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1010061.

Prechelt:2001:CEM

[PUT^{+}01] L. Prechelt, B. Unger, W. F. Tichy, P. Brossler, and L. G. Votta. A controlled experiment in maintenance: comparing design patterns to simpler solutions. IEEE Transactions on Software Engineering, 27(12):1134–1144, December 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=988711.

Plasil:2002:BPS

[PV02] F. Plasil and S. Visnovsky.

Behavior protocols for software components. IEEE Transactions on Software Engineering, 28(11):1056–1076, November 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1049404.

Porter:2007:SPI

[PYM^{+}07] A. Porter, C. Yilmaz, A. M.

Memon, D. C. Schmidt, and B. Natarajan. Skoll: A process and infrastructure for distributed continuous quality assurance. IEEE Transactions on Software Engineering, 33(8): 510–525, August 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4267023.

Ravindran:2002:EDR

[Rav02] B. Ravindran. Engineering dynamic real-time distributed systems: architecture, system description language, and middleware. IEEE Transactions on Software Engineering, 28 (1):30–57, January 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=979988.

Rovegaard:2008:ESV

[RAW08] P. Rovegård, L. Angelis, and C. Wohlin. An empirical study on views of importance of change impact analysis issues. IEEE Transactions on Software Engineering, 34(4):516– 530, July/August 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4527253.

Ruthruff:2006:IFL

[RBR06] J. R. Ruthruff, M. Burnett, and G. Rothermel. Interactive fault localization techniques in a spreadsheet environment. IEEE Transactions on Software Engineering, 32 (4):213–239, April 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1628969.

Robillard:2004:HED

[RCM04] M. P. Robillard, W. Coelho, and G. C. Murphy. How effective developers investigate source code: an exploratory study. IEEE Transactions on Software Engineering, 30(12): 889–903, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377187.

Ren:2006:IFC

[RCR06] Xiaoxia Ren, O. C. Chesley, and B. G. Ryder. Identifying failure causes in Java programs: An application of change impact analysis. IEEE Transactions on Software Engineering, 32(9):718– 732, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1707669.

Rutherford:2008:ETS

[RCW08] M. J. Rutherford, A. Carzaniga, and A. L. Wolf. Evaluating test suites and adequacy criteria using simulation-based models of distributed systems. IEEE Transactions on Software Engineering, 34(4):452– 470, July/August 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4527254.

Rothenberger:2003:SSR

[RDKN03] M. A. Rothenberger, K. J.

Dooley, U. R. Kulkarni, and N. Nada. Strategies for software reuse: a principal component analysis of reuse practices. IEEE Transactions on Software Engineering, 29(9): 825–837, September 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1232287.

Reiss:2006:IMS

[Rei06] S. P. Reiss. Incremental maintenance of software artifacts. IEEE Transactions on Software Engineering, 32(9):682– 697, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1707667.

Roberts:1998:FII

[RGFR98] T. L. Roberts, M. L. Gibson, K. T. Fields, and R. K. Rainer. Factors that impact implementing a system development methodology. IEEE Transactions on Software Engineering, 24(8): 640–649, August 1998. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=707699. See com-

ments [YSW^{+}01].

Roberts:2001:RCF

[RGRF01] T. L. Roberts, M. L. Gibson, R. K. Rainer, and K. T. Fields. Response to “Comments on factors that impact the implementation of a systems development methodology”. IEEE Transactions on Software Engineering, 27(3): 282–286, March 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=910864. See [RGFR98,

YSW^{+}01].

Riemenschneider:2002:ESD

[RHD02] C. K. Riemenschneider, B. C.

Hardgrave, and F. D. Davis. Explaining software developer acceptance of methodologies: a comparison of five theoretical models. IEEE Transactions on Software Engineering, 28(12):1135–1145, December 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1158287.

Ramesh:2001:TRM

[RJ01] B. Ramesh and M. Jarke. Toward reference models for requirements traceability. IEEE Transactions on Software Engineering, 27(1):58–93, January 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=895989.

Ropponen:2000:CSD

[RL00] J. Ropponen and K. Lyytinen. Components of software development risk: how to address them? A project manager survey. IEEE Transactions on Software Engineering, 26(2): 98–112, February 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=841112.

Rallis:2001:RES

[RL01] N. E. Rallis and Z. F. Lansdowne. Reliability estimation for a software system with sequential independent reviews. IEEE Transactions on Software Engineering, 27(12):1057– 1061, December 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=988707.

Rajgopal:2002:MOT

[RM02] J. Rajgopal and M. Mazumdar. Modular operational test plans for inferences on software reliability based on a Markov model. IEEE Transactions on Software Engineering, 28 (4):358–363, April 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=995424.

Rountev:2004:FCA

[RMR04] A. Rountev, A. Milanova, and B. G. Ryder. Fragment class analysis for testing of polymorphism in Java software. IEEE Transactions on Software Engineering, 30(6):372– 387, June 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1321060.

Rothermel:2005:GEI

[Rot05] G. Rothermel. Guest Editor’s introduction: 2004 IEEE International Symposium on Software Testing and Analysis. IEEE Transactions on Software Engineering, 31(4): 273–274, April 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1435349.

Ramesh:2000:MCS

[RP00] Sridhar Ramesh and H. G.

Perros. A multilayer clientserver queueing network model with synchronous and asynchronous messages. IEEE Transactions on Software Engineering, 26(11):1086–1100, November 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=881719.

Rauber:2000:TAD

[RR00] T. Rauber and G. Runger. A transformation approach to derive efficient parallel implementations. IEEE Transactions on Software Engineering, 26 (4):315–339, April 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=844492.

Rosti:2002:MPA

[RSSS02] E. Rosti, G. Serazzi, E. Smirni, and M. S. Squillante. Models of parallel applications with large computation and I/O requirements. IEEE Transactions on Software Engineering, 28(3): 286–307, March 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=991321.

Rothermel:2001:PTC

[RUCH01] G. Rothermel, R. H. Untch, Chengyun Chu, and M. J. Harrold. Prioritizing test cases for regression testing. IEEE Transactions on Software Engineering, 27(10):929– 948, October 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=962562.

Rushby:1999:SFV

[Rus99] J. Rushby. Systematic formal verification for faulttolerant time-triggered algorithms. IEEE Transactions on Software Engineering, 25(5):651–660, September/October 1999. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?

arnumber=815324. See note

[Pik06].

Siegel:2002:IPI

[SA02] S. F. Siegel and G. S. Avrunin.

Improving the precision of INCA by eliminating solutions with spurious cycles. IEEE Transactions on Software Engineering, 28(2):115– 128, February 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=988494.

Stewart:2003:TAF

[SA03] D. B. Stewart and G. Arora.

A tool for analyzing and fine tuning the real-time properties of an embedded system. IEEE Transactions on Software Engineering, 29(4):311– 326, April 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1191796.

Schmerl:2006:DAR

[SAG^{+}06] B. Schmerl, J. Aldrich, D. Garlan, R. Kazman, and Hong Yan. Discovering architectures from running systems. IEEE Transactions on Software Engineering, 32(7):454– 466, July 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1677532.

Santone:2003:HSL

[San03] A. Santone. Heuristic search + local model checking in selective mu-calculus. IEEE Transactions on Software Engineering, 29(6):510–523, June 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1205179.

Satoh:2003:TFM

[Sat03] I. Satoh. A testing framework for mobile computing software. IEEE Transactions on Software Engineering, 29(12):1112– 1121, December 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1265525.

Stuart:2001:SVB

[SBMJ01] D. A. Stuart, M. Brockmeyer, A. K. Mok, and F. Jahanian. Simulation-verification: biting at the state explosion problem. IEEE Transactions on Software Engineering, 27 (7):599–617, July 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=935853.

Succi:2001:AES

[SBV01] G. Succi, L. Benedicenti, and T. Vernazza. Analysis of the effects of software reuse on customer satisfaction in an RPG environment. IEEE Transactions on Software Engineering, 27(5):473–479, May 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=922717.

Shepperd:2001:PSD

[SC01] M. Shepperd and M. Cartwright.

Predicting with sparse data. IEEE Transactions on Software Engineering, 27(11):987– 998, November 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=965339.

Sobel:2002:FMA

[SC02] A. E. K. Sobel and M. R.

Clarkson. Formal methods application: an empirical tale of software development. IEEE Transactions on Software Engineering, 28(3): 308–320, March 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=991322. See com-

ments [BT03].

Sobel:2003:RCF

[SC03] A. E. K. Sobel and M. R.

Clarkson. Response to “Comments on ‘Formal methods application: an empirical tale of software development’’. IEEE Transactions on Software Engineering, 29(6):572– 575, June 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1205184. See [SC02,

BT03].

Sengupta:2006:TMS

[SC06] B. Sengupta and R. Cleaveland. Triggered message sequence charts. IEEE Transactions on Software Engineering, 32(8): 587–607, August 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1703389.

Spiekermann:2009:EP

[SC09] S. Spiekermann and L. F. Cranor. Engineering privacy. IEEE Transactions on Software Engineering, 35(1):67–82, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4657365.

Stuermer:2007:STM

[SCDP07] I. Stuermer, M. Conrad, H. Doerr, and P. Pepper. Systematic testing of model-based code generators. IEEE Transactions on Software Engineering, 33(9): 622–634, September 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4288195.

Sproston:2006:BBM

[SD06a] J. Sproston and S. Donatelli. Backward bisimulation in Markov chain model checking. IEEE Transactions on Software Engineering, 32(8): 531–546, August 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1703385.

Sun:2006:DSI

[SD06b] J. Sun and J. S. Dong. Design synthesis from interaction and state-based specifications. IEEE Transactions on Software Engineering, 32 (6):349–364, June 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1650212.

Snavely:2005:UUU

[SDA05] N. Snavely, S. Debray, and G. R. Andrews. Unpredication, unscheduling, unspeculation: reverse engineering Itanium executables. IEEE Transactions on Software Engineering, 31(2): 99–115, February 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1401927.

Schafer:2006:SSE

[SEHM06] T. Schafer, M. Eichberg, M. Haupt, and M. Mezini.

The SEXTANT software exploration tool. IEEE Transactions on Software Engineering, 32(9): 753–768, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1707671.

Selby:2005:ERB

[Sel05]  R. W. Selby.  Enabling

reuse-based software development of large-scale systems. IEEE Transactions on Software Engineering, 31(6):495– 510, June 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1463232.

Strike:2001:SCE

[SEM01] K. Strike, K. El Emam, and N. Madhavji. Software cost estimation with incomplete data. IEEE Transactions on Software Engineering, 27(10):890– 908, October 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=962560.

Shin:2000:EDM

[SG00] Miyoung Shin and A. L. Goel. Empirical data modeling in software engineering using radial basis functions. IEEE Transactions on Software Engineering, 26(6):567– 576, June 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=852743.

Shaw:2003:MDR

[SGMB03] S. C. Shaw, M. Goldstein, M. Munro, and E. Burd. Moral dominance relations for program comprehension. IEEE Transactions on Software Engineering, 29(9):851– 863, September 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1232289.

Sinha:2000:ATP

[SH00] S. Sinha and M. J. Harrold.

Analysis and testing of programs with exception handling constructs. IEEE Transactions on Software Engineering, 26(9): 849–871, September 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=877846.

Srivastava:2008:MIH

[SHFJ08] S. Srivastava, M. Hicks, J. S. Foster, and P. Jenkins. Modular information hiding and type-safe linking for C. IEEE Transactions on Software Engineering, 34(3):357– 376, May/June 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4497211.

Sjoeberg:2005:SCE

[SHH^{+}05] Dag I. K. Sjøberg, J. E. Hannay, O. Hansen, V. B. Kampenes, A. Karahasanovic, N. K. Liborg, and A. C. Rekdal. A survey of controlled experiments in software engineering. IEEE Transactions on Software Engineering, 31(9):733– 753, September 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1514443.

Smith:2001:ESU

[SHP01] R. K. Smith, J. E. Hale, and A. S. Parrish. An empirical study using task assignment patterns to improve the accuracy of software effort estimation. IEEE Transactions on Software Engineering, 27(3): 264–271, March 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=910861.

Sharma:2006:AWP

[SJJ06] Naveen Sharma, Padmaja Joshi, and R. K. Joshi. Applicability of Weyuker’s Property 9 to object oriented metrics. IEEE Transactions on Software Engineering, 32(3): 209–211, March 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1610611.

Sauer:2000:ESD

[SJLY00] C. Sauer, D. R. Jeffery, L. Land, and P. Yetton. The effectiveness of software development technical reviews: a behaviorally motivated program of research. IEEE Transactions on Software Engineering, 26(1):1–14, January 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=825763.

Shepperd:2001:CSP

[SK01] M. Shepperd and G. Kadoda.

Comparing software prediction techniques using simulation. IEEE Transactions on Software Engineering, 27(11):1014– 1022, November 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=965341.

Subramanyam:2003:EAC

[SK03] R. Subramanyam and M. S.

Krishnan. Empirical analysis of CK metrics for objectoriented design complexity: implications for software defects. IEEE Transactions on Software Engineering, 29(4):297– 310, April 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1191795.

Sarkar:2008:MMQ

[SKR08] S. Sarkar, A. C. Kak, and G. M. Rama. Metrics for measuring the quality of modularization of large-scale objectoriented software. IEEE Transactions on Software Engineering, 34(5):700–720, September/October 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4623684.

Shull:2000:IRT

[SLB00] F. Shull, F. Lanubile, and V. R. Basili. Investigating reading techniques for objectoriented framework learning. IEEE Transactions on Software Engineering, 26(11):1101– 1118, November 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=881720.

Strein:2007:EMM

[SLLL07a] D. Strein, R. Lincke, J. Lundberg, and W. Löwe. An extensible meta-model for program analysis. IEEE Transactions on Software Engineering, 33(9): 592–607, September 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4288193.

Strein:2007:EMP

[SLLL07b] D. Strein, R. Lincke, J. Lundberg, and W. Löwe. An extensible metamodel for program analysis (abstract only). IEEE Transactions on Software Engineering, 33(12):799, December 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4375379.

Sherman:2003:KEL

[SM03a] A. T. Sherman and D. A.

McGrew. Key establishment in large dynamic groups using one-way function trees. IEEE Transactions on Software Engineering, 29(5):444– 458, May 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1199073.

Stensrud:2003:IHP

[SM03b] E. Stensrud and I. Myrtveit.

Identifying high performance ERP projects. IEEE Transactions on Software Engineering, 29(5):398–416, May 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1199070.

Seaman:2003:UIE

[SMBK03] C. B. Seaman, M. G. Mendonca, V. R. Basili, and Y. M.

Kim. User interface evaluation and empirically-based evolution of a prototype experience management tool. IEEE Transactions on Software Engineering, 29(9):838– 850, September 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1232288.

Saougkos:2007:RJB

[SMBZ07] D. Saougkos, G. Manis, K. Blekas, and A. V. Zarras. Revisiting Java bytecode compression for embedded and mobile computing environments. IEEE Transactions on Software Engineering, 33(7):478– 495, July 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4227829.

Sillito:2008:AAQ

[SMD08] J. Sillito, G. C. Murphy, and K. De Volder. Asking and answering questions during a programming change task. IEEE Transactions on Software Engineering, 34(4):434– 451, July/August 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4497212.

Souter:2003:CCD

[SP03] A. L. Souter and L. L. Pollock. The construction of contextual def-use associations for object-oriented systems. IEEE Transactions on Software Engineering, 29(11):1005– 1018, November 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1245302.

Shinjo:2005:AEP

[SP05] Y. Shinjo and C. Pu. Achieving efficiency and portability in systems software: a case study on POSIXcompliant multithreaded programs. IEEE Transactions on Software Engineering, 31(9): 785–800, September 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1514446.

Spinellis:2003:GAT

[Spi03] D. Spinellis. Global analysis and transformations in preprocessed languages. IEEE Transactions on Software Engineering, 29(11):1019–1030, November 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1245303.

Sharp:2006:SAO

[SR06] M. Sharp and A. Rountev.

Static analysis of object references in RMI-based Java software. IEEE Transactions on Software Engineering, 32(9): 664–681, September 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1707666.

Sawyer:2005:SKA

[SRC05] P. Sawyer, P. Rayson, and K. Cosh. Shallow knowledge as an aid to deep understanding in early phase requirements engineering. IEEE Transactions on Software Engineering, 31(11): 969–981, November 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1556555.

Sarkar:2007:ABI

[SRK07] S. Sarkar, G. M. Rama, and A. C. Kak. API-based and information-theoretic metrics for measuring the quality of software modularization. IEEE Transactions on Software Engineering, 33(1): 14–32, January 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4027146.

Storey:2009:HSD

[SRS^{+}09] M. A. Storey, J. Ryall, J. Singer, D. Myers, L. T. Cheng, and M. Muller. How software developers use tagging to support reminding and refinding. IEEE Transactions on Software Engineering, 35(4):470–483, July/ August 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4782972.

Song:2006:SDA

[SSCM06] Qinbao Song, M. Shepperd, M. Cartwright, and C. Mair. Software defect association mining and defect correction effort prediction. IEEE Transactions on Software Engineering, 32(2):69–82, February 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1599417.

Silva:2001:CCO

[SSE01] D. M. Silva, K. Schwan, and G. Eisenhauer. CTK: configurable object abstractions for multiprocessors. IEEE Transactions on Software Engineering, 27(6):531–549, June 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=926175.

Sampath:2007:ACA

[SSG^{+}07] S. Sampath, S. Sprenkle, E. Gibson, L. Pollock, and A. Souter Greenwald. Applying concept analysis to usersession-based testing of Web applications. IEEE Transactions on Software Engineering, 33(10):643–658, October 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=4302777.

Singer:2002:EIE

[SV02] J. Singer and N. G. Vinson. Ethical issues in empirical studies of software engineering. IEEE Transactions on Software Engineering, 28(12):1171– 1180, December 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1158289.

Stubblebine:2002:ALF

[SW02] S. G. Stubblebine and R. N. Wright. An authentication logic with formal semantics supporting synchronization, revocation, and recency. IEEE Transactions on Software Engineering, 28(3):256– 285, March 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=991320.

Shen:2009:ESA

[SWE09] W. Shen, K. Wang, and A. Egyed. An efficient and scalable approach to correct class model refinement. IEEE Transactions on Software Engineering, 35(4):515– 533, July/August 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4815278.

Shoham:2008:SSM

[SYFP08] S. Shoham, E. Yahav, S. J.

Fink, and M. Pistoia. Static specification mining using automata-based abstractions. IEEE Transactions on Software Engineering, 34(5):651–666, September/October 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4589215.

Stankovic:2002:DPP

[SZ02] N. Stankovic and Kang Zhang.

A distributed parallel programming framework. IEEE Transactions on Software Engineering, 28(5):478–493, May 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1000451.

Traore:2004:ESR

[TA04] I. Traore and D. B. Aredo.

Enhancing structured review with model-based verification. IEEE Transactions on Software Engineering, 30(11):736– 753, November 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1359768.

Tonella:2005:RAI

[TC05] P. Tonella and M. Ceccato. Refactoring the aspectizable interfaces: an empirical assessment. IEEE Transactions on Software Engineering, 31(10): 819–832, October 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1542065.

Tsantalis:2009:IMM

[TC09] N. Tsantalis and A. Chatzigeorgiou. Identification of move method refactoring opportunities. IEEE Transactions on Software Engineering, 35(3): 347–367, May/June 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4752842.

Teasley:2002:RSD

[TCKO02] S. D. Teasley, L. A. Covi, M. S. Krishnan, and J. S. Olson. Rapid software development through team collocation. IEEE Transactions on Software Engineering, 28 (7):671–683, July 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1019481.

Tsantalis:2005:PPC

[TCS05] N. Tsantalis, A. Chatzigeorgiou, and G. Stephanides. Predicting the probability of change in object-oriented systems. IEEE Transactions on Software Engineering, 31(7): 601–614, July 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1492374.

Tsantalis:2006:DPD

[TCSH06] N. Tsantalis, A. Chatzigeorgiou, G. Stephanides, and S. T. Halkidis. Design pattern detection using similarity scoring. IEEE Transactions on Software Engineering, 32(11): 896–909, November 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4015512.

Teruel:2003:WDG

[TFD03] E. Teruel, G. Franceschinis, and M. De Pierro. Welldefined generalized stochastic Petri nets: a net-level method to specify priorities. IEEE Transactions on Software Engineering, 29(11):962– 973, November 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1245298.

Tomer:2004:ESR

[TGK^{+}04] A. Tomer, L. Goldin, T. Kuflik, E. Kimchi, and S. R. Schach. Evaluating software reuse alternatives: a model and its application to an industrial case study. IEEE Transactions on Software Engineering, 30(9): 601–612, September 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1324647.

Tan:2003:FWP

[THY03] Hee Beng Kuan Tan, Lun Hao, and Yong Yang. On formalization of the wholepart relationship in the unified modeling language. IEEE Transactions on Software Engineering, 29(11):1054–1055, November 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1245307.

Tian:2002:BRA

[Tia02]  J. Tian.  Better reliabil-

ity assessment and prediction through data clustering. IEEE Transactions on Software Engineering, 28(10):997– 1007, October 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1041055.

Tai:2002:TGS

[TL02] Kuo-Chung Tai and Yu Lei.

A test generation strategy for pairwise testing. IEEE Transactions on Software Engineering, 28(1):109–111, January 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=979992.

Tan:2005:CUF

[TM05] Y. Tan and V. S. Mookerjee. Comparing uniform and flexible policies for software maintenance and replacement. IEEE Transactions on Software Engineering, 31(3):238– 255, March 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1423995.

Tonella:2001:CAM

[Ton01] P. Tonella. Concept analysis for module restructuring. IEEE Transactions on Software Engineering, 27(4): 351–363, April 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=917524.

Tonella:2003:UCL

[Ton03] P. Tonella. Using a concept lattice of decomposition slices for program understanding and impact analysis. IEEE Transactions on Software Engineering, 29(6): 495–509, June 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1205178.

Tian:2004:EWS

[TRL04] J. Tian, S. Rudraraju, and Zhao Li. Evaluating Web software reliability based on workload and failure data extracted from server logs. IEEE Transactions on Software Engineering, 30(11):754– 769, November 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1359769.

Thelin:2003:ECU

[TRW03] T. Thelin, P. Runeson, and C. Wohlin. An experimental comparison of usage-based and checklist-based reading. IEEE Transactions on Software Engineering, 29(8):687– 704, August 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1223644.

Talia:2000:GEIb

[TSJ00a] D. Talia, P. K. Srimani, and M. Jazayeri. Guest Editors’ introduction: special issues on architectureindependent languages and software tools for parallel processing. IEEE Transactions on Software Engineering, 26 (3):193–196, March 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=842946.

Talia:2000:GEIa

[TSJ00b] D. Talia, P. K. Srimani, and M. Jazayeri. Guest Editors’ introduction: special issues on architectureindependent languages and software tools for parallel processing. IEEE Transactions on Software Engineering, 26 (4):289–292, April 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=844490.

Turski:2002:RMS

[Tur02] W. M. Turski. The reference model for smooth growth of software systems revisited. IEEE Transactions on Software Engineering, 28(8):814– 815, August 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1027802.

Telang:2007:EAI

[TW07] R. Telang and S. Wattal. An empirical analysis of the impact of software vulnerability announcements on firm stock price. IEEE Transactions on Software Engineering, 33(8): 544–557, August 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4267025.

Tan:2006:STF

[TYB06] H. B. K. Tan, Y. Yang, and L. Bian. Systematic transformation of functional analysis model into OO design and implementation. IEEE Transactions on Software Engineering, 32(2):111– 135, February 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1599420.

Uchitel:2009:SPB

[UBC09] S. Uchitel, G. Brunet, and M. Chechik. Synthesis of partial behavior models from properties and scenarios. IEEE Transactions on Software Engineering, 35(3):384–406, May/ June 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=4721439.

Uchitel:2005:GES

[UBKW05] S. Uchitel, M. Broy, I. H. Kruger, and J. Whittle. Guest editorial: Special section on interaction and state-based modeling. IEEE Transactions on Software Engineering, 31(12):

997–998, December 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1566602.

Uchitel:2003:SBM

[UKM03] S. Uchitel, J. Kramer, and J. Magee. Synthesis of behavioral models from scenarios. IEEE Transactions on Software Engineering, 29(2): 99–115, February 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1178048.

VanRompaey:2007:DTS

[VDDR07] B. Van Rompaey, B. Du Bois, S. Demeyer, and M. Rieger. On the detection of test smells: A metrics-based approach for general fixture and eager test. IEEE Transactions on Software Engineering, 33(12):800– 817, December 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4359471.

Verdickt:2005:AIM

[VDGD05] T. Verdickt, B. Dhoedt, F. Gielen, and P. Demeester. Automatic inclusion of middleware performance attributes into architectural UML software models. IEEE Transactions on Software Engineering, 31(8): 695–711, August 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1498773.

vanderHoek:2002:TCM

[vdHCHW02] A. van der Hoek, A. Carzaniga, D. Heimbigner, and A. L. Wolf. A testbed for configuration management policy programming. IEEE Transactions on Software Engineering, 28(1): 79–99, January 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=979990.

vanderMeulen:2008:ESD

[vdMR08] M. J. P. van der Meulen and M. A. Revilla. The effectiveness of software diversity in a large population of programs. IEEE Transactions on Software Engineering, 34(6):753–764, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4604670.

vanDeursen:2005:GEI

[vDS05] A. van Deursen and E. Stroulia. Guest Editor’s introduction: 10th working conference on reverse engineering. IEEE Transactions on Software Engineering, 31(2): 97–98, February 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1401926.

Vandewoude:2007:TLD

[VEBD07] Y. Vandewoude, P. Ebraert, Y. Berbers, and T. D’Hondt. Tranquility: A low disruptive alternative to quiescence for ensuring safe dynamic updates. IEEE Transactions on Software Engineering, 33(12): 856–868, December 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4359466.

Villadangos:2003:SAR

[VFdM^{+}03] J. Villadangos, F. Farina, J. R. Gonzalez de Mendivil, J. R. Garitagoitia, and A. Cordoba. A safe algorithm for resolving OR deadlocks. IEEE Transactions on Software Engineering, 29(7):608– 622, July 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1214325.

Vicario:2001:SAD

[Vic01]  E. Vicario.  Static anal-

ysis and dynamic steering of time-dependent systems. IEEE Transactions on Software Engineering, 27(8):728– 748, August 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=940727.

Vegas:2009:MSE

[VJB09] S. Vegas, N. Juristo, and V. R. Basili. Maturing software engineering knowledge through classifications: A case study on unit testing techniques. IEEE Transactions on Software Engineering, 35(4):551– 565, July/August 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4775907.

vanLamsweerde:2000:HOG

[vLL00] A. van Lamsweerde and E. Letier. Handling obstacles in goal-oriented requirements engineering. IEEE Transactions on Software Engineering, 26(10):978–1005, October 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=879820.

Volzer:2004:STI

[VMA^{+}04] H. Volzer, A. MacDonald, B. Atchison, A. Hanlon, P. Lindsay, and P. Strooper. SubCM: a tool for improved visibility of software change in an industrial setting. IEEE Transactions on Software Engineering, 30(10):675–693, October 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1339278.

vanMoorsel:2006:ARM

[vMW06] A. P. A. van Moorsel and K. Wolter. Analysis of restart mechanisms in software systems. IEEE Transactions on Software Engineering, 32(8): 547–558, August 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1703386.

vanOmmering:2005:SRP

[vO05]  R. van Ommering.  Soft-

ware reuse in product populations. IEEE Transactions on Software Engineering, 31 (7):537–550, July 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1492370.

Vokac:2004:DFD

[Vok04] M. Vokac. Defect frequency and design patterns: an empirical study of industrial code. IEEE Transactions on Software Engineering, 30(12):904– 917, December 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377188.

Vitharana:2003:CMG

[VR03] P. Vitharana and K. Ramamurthy. Computer-mediated group support, anonymity, and the software inspection process: an empirical investigation. IEEE Transactions on Software Engineering, 29(2): 167–180, February 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1178054.

Vicario:2009:USS

[VSC09] E. Vicario, L. Sassoli, and L. Carnevali. Using stochastic state classes in quantitative evaluation of dense-time reactive systems. IEEE Transactions on Software Engineering, 35(5):703–719, September/October 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5010440.

Vitharana:2003:KBR

[VZJ03] P. Vitharana, F. M. Zahedi, and H. Jain. Knowledgebased repository scheme for storing and retrieving business components: a theoretical design and an empirical analysis. IEEE Transactions on Software Engineering, 29 (7):649–664, July 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1214328.

Wang:2005:SPS

[Wan05] Farn Wang. Symbolic parametric safety analysis of linear hybrid systems with BDDlike data-structures. IEEE Transactions on Software Engineering, 31(1):38–51, January 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1392719.

Wang:2005:AAO

[WBY05] D. Wang, F. B. Bastani, and L. I. Yen. Automated aspect-oriented decomposition of process-control systems for ultra-high dependability assurance. IEEE Transactions on Software Engineering, 31(9): 713–732, September 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1514442.

Weber:2005:FSB

[WCD^{+}05] S. Weber, H. Chan, L. Degenaro, J. Diament, A. Fokoue-Nkoutche, and I. Rouvellou. Fusion: a system for business users to manage program variability. IEEE Transactions on Software Engineering, 31 (7):570–587, July 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1492372.

Wu:2002:EDD

[WCJ02] Hui Wu, Wei-Ngan Chin, and J. Jaffar. An efficient distributed deadlock avoidance algorithm for the AND model. IEEE Transactions on Software Engineering, 28(1):18– 29, January 2002. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=979987.

Wong:2004:XMS

[WCL04] E. Y. C. Wong, A. T. S. Chan, and Hong Va Leong. Xstream: a middleware for streaming XML contents over wireless environments. IEEE Transactions on Software Engineering, 30(12):918–935, December 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1377189.

Wooff:2002:BGM

[WGC02] D. A. Wooff, M. Goldstein, and F. P. A. Coolen. Bayesian graphical models for software testing. IEEE Transactions on Software Engineering, 28 (5):510–525, May 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1000453.

Williams:2005:AMS

[WH05] C. C. Williams and J. K.

Hollingsworth. Automatic mining of source code repositories to improve bug finding techniques. IEEE Transactions on Software Engineering, 31 (6):466–480, June 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1463230.

Wang:2006:TIA

[WHY06] Farn Wang, Geng-Dian Huang, and Fang Yu. TCTL inevitability analysis of dense-time systems: From theory to engineering. IEEE Transactions on Software Engineering, 32 (7):510–526, July 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1677536.

Whisnant:2004:EAB

[WIK^{+}04] K. Whisnant, R. K. Iyer, Z. T. Kalbarczyk, P. H. Jones, D. A. Rennels, and R. Some. The effects of an ARMORbased SIFT environment on the performance and dependability of user applications. IEEE Transactions on Software Engineering, 30(4):257– 277, April 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1274045.

Weimer:2008:PFS

[WM08] W. Weimer and N. Mishra. Privately finding specifications. IEEE Transactions on Software Engineering, 34(1):21–32, January/February 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4359470.

Westfechtel:2001:LAU

[WMC01] B. Westfechtel, B. P. Munch, and R. Conradi. A layered architecture for uniform version management. IEEE Transactions on Software Engineering, 27(12):1111–1133, December 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=988710.

Wing:2000:RTC

[WO00] J. M. Wing and J. Ockerbloom. Respectful type converters. IEEE Transactions on Software Engineering, 26(7):579– 593, July 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=859529.

Wohlin:2001:GEI

[Woh01] C. Wohlin. Guest Editor’s introduction: Seventh International Software Metrics Symposium. IEEE Transactions on Software Engineering, 27(11): 961–962, November 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=965337.

Wang:2006:RAA

[WS06a] L. Wang and S. D. Stoller.

Runtime analysis of atomicity for multithreaded programs. IEEE Transactions on Software Engineering, 32(2):93– 110, February 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1599419.

Wang:2006:TCM

[WS06b] S. Wang and K. G. Shin. Task construction for model-based design of embedded control software. IEEE Transactions on Software Engineering, 32 (4):254–264, April 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1628971.

Wilson:2007:NAO

[WS07] S. P. Wilson and F. J.

Samaniego. Nonparametric analysis of the order-statistic model in software reliability. IEEE Transactions on Software Engineering, 33(3):198– 208, March 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4084137.

Wang:2004:BBS

[WSY^{+}04] Farn Wang, K. Schmidt, Fang Yu, Geng-Dian Huang, and Bow-Yaw Wang. BDD-based safety-analysis of concurrent software with pointer data structures using graph automorphism symmetry reduction. IEEE Transactions on Software Engineering, 30(6): 403–417, June 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1321062.

Weyuker:2000:EPT

[WV00] E. J. Weyuker and F. I. Vokolos. Experience with performance testing of software systems: issues, an approach, and case study. IEEE Transactions on Software Engineering, 26(12):1147–1156, December 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=888628.

Wing:2000:SIF

[WW00] J. M. Wing and J. Woodcock. Special issues for FM ’99: the First World Congress on Formal Methods in the Development of Computing Systems. IEEE Transactions on Software Engineering, 26(8): 673–674, August 2000. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=879806.

Xu:2003:TBD

[XC03] Songwen Xu and Yam San Chee. Transformation-based diagnosis of student programs for programming tutoring systems. IEEE Transactions on Software Engineering, 29(4): 360–384, April 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1191799.

Xie:2003:URF

[XE03] Yichen Xie and D. Engler. Using redundancies to find errors. IEEE Transactions on Software Engineering, 29(10): 915–928, October 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1237172.

Xu:2002:CSA

[XHD02] Dianxiang Xu, Xudong He, and Yi Deng. Compositional schedulability analysis of realtime systems using time Petri nets. IEEE Transactions on Software Engineering, 28(10): 984–996, October 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1041054.

Xie:2005:CIB

[XN05] T. Xie and D. Notkin. Checking inside the black box: regression testing by comparing value spectra. IEEE Transactions on Software Engineering, 31(10): 869–883, October 2005. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1542068.

Xu:2006:TDM

[XN06] D. Xu and K. E. Nygard. Threat-driven modeling and verification of secure software using aspect-oriented Petri nets. IEEE Transactions on Software Engineering, 32(4): 265–278, April 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1628972.

Xu:2003:FMB

[XS03] Haiping Xu and S. M. Shatz. A framework for model-based design of agent-oriented software. IEEE Transactions on Software Engineering, 29(1): 15–30, January 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1166586.

Xing:2005:AEH

[XS05] Z. Xing and E. Stroulia. Analyzing the evolutionary history of the logical design of object-oriented software. IEEE Transactions on Software Engineering, 31(10):850– 868, October 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1542067.

Xing:2007:AES

[XS07] Z. Xing and E. Stroulia.

API-evolution support with Diff–CatchUp. IEEE Transactions on Software Engineering, 33(12):818–836, December 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4359473.

Xu:2003:IVS

[Xu03]  J. Xu.  On inspection

and verification of software with timing requirements. IEEE Transactions on Software Engineering, 29(8):705– 720, August 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1223645.

Xie:2003:SEI

[XY03] Min Xie and Bo Yang. A study of the effect of imperfect debugging on software development cost. IEEE Transactions on Software Engineering, 29 (5):471–473, May 2003. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1199075.

Xu:2003:FAM

[XYDD03] Dianxiang Xu, Jianwen Yin, Yi Deng, and Junhua Ding. A formal architectural model for logical agent mobility. IEEE Transactions on Software Engineering, 29(1):31– 45, January 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1166587.

Yacoub:2002:MAL

[YA02] S. M. Yacoub and H. H. Ammar. A methodology for architecture-level reliability risk analysis. IEEE Transactions on Software Engineering, 28(6):529–547, June 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1010058.

Yang:2002:CNA

[Yan02] Wuu Yang. A classification of noncircular attribute grammars based on the look-ahead behavior. IEEE Transactions on Software Engineering, 28 (3):210–227, March 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=991318.

Yen:2001:DMI

[YBT01] I-Ling Yen, F. B. Bastani, and D. J. Taylor. Design of multi-invariant data structures for robust shared accesses in multiprocessor systems. IEEE Transactions on Software Engineering, 27(3): 193–207, March 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=910857.

Ye:2009:AAS

[YCCX09] C. Ye, S. C. Cheung, W. K.

Chan, and C. Xu. Atomicity analysis of service composition across organizations. IEEE Transactions on Software Engineering, 35(1):2–28, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4641941.

Yilmaz:2006:CAE

[YCP06] C. Yilmaz, M. B. Cohen, and A. A. Porter. Covering arrays for efficient fault characterization in complex configuration spaces. IEEE Transactions on Software Engineering, 32 (1):20–34, January 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1583600.

Yang:2008:SUS

[YHJ08] B. Yang, H. Hu, and L. Jia. A study of uncertainty in software cost and its impact on optimal software release time. IEEE Transactions on Software Engineering, 34(6):813–825, November/December 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4553721.

Ying:2004:PSC

[YMNCC04] A. T. T. Ying, G. C. Murphy, R. Ng, and M. C. Chu-Carroll. Predicting source code changes by mining change history. IEEE Transactions on Software Engineering, 30(9): 574–586, September 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1324645.

Yilmaz:2007:RES

[YPK^{+}07] C. Yilmaz, A. Porter, A. S. Krishna, A. M. Memon, D. C. Schmidt, A. S. Gokhale, and B. Natarajan. Reliable effects screening: A distributed continuous quality assurance process for monitoring performance degradation in evolving software systems.

IEEE Transactions on Software Engineering, 33(2):124– 141, February 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4052587.

Yu:2004:CCC

[YSCO04] L. Yu, S. R. Schach, K. Chen, and J. Offutt. Categorization of common coupling and its application to the maintainability of the Linux kernel. IEEE Transactions on Software Engineering, 30(10): 694–706, October 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339279.

Yadav:2001:CFI

[YSW^{+}01] S. B. Yadav, N. G. Shaw, L. Webb, C. Sutcu, T. L. Roberts, M. L. Gibson, R. K. Rainer, and K. T. Fields. Comments on “Factors that impact implementing a system development methodology” [with reply]. IEEE Transactions on Software Engineering, 27(3): 279–286, March 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=910863. See [RGFR98].

Zeng:2004:QAM

[ZBN^{+}04] Liangzhao Zeng, B. Benatallah, A. H. H. Ngu, M. Dumas, J. Kalagnanam, and H. Chang. QoS-aware middleware for Web services composition. IEEE Transactions on Software Engineering, 30 (5):311–327, May 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1291834.

Zheng:2006:MMC

[ZC06] Y. Zheng and A. T. S. Chan. MobiGATE: a mobile computing middleware for the active deployment of transport services. IEEE Transactions on Software Engineering, 32(1): 35–50, January 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1583601.

Zave:2009:CCI

[ZC09] P. Zave and E. Cheung. Compositional control of IP media. IEEE Transactions on Software Engineering, 35(1):46–66, January/February 2009. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4569853.

Zhang:2004:UHG

[ZD04] F. Zhang and E. H. D’Hollander. Using hammock graphs to structure programs. IEEE Transactions on Software Engineering, 30(4):231–245, April 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1274043.

Zeller:2002:SIF

[ZH02] A. Zeller and R. Hildebrandt. Simplifying and isolating failure-inducing input. IEEE Transactions on Software Engineering, 28(2):183– 200, February 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=988498.

Zhang:2008:DSF

[Zha08]  H. Zhang.  On the dis-

tribution of software faults. IEEE Transactions on Software Engineering, 34(2):301– 302, March/April 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4407730.

Zhou:2006:EAO

[ZL06] Yuming Zhou and Hareton Leung. Empirical analysis of object-oriented design metrics for predicting high and low severity faults. IEEE Transactions on Software Engineering, 32(10):771– 789, October 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1717471.

Zhou:2007:MMM

[ZLW07] Y. Zhou, H. Leung, and P. Winoto. MNav: A Markov model-based Web site navigability measure. IEEE Transactions on Software Engineering, 33(12):869–890, December 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4359469.

Zhou:2009:EPC

[ZLX09] Y. Zhou, H. Leung, and B. Xu. Examining the potentially confounding effect of class size on the associations between object-oriented metrics and change-proneness. IEEE Transactions on Software Engineering, 35(5):607–623, September/October 2009. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4967613.

Zachariadis:2006:SCS

[ZME06] S. Zachariadis, C. Mascolo, and W. Emmerich. The SATIN Component System — a metamodel for engineering adaptable mobile systems. IEEE Transactions on Software Engineering, 32(11):910– 927, November 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4015513.

Zuberi:2001:ESM

[ZS01] K. M. Zuberi and K. G.

Shin. EMERALDS: a smallmemory real-time microkernel. IEEE Transactions on Software Engineering, 27(10): 909–928, October 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=962561.

Zheng:2008:PME

[ZWL08] T. Zheng, C. M. Woodside, and M. Litoiu. Performance model estimation and tracking using optimal filters. IEEE Transactions on Software Engineering, 34(3):391– 406, May/June 2008. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4515874.

Zheng:2006:VSA

[ZWN^{+}06] J. Zheng, L. Williams, N. Nagappan, W. Snipes, J. P. Hudepohl, and M. A. Vouk. On the value of static analysis for fault detection in software. IEEE Transactions on Software Engineering, 32(4): 240–253, April 2006. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

ieee.org/stamp/stamp.jsp?
arnumber=1628970.

Zhang:2002:CAW

[ZX02] Lu Zhang and Dan Xie. Comments on “On the applicability of Weyuker property 9 to object-oriented structural inheritance complexity metrics”. IEEE Transactions on Software Engineering, 28(5): 526–527, May 2002. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1000454. See com-

ments [GR01].

Zhang:2003:OTT

[ZyC03] Fan Zhang and To yat Cheung. Optimal transfer trees and distinguishing trees for testing observable nondeterministic finite-state machines. IEEE Transactions on Software Engineering, 29(1):1– 14, January 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1166585.

Zhang:2007:CDM

[ZZ07] H. Zhang and X. Zhang. Comments on “Data Mining Static Code Attributes to Learn Defect Predictors”. IEEE Transactions on Software Engineering, 33(9):635– 637, September 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=4288196.  See

[MGF07] and response [MDDG07].

Zhang:2001:DCA

[ZZC01] K. Zhang, D. Q. Zhang, and J. Cao. Design, construction, and application of a generic visual language generation environment. IEEE Transactions on Software Engineering, 27 (4):289–307, April 2001. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=917521.

Zimmermann:2005:MVH

[ZZWD05] T. Zimmermann, A. Zeller, P. Weissgerber, and S. Diehl. Mining version histories to guide software changes. IEEE Transactions on Software Engineering, 31(6):429–445, June 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1463228.

Zou:2007:IUC

[ZZZ07] Y. Zou, Q. Zhang, and X. Zhao. Improving the usability of E-commerce applications using business processes. IEEE Transactions on Software Engineering, 33(12):837– 855, December 2007. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.

                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=4359465.